### v1

# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

Abstract
Understanding the dynamical origin of high–temperature superconductivity remains a central challenge of strongly correlated quantum matter. Conventional approaches to quantum criticality assume overdamped Markovian dissipation governed by Ohmic Landau damping. Here we show that infrared collective dynamics is instead generically controlled by the time–scale density of states (TDOS) of relaxation modes. Within the Martin–Siggia–Rose–Janssen–De Dominicis formalism, we derive an exact spectral representation of the collective susceptibility in terms of the TDOS. A finite infrared TDOS defines a new universality class of memory–dominated critical dynamics characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. This spectral reorganization produces a strong infrared amplification of the particle–particle channel, converting the marginal instabilities of BCS and Eliashberg theories into algebraic superconducting transitions. As a result, the transition temperature scales linearly with the infrared spectral weight of slow collective modes, naturally yielding superconducting domes and Uemura scaling without invoking bosonic glue or fine tuning. The same slow–mode reservoir governs anomalous normal–state dynamics, including long–time correlations and strange–metal behavior, providing a unified description of thermodynamic and dynamical phenomena in correlated superconductors. Our results establish dynamical spectral organization as a fundamental principle of quantum critical matter and identify memory–dominated criticality as a generic route to high–temperature superconductivity.
I I. Introduction
High–temperature superconductivity remains one of the central unsolved problems of strongly correlated quantum matter [1, 2]. Despite decades of intensive theoretical and experimental effort, no consensus has emerged on the microscopic mechanism responsible for robust pairing far above conventional phonon scales. A central difficulty lies in the intimate proximity of superconductivity to quantum critical regimes, where strong collective fluctuations dominate low–energy dynamics.
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such low-dimensional reductions necessarily generate a finite set of soft modes and therefore preclude the emergence of a dense continuum of relaxation time scales. A genuine slow-mode spectrum can only arise when the full high-dimensional collective dynamical state space is retained.
In this work we show that infrared collective dynamics is generically controlled not by a single overdamped mode but by the time–scale density of states (TDOS) of correlated internal relaxation modes. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [12, 13, 14], we derive an exact spectral representation of the collective susceptibility in terms of the TDOS.
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a new universality class governed by self–organized slow modes rather than fast external baths.
Crucially, we show that such flat relaxation mode spectra arise naturally in strongly correlated collective systems, including self–organized dynamical manifolds explored in our previous work [15, 16, 17]. The resulting long–memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a natural route to strong superconductivity without invoking finely tuned bosonic mediators.
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation mode spectra determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 18, 19]. Charge, spin, and pairing channels reorganize simultaneously within the same slow–mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a common infrared spectral continuum.
Our results establish the relaxation mode spectra as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified perspective on collective behavior across strongly correlated quantum matter [5, 7, 10].
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. We conclude in Section VI with a summary and outlook.
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory. (c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
Strongly correlated electronic matter hosts many intertwined collective fluctuation channels, including charge, spin, current, pairing, and phase–coherence sectors. These degrees of freedom continuously exchange energy and information and cannot, in general, be reduced to a small set of independent order parameters.
To capture this collective complexity in a unified manner, we represent the correlated system as a high–dimensional interacting dynamical system whose instantaneous collective configuration is encoded in a state vector
$x(t)\in\mathbb{R}^{N},$ (1)
which evolves under nonlinear many–body feedback.
The coarse-grained dynamics may therefore be written as a stochastic field theory
$\dot{x}(t)=F(x(t))+\xi(t),$ (2)
where $F$ encodes correlated interactions and $\xi$ is Gaussian noise with
$\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=2D\,\delta_{ij}\delta(t-t^{\prime}).$ (3)
II.1 A. Dynamical stability spectrum and time-scale density of states
Near correlated critical regimes the system does not soften a single collective mode but undergoes a reorganization of its full dynamical stability structure. Linearizing around a fluctuating metastable configuration $x^{*}$,
$\delta\dot{x}(t)=-A\,\delta x(t)+\xi(t),$ (4)
with
$A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}},$ (5)
yields a spectrum of collective relaxation modes
$Av_{i}=\lambda_{i}v_{i},$ (6)
whose retarded propagators are
$G_{i}(\omega)=\frac{1}{-i\omega+\lambda_{i}}.$ (7)
The infrared structure of correlated dynamics is therefore controlled by the distribution of relaxation rates $\lambda_{i}$. We characterize this spectrum by the time-scale density of states,
$\rho(\lambda)=\Big\langle\sum_{i}\delta(\lambda-\lambda_{i})\Big\rangle.$ (8)
In weakly interacting systems the TDOS vanishes as $\lambda\to 0$, reflecting a finite number of quasiparticle-like modes. By contrast, strongly correlated critical regimes generically generate an extensive accumulation of near-marginal directions in phase space,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (9)
forming a slow-mode reservoir that governs infrared collective dynamics.
This accumulation is the dynamical signature of correlated criticality: the system approaches marginal stability along many collective directions rather than along a single order-parameter coordinate.
II.2 B. Exact spectral representation from MSRJD field theory
To compute collective response functions while retaining the full dynamical spectrum, we employ the Martin–Siggia–Rose–Janssen–De Dominicis formalism associated with Eq. (2),
$S[x,\tilde{x}]=\int dt\,\tilde{x}(t)\big[\dot{x}(t)-F(x(t))\big]-D\tilde{x}^{2}(t).$ (10)
We focus on a slow collective field $\phi(\mathbf{r},t)$ obtained by coarse graining the full state vector and regard the remaining fast correlated degrees of freedom as internal relaxation modes.
The reorganized stability spectrum may be represented exactly as a continuum of Ornstein–Uhlenbeck modes $X_{\lambda}$ labeled by their decay rates,
$\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (11)
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda}.$ (12)
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action becomes
$S=$ $\int dt\,d^{d}r\,\tilde{\phi}\Big[\dot{\phi}+(r+c\nabla^{2})\phi-\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}\Big]-D_{\phi}\tilde{\phi}^{2}$
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi]$
$-\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,D_{\lambda}\tilde{X}_{\lambda}^{2}.$ (13)
Since the reservoir variables enter quadratically, they can be integrated out exactly. In Fourier space this yields the effective collective action
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots,$ (14)
with retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (15)
Separating real and imaginary parts,
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\frac{\omega}{\lambda^{2}+\omega^{2}},$ (16)
gives
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (17)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{1}{\lambda^{2}+\omega^{2}}.$ (18)
Identifying the TDOS as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (19)
the collective inverse susceptibility becomes
$\chi_{R}^{-1}(\omega,\mathbf{q})$ $=r+cq^{2}+\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}$ (20)
$\quad+i\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$
Equation (20) constitutes an exact spectral representation of collective infrared dynamics in terms of the dynamical stability spectrum. The complete field-theoretic derivation of the collective susceptibility kernel from the MSRJD formalism, including the explicit Gaussian integration over the relaxation-mode reservoir and the resulting self-energy structure, is presented in Appendix A.
II.3 C. Infrared universality classes
[…]
$\rho(\lambda)\sim\lambda^{\alpha},\qquad\lambda\to 0,$ (21)
the dissipative part of the collective kernel follows directly from the spectral representation in Eq. (20),
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (22)
Here $\omega$ denotes the frequency conjugate to the collective relaxation dynamics obtained after integrating out the slow–mode reservoir in the MSRJD formulation, rather than a microscopic quasiparticle energy scale.
It is therefore natural to write the long–wavelength dynamical response in the scaling form
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (23)
in direct analogy with the conventional Hertz–Millis representation. The TDOS scaling immediately implies the universal identification $s=\alpha$.
Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$. Requiring both contributions to remain of comparable order at criticality ($r=0$) fixes the dynamical exponent to
$z=\frac{2}{s}.$ (24)
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the presence of an extensive reservoir of slow collective relaxation modes. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy organization of relaxation modes encoded in the TDOS.
In particular, when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (25)
the dissipative kernel simplifies to
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{0}}{\lambda^{2}+\omega^{2}}$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (26)
while the reactive part acquires a logarithmic infrared singularity,
$\mathrm{Re}\,\chi_{R}^{-1}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{\lambda}{\lambda^{2}+\omega^{2}}\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (27)
Collective infrared dynamics is therefore governed by the marginal non-Markovian kernel
$\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (28)
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, which is derived explicitly in Appendix B.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow internal modes.
The TDOS thus provides a unified organizing principle for infrared critical dynamics, with different universality classes emerging from the scaling of the dynamical stability spectrum itself.
III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
The emergence of a slow–mode reservoir qualitatively reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
III.1 A. Marginal pairing from Ohmic collective dynamics
Near metallic quantum critical points, fermions couple to overdamped collective modes with retarded propagator
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (29)
The linear frequency dependence reflects Landau damping by the electronic continuum and produces a pairing kernel scaling as
$D(\omega)\sim\frac{1}{|\omega|}.$ (30)
The infrared shell correction to the dimensionless Cooper–channel coupling $g$ therefore reads
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=g^{2}\ln b.$ (31)
Introducing the RG scale $l=\ln b$ yields
$\frac{dg}{dl}=Cg^{2},$ (32)
with $C>0$.
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes rapidly dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
III.2 B. Infrared reorganization by slow collective modes
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir within the MSRJD framework, the retarded collective susceptibility acquires the exact spectral form
$\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (33)
For a flat TDOS in the infrared, $\rho(\lambda\to 0)=\rho_{0}$, the effective polarization kernel in the Cooper channel inherits the infrared scaling
$\Pi(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{1}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},\qquad|\omega|\to 0.$ (34)
Unlike the conventional Ohmic case, this singular behavior does not originate from Landau damping by fast electronic continua. Instead, it emerges from the accumulation of an extensive continuum of near–marginal relaxation modes encoded in a flat TDOS. The resulting kernel corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$.
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects instantaneous overdamped dissipation arising from coupling to a fast electronic bath [20, 21]. Here, by contrast, the same infrared scaling is a consequence of self–organized spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales. The slow–mode reservoir therefore converts the Cooper channel into a relevant infrared interaction.
III.3 C. Relevant infrared amplification of the Cooper channel
The infrared shell correction in the memory–dominated regime becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}\ln(\Lambda/\omega^{\prime})}.$ (35)
Introducing $u=\ln(\Lambda/\omega^{\prime})$ yields
$dg\simeq\alpha g\,dl,$ (36)
with $\alpha>0$ proportional to the infrared TDOS weight.
The resulting RG flow reads
$\frac{dg}{dl}=\alpha g,$ (37)
with solution $g(l)=g_{0}e^{\alpha l}$.
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix C.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and dynamically amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral self–organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
IV IV. Pairing from Memory–Dominated Critical Dynamics
The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes. Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum. The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes. Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
In this section we show explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, yielding power–law growth of the effective pairing interaction and algebraic scaling of the transition temperature.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes. Starting from the Hubbard Hamiltonian,
$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (38)
projection to the strong–coupling sector yields the superexchange interaction
$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (39)
favoring local singlet formation.
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (40)
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies this preexisting pairing tendency.
IV.2 B. Memory–dominated Cooper kernel from flat TDOS
Because single–particle propagators admit a relaxational spectral representation controlled by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (41)
When the slow–mode reservoir develops,
$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (42)
the infrared behavior becomes
$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (43)
This nonanalytic enhancement reflects the accumulation of collective memory across all time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel is dominated by long–time memory effects generated by the slow–mode TDOS.
IV.3 C. Algebraic superconducting instability and universality
The superconducting transition follows from the Thouless criterion
$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (44)
which yields
$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (45)
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir. High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
Because the TDOS exponent controls infrared dynamics independently of microscopic details, the pairing amplification mechanism is universal across correlated systems exhibiting flat relaxation spectra. The slow–mode reservoir therefore acts as a universal dynamical amplifier, converting electronic correlation energy into enhanced superconducting transition temperatures without invoking specific bosonic glue.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality. The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest. Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning. (b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$. Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
V V. Experimental consequences of memory–dominated critical dynamics
[…]
V.1 A. Infrared spectral control and automatic emergence of the dome
The linearized transition scale derived in Eq. (45) shows that the superconducting temperature is controlled by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. Near correlated quantum criticality, $\rho_{0}$ is not constant but is governed by universal scaling.
Let $p$ denote a tuning parameter (e.g., doping or pressure) and $p_{c}$ the critical point. The mass term of the collective kernel obeys
$r(p)\sim|p-p_{c}|,$ (46)
so that the corresponding infrared cutoff scale takes the universal form
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{c}|^{z\nu},$ (47)
with $z$ the dynamical exponent and $\nu$ the correlation-length exponent. Away from criticality the slow–mode reservoir is therefore truncated in the infrared, yielding an effective low-energy TDOS weight
$\rho_{0}(p)\sim\int_{0}^{\Lambda(p)}d\lambda\,\rho(\lambda).$ (48)
For a memory–dominated regime in which the TDOS remains finite at the origin, $\rho(\lambda\to 0)=\rho_{*}$, one obtains parametrically
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{p_{0}}\right),$ (49)
where $f$ is a universal decreasing scaling function and $p_{0}$ denotes a nonuniversal crossover scale.
Substituting Eq. (49) into Eq. (45) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{x_{0}}\right),$ (50)
which is maximal near criticality and suppressed on both sides, producing a superconducting dome without fine tuning.
The dome structure is thus a direct consequence of the competition between intrinsic short–range electronic pairing and the finite infrared extent of the slow–mode reservoir away from criticality. Importantly, no competing order parameters or additional microscopic suppression mechanisms are required: suppression of $T_{c}$ arises purely from the truncation of long–time memory encoded in the TDOS. In this interpretation, optimal doping corresponds to maximal infrared spectral reorganization rather than to the peak of any specific bosonic fluctuation.
V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
While the microscopic pairing seed is set by the intrinsic electronic scale $E_{\rm pair}$, the establishment of superconductivity requires global phase coherence of the Cooper collective field. In systems with low superfluid density, the transition may be controlled by phase stiffness rather than by gap formation.
The effective phase–only action for the Cooper field reads
$S_{\theta}=\frac{\rho_{s}}{2}\int d\tau d^{d}r\left[(\partial_{\tau}\theta)^{2}+v_{s}^{2}(\nabla\theta)^{2}\right],$ (51)
where $\rho_{s}$ is the superfluid stiffness. The phase–ordering temperature therefore scales as
$T_{\theta}\sim\rho_{s}.$ (52)
Within the slow–mode reservoir framework, the stiffness is controlled by the same infrared spectral weight that enhances pairing. Integrating out amplitude fluctuations yields the scaling
$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (53)
reflecting the collective participation of slow modes in coherent transport. Combining Eqs. (52) and (53) gives
$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (54)
establishing a direct proportionality between the superconducting transition temperature and superfluid density.
Equation (54) reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors. In contrast to BCS superconductors, where phase stiffness is large and irrelevant to $T_{c}$, correlated systems operate in a regime where infrared collective dynamics simultaneously controls pairing and coherence. In the present framework, this is not an additional assumption but follows directly from the single infrared control parameter $\rho_{0}$.
V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
A distinctive prediction of the present theory concerns dynamical response in the normal state. As shown in Sec. II, a flat TDOS produces a marginal nonanalytic retarded kernel with
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)\propto\mathrm{sgn}(\omega),$ (55)
corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
Such long–time tails naturally generate 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
V.4 D. Unified infrared scaling picture
We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The TDOS controls: (i) the extent of the slow–mode reservoir away from criticality [Eq. (47)], (ii) the dynamical amplification of intrinsic pairing tendencies [Eq. (50)], (iii) the superfluid stiffness and phase coherence [Eq. (54)], and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a self–organized dynamical amplifier that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) should simultaneously display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
VI VI. Discussion
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is governed not by local Markovian dissipation but by the time–scale density of states of collective relaxation modes. Criticality corresponds to a reorganization of the entire dynamical spectrum, and universality classes are naturally classified by the low–energy scaling of the TDOS. In particular, a finite TDOS at vanishing relaxation rate defines a memory–dominated infrared fixed point characterized by long–time kernels $K(t)\sim 1/t$.
This perspective clarifies the limited domain of validity of conventional coarse–grained field theories. Away from criticality, relaxation spectra are gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective order parameters. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending to zero. Internal modes cannot be integrated out: the full relaxation spectrum remains dynamically active at all time scales. The breakdown of Markovian coarse–graining is therefore a universal consequence of dynamical criticality rather than a model-dependent artifact.
This observation explains why Hertz–Millis–type approaches, which assume overdamped local dynamics of a small set of collective fields, cannot capture the infrared structure of strongly correlated critical matter. Criticality does not correspond to the softening of a single bosonic mode but to a spectral condensation of near–marginal collective degrees of freedom. The TDOS thus emerges as the natural organizing variable of infrared dynamics, replacing phenomenological damping parameters and providing a microscopic foundation for anomalous relaxation and transport.
A central physical consequence of this spectral reorganization is the qualitative enhancement of superconducting instabilities. Rather than requiring finely tuned bosonic mediators, superconductivity emerges generically from dynamical amplification of intrinsic electronic pairing tendencies by the slow–mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
A single infrared parameter, $\rho_{0}$, governs thermodynamic and dynamical phenomena alike. The superconducting dome reflects truncation of long–time memory away from criticality, Uemura scaling follows from the shared spectral origin of pairing and phase coherence, and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a self-organized dynamical amplifier that simultaneously controls fluctuations, transport, and superconducting order.
Although electrons propagate as Bloch quasiparticles on a crystalline lattice, collective particle–hole and particle–particle excitations form extended many-body coordinates not constrained to lattice energy eigenstates. In strongly correlated regimes these collective modes are generically overdamped by the electronic continuum, rendering a relaxation-rate description more fundamental than an energy-eigenmode picture. While conventional lattice theories organize collective behavior in terms of Hamiltonian eigenmodes, the present framework shows that near criticality the relevant organizing structure is instead the dynamically self-organized spectrum of relaxation rates.
This relaxation–rate organization is not a phenomenological assumption but a direct consequence of the fundamental structure of nonequilibrium quantum and classical dynamics. As shown in the Appendix F, both Liouvillian evolution in open quantum systems and generalized memory–kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues rather than by Hamiltonian energy eigenstates.
Even in the absence of spatial degrees of freedom, the Laplace–domain representation of non-Markovian dynamics generically yields continuous spectra of relaxation rates whenever long–time memory kernels are present. The emergence of a dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, not a model-dependent artifact of coarse–graining.
From this perspective, energy eigenmodes describe conservative reversible dynamics, while relaxation spectra encode the physically relevant irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points is therefore more naturally classified by its relaxation-rate spectrum than by its Hamiltonian excitation spectrum.
The TDOS framework introduced here thus elevates relaxation modes to the status of fundamental collective coordinates, providing a principled dynamical basis for infrared universality, memory effects, and emergent macroscopic coherence.
Away from criticality the dynamical spectrum collapses onto a trivial fast mode. Near memory-dominated criticality an extensive continuum of slow collective modes emerges and governs macroscopic coherence. Superconducting pairing in this regime corresponds to condensation within a slow-mode reservoir rather than binding mediated by a single sharp bosonic excitation.
This unified picture resolves long-standing puzzles of unconventional superconductivity, including the coexistence of strange-metal behavior with robust pairing and the apparent universality of dome-shaped phase diagrams across disparate materials. Rather than invoking competing orders or material-specific bosonic glue, these phenomena arise naturally from universal infrared spectral structure.
More broadly, the TDOS framework establishes dynamical spectral organization as a fundamental principle of correlated quantum matter. Memory-dominated criticality represents a new universality class beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and self-organized critical systems.
Future directions include quantitative determination of TDOS spectra in microscopic models and experiments, extension to nonequilibrium driven systems, and exploration of engineered platforms in which slow-mode reservoirs can be dynamically controlled. In this light, high-temperature superconductivity appears not as an anomaly requiring special microscopic ingredients, but as a generic consequence of infrared spectral self-organization in strongly correlated matter.
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Recent experimental efforts toward room-temperature superconductivity have largely focused on engineering stronger microscopic pairing interactions through pressure, lattice design, or exotic chemical compositions. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. Our results show that the decisive control parameter for high transition temperatures is not the magnitude of an attractive interaction but the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
A dense reservoir of near-marginal slow modes fundamentally reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this perspective, many existing high-pressure hydride systems and correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
The present theory therefore suggests a new design principle for superconductivity beyond material-specific glue engineering: the deliberate stabilization of dynamical phases supporting extensive slow-mode continua. Controlling electronic structure, frustration, dimensionality, and feedback mechanisms to promote near-marginal collective dynamics may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this view, corresponds to a dynamical phase characterized by maximal infrared memory rather than by extreme coupling strength.
VII VII. Conclusion
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. Formulating collective dynamics in terms of the time–scale density of states reveals that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting the marginal instabilities of conventional theories into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified explanation of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise generically from infrared spectral reorganization, without invoking specific bosonic glue, competing orders, or fine tuning.
More broadly, our results recast quantum criticality as a constructive dynamical process that enhances collective quantum phenomena through self–organized memory effects. Memory–dominated criticality thus represents a new paradigm beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity emerges not as an exception requiring special microscopic ingredients, but as a natural consequence of infrared spectral self–organization in strongly correlated matter. Future exploration of TDOS-controlled dynamics may enable systematic design of materials and devices that exploit memory–dominated criticality to stabilize and enhance collective quantum states.
Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during early development.
Supplementary Materials
Appendix A Appendix A: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
In this appendix we derive, without omitting algebraic steps, how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
(I) Coupled stochastic dynamics: collective variable plus OU reservoir: We consider a collective degree of freedom $x(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{x}(t)$ $=-r\,x(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S1)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,x(t)+\xi_{\lambda}(t).$ (S2)
The noises are assumed Gaussian, stationary, and white:
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{x}\,\delta(t-t^{\prime}),$ (S3)
$\langle\xi_{\lambda}(t)\,\xi_{\lambda^{\prime}}(t^{\prime})\rangle$ $=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\,\delta(t-t^{\prime}).$ (S4)
No equilibrium fluctuation–dissipation relation is assumed; $D_{\lambda}$ may be arbitrary. Our goal is to integrate out $\{X_{\lambda}\}$ and obtain an exact effective MSRJD action for (x,x~)($\tilde{x}$) containing a memory kernel.
(II) MSRJD action for the coupled system: For a generic additive-noise Langevin equation $\dot{y}=f(y)+\zeta$ with $\langle\zeta(t)\zeta(t^{\prime})\rangle=2D\,\delta(t-t^{\prime})$, the MSRJD functional integral may be written (Ito convention) as
$Z$ $=\int\mathcal{D}y\,\mathcal{D}\tilde{y}\;$ (S5)
$\exp\!\left[-\int dt\;\tilde{y}(t)\big(\dot{y}(t)-f(y(t))\big)+\int dt\;D\,\tilde{y}(t)^{2}\right],$
up to an overall normalization independent of $y$. Applying (S5) to (S1)–(S2) gives
$Z=\int\mathcal{D}x\,\mathcal{D}\tilde{x}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[x,\tilde{x},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S6)
with total action
$S=S_{x}+\int d\lambda\;S_{\lambda},$ (S7)
where
$S_{x}$ $=\int dt\;\tilde{x}(t)\!\left[\dot{x}(t)+r\,x(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{x}\,\tilde{x}(t)^{2},$ (S8)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,x(t)\right]$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S9)
(III) Isolating a single $\lambda$-mode contribution: Fix $\lambda$ and collect all terms in $S$ that contain $X_{\lambda}$ or $\tilde{X}_{\lambda}$. From (S8) we extract
$S_{x}\supset-\int dt\;\tilde{x}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S10)
Combining (S10) with (S9) yields the $\lambda$-sector action
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};x,\tilde{x}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S11)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;g(\lambda)\,\tilde{x}(t)\,X_{\lambda}(t)$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
In what follows, $x,\tilde{x}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: emergence of a delta constraint: Define the linear operator
$\mathcal{L}_{\lambda}\equiv\partial_{t}+\lambda.$ (S12)
Then the first term in (S11) is
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=\int dt\;\tilde{X}_{\lambda}(t)\big(\partial_{t}X_{\lambda}(t)+\lambda X_{\lambda}(t)\big).$ (S13)
We now rewrite this term so that $X_{\lambda}$ appears without derivatives. Using integration by parts,
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)$ (S14)
$\quad=\big[\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)\big]_{t_{i}}^{t_{f}}-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$
Assuming boundary terms vanish (e.g. $t_{i}\to$-$\infty$, $t_{f}\to+\infty$ with sufficient decay, or appropriate causal boundary conditions), we drop the bracketed term and obtain
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$ (S15)
Therefore,
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t)$ (S16)
$+\int dt\;\lambda\,\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)=\int dt\;X_{\lambda}(t)\big(-\partial_{t}+\lambda\big)\tilde{X}_{\lambda}(t).$
Define the adjoint operator
$\mathcal{L}_{\lambda}^{\dagger}\equiv-\partial_{t}+\lambda.$ (S17)
Then (S11) becomes
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t)\Big)$ (S18)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
Now $X_{\lambda}$ enters only linearly. Hence its functional integral produces a functional delta constraint:
$\int\mathcal{D}X_{\lambda}\;\exp\!\left\{-\int dt\;X_{\lambda}(t)\,A(t)\right\}$ $\propto\delta[A(t)],$ (S19)
where here
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t).$ (S20)
Therefore,
[…]
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $x,\tilde{x}$ and can therefore be absorbed into the overall normalization of $Z$.
(V) Solving the delta constraint for $\tilde{X}_{\lambda}$ via Green functions: The constraint
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{x}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S22)
is a linear inhomogeneous ODE for $\tilde{X}_{\lambda}(t)$. Introduce the Green function $G_{\lambda}^{A}(t-t^{\prime})$ solving
$(-\partial_{t}+\lambda)\,G_{\lambda}^{A}(t-t^{\prime})=\delta(t-t^{\prime}).$ (S23)
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $x$-sector response).
The solution with advanced support is
$G_{\lambda}^{A}(t)=\Theta(-t)\,e^{\lambda t}.$ (S24)
(Indeed, for $t\neq 0$, $(-\partial_{t}+\lambda)G_{\lambda}^{A}=0$ and the jump at $t=0$ produces the delta distribution.) Multiplying (S22) by $G_{\lambda}^{A}$ and integrating, we obtain
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime}).$ (S25)
Substituting (S25) into (S21) eliminates $\tilde{X}_{\lambda}$ entirely and yields an effective contribution to the (x,x~)($\tilde{x}$) action.
(VI) Effective action: memory kernel (self-energy) and induced colored noise: We now evaluate the two remaining terms in the exponent of (S21) using (S25).
1. The $\tilde{X}_{\lambda}x$ term (deterministic memory kernel).
Start from
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t).$ (S26)
Insert (S25):
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime})\,x(t).$ (S27)
Relabel integration variables $(t,t^{\prime})\to(t^{\prime},t)$:
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{x}(t)\,x(t^{\prime}).$ (S28)
Using the identity $G_{\lambda}^{A}(t^{\prime}-t)=G_{\lambda}^{R}(t-t^{\prime})$ with the retarded Green function
$G_{\lambda}^{R}(t)\equiv\Theta(t)\,e^{-\lambda t},\qquad(\partial_{t}+\lambda)G_{\lambda}^{R}(t)=\delta(t),$ (S29)
we obtain
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)$ $=\int dt\,dt^{\prime}\;\tilde{x}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,x(t^{\prime}).$ (S30)
Thus a single OU mode contributes the causal memory kernel
$K_{\lambda}(t-t^{\prime})=g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S31)
2. The $\tilde{X}_{\lambda}^{2}$ term (induced colored noise kernel).
Next consider
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S32)
Insert (S25):
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$ (S33)
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{x}(t^{\prime})\,\tilde{x}(t^{\prime\prime}).$ (S34)
Define the induced (generally nonlocal) noise kernel
$\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\equiv D_{\lambda}\,g(\lambda)^{2}\int dt\;G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime}),$ (S35)
so that
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{x}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{x}(t^{\prime\prime}).$ (S36)
In equilibrium one may impose a fluctuation–dissipation relation linking $\mathcal{N}_{\lambda}$ and $K_{\lambda}$, but this is not required for the deterministic self-energy structure derived below.
3. Summing the reservoir and assembling $S_{\mathrm{eff}}$.
Combining (S30)–(S36) for all $\lambda$ yields the exact effective action
$S_{\mathrm{eff}}[x,\tilde{x}]=\int dt\;\tilde{x}(t)\big(\dot{x}(t)+rx(t)\big)-\int dt\;D_{x}\,\tilde{x}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{x}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{x}(t^{\prime}),$ (S37)
where the total memory kernel is the reservoir sum
$K(t-t^{\prime})=\int d\lambda\;K_{\lambda}(t-t^{\prime})=\int d\lambda\;g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})},$ (S38)
and likewise $\mathcal{N}=\int d\lambda\,\mathcal{N}_{\lambda}$.
(VI) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S37) to the standard response-function form, we Fourier transform using
$f(\omega)=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,f(t),\qquad f(t)=\int\frac{d\omega}{2\pi}\;e^{-i\omega t}\,f(\omega).$ (S39)
The bilinear term involving $K$ becomes
$\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{x}(-\omega)\,K(\omega)\,x(\omega),$ (S40)
where $K(\omega)$ is the Fourier transform of the causal kernel (S38):
$K(\omega)$ $=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,K(t)=\int d\lambda\;g(\lambda)^{2}\int_{0}^{\infty}dt\;e^{i\omega t}e^{-\lambda t}.$ (S41)
The time integral is elementary:
$\int_{0}^{\infty}dt\;e^{-(\lambda-i\omega)t}$ $=\left[\frac{-1}{\lambda-i\omega}e^{-(\lambda-i\omega)t}\right]_{0}^{\infty}=\frac{1}{\lambda-i\omega},$
$\qquad(\mathrm{Re}\,\lambda>0).$ (S42)
Substituting (S42) into (S41) gives the exact spectral representation
$K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S43)
In the MSRJD response sector, the retarded inverse propagator for $x$ is read off from the $\tilde{x}(-\omega)x(\omega)$ coefficient in $S_{\mathrm{eff}}$:
$G_{R}^{-1}(\omega)=-i\omega+r+K(\omega).$ (S44)
It is therefore natural to define the (retarded) self-energy as
$\Sigma_{R}(\omega)\equiv K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S45)
(Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
Equations (S43)–(S45) demonstrate that integrating out a continuum of OU relaxation modes produces a frequency- dependent self-energy whose analytic structure is determined by the relaxation-rate spectrum and couplings.
(VII) Connection to TDOS (relaxation-rate density of states): If the couplings are sufficiently structureless over the low-$\lambda$ sector, one may absorb $g(\lambda)^{2}$ into an effective spectral density. Defining
$\mathcal{J}(\lambda)\equiv g(\lambda)^{2},$ (S46)
we can write
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega}.$ (S47)
Equivalently, starting from a discrete set of modes labeled by $\alpha$,
$\Sigma_{R}(\omega)=\sum_{\alpha}\frac{g_{\alpha}^{2}}{\lambda_{\alpha}-i\omega},$ (S48)
the continuum limit gives
$\Sigma_{R}(\omega)$ $=\int d\lambda\;\rho(\lambda)\,\frac{\overline{g^{2}}(\lambda)}{\lambda-i\omega}\equiv\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega},$
$\qquad\mathcal{J}(\lambda)\equiv\rho(\lambda)\,\overline{g^{2}}(\lambda),$ (S49)
where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates (e.g. from Jacobian eigenvalues). In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
$\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S50)
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective variable.
(VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S51)
Coupling $\phi$ locally to an OU reservoir $X_{\lambda}(\mathbf{r},t)$ and integrating out $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the nonlocal-in-time term
$S_{\rm eff}\supset-\int dt\,dt^{\prime}\,d^{d}r\;\tilde{\phi}(\mathbf{r},t)\,K(t-t^{\prime})\,\phi(\mathbf{r},t^{\prime}),$ (S52)
with $K(\omega)=\int d\lambda\,g(\lambda)^{2}/(\lambda-i\omega)\equiv\Sigma_{R}(\omega)$. Hence,
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega)$ (S53)
(up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response).
(IX) Summary of the exact Gaussian result: Starting from the coupled Langevin system (S1)–(S2), the MSRJD action (S8)–(S9) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S37) with a causal memory kernel (S38). In frequency space the kernel takes the spectral form (S43), which is naturally identified with the retarded self-energy (S45). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
Appendix B Appendix B: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$.
(I) Coupled Markovian dynamics: collective field plus relaxational reservoir: For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S54)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\zeta_{\lambda}(t).$ (S55)
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. Equations (S54)–(S55) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
(II) Exact solution for the OU reservoir modes: Equation (S55) is linear and can be solved exactly by the integrating-factor method. Multiply (S55) by $e^{\lambda t}$:
$e^{\lambda t}\dot{X}_{\lambda}(t)+\lambda e^{\lambda t}X_{\lambda}(t)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S56)
The left-hand side is the total derivative
$\frac{d}{dt}\Big(e^{\lambda t}X_{\lambda}(t)\Big)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S57)
Integrate from an initial time $t_{0}$ to $t$:
$e^{\lambda t}X_{\lambda}(t)-e^{\lambda t_{0}}X_{\lambda}(t_{0})=\int_{t_{0}}^{t}ds\,e^{\lambda s}\big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\big].$ (S58)
Solving for $X_{\lambda}(t)$ yields the exact causal representation
$X_{\lambda}(t)$ (S59)
$\quad=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
This expression already shows that the reservoir coordinate at time $t$ depends on the entire history $\{\phi(s)\}_{s<t}$.
(III) Substitution into the collective equation and emergence of a memory kernel: Substituting (S59) into the collective equation (S54) gives
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,$
$\Bigg[e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big(g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big)\Bigg]$
$+\eta(t)$
$=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s)+\eta(t).$ (S60)
We now isolate the term proportional to the history $\phi(s)$:
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\quad=\int_{t_{0}}^{t}ds\,\Bigg[\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}e^{-\lambda(t-s)}\Bigg]\phi(s).$ (S61)
This motivates the definition of the (causal) memory kernel
$K(t-s)\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},\qquad t\geq s,$ (S62)
so that (S61) becomes simply
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S63)
With this definition, Eq. (S60) can be written as
$\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S64)
where we have collected the remaining contributions into an effective noise term (generated by the reservoir noise) and an initial-condition transient:
$\xi_{\rm eff}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s),$ (S65)
$J_{\rm ic}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0}).$ (S66)
Equation (S64) is an exact generalized Langevin equation for $\phi(t)$. Although the enlarged system (S54)–(S55) is Markovian, eliminating the hidden reservoir coordinates produces a non-Markovian (memory) term and, generally, colored effective noise.
Long-time limit.
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure: Defining the time-scale density of states as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (S67)
(Here $\rho(\lambda)$ should be understood as an effective spectral weight for the coupling of relaxation rates to $\phi$; in a microscopic discrete-mode representation one naturally has $\rho(\lambda)\overline{g^{2}}(\lambda)$, and we absorb any slowly varying $\overline{g^{2}}(\lambda)$ into $\rho(\lambda)$ for notational simplicity.)
the memory kernel (S62) becomes
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S68)
Thus $K(t)$ is precisely the Laplace transform of the TDOS (with cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object, and the time-domain memory kernel is its direct transform.
(V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (S69)
then for times $t\gg\Lambda^{-1}$ the infrared part of (S68) dominates and we may approximate $\rho(\lambda)\simeq\rho_{0}$ over the relevant range. In that case the kernel can be evaluated exactly:
$K(t)$ $=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}=\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}.$ (S70)
Perform the change of variables $u=\lambda t$ (so $d\lambda=du/t$):
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\int_{0}^{\Lambda t}du\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S71)
Therefore, at long times $t\gg\Lambda^{-1}$,
$K(t)\simeq\frac{\rho_{0}}{t},\qquad(t\gg\Lambda^{-1}),$ (S72)
which is the universal long-memory tail discussed in the main text.
(VI) Frequency-space form and connection to the retarded self-energy: For completeness we relate the time-domain memory kernel to the standard frequency-space self-energy. Taking the Fourier transform of the reservoir solution (S59) in the stationary (long-time) limit yields
$X_{\lambda}(\omega)=\frac{g(\lambda)}{\lambda-i\omega}\,\phi(\omega)+\frac{1}{\lambda-i\omega}\,\zeta_{\lambda}(\omega),$ (S73)
and substituting into (S54) gives
$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S74)
with the retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S75)
(We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
Separating real and imaginary parts using
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\,\frac{\omega}{\lambda^{2}+\omega^{2}},$ (S76)
one obtains
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S77)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S78)
For a flat TDOS $\rho(\lambda)\simeq\rho_{0}$ and $|\omega|\ll\Lambda$,
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (S79)
and
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
$=\frac{\rho_{0}}{2}\Big(\ln(\Lambda^{2}+\omega^{2})-\ln(\omega^{2})\Big)$
$=\frac{\rho_{0}}{2}\ln\!\left(\frac{\Lambda^{2}+\omega^{2}}{\omega^{2}}\right)\simeq\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$
$\qquad(|\omega|\ll\Lambda),$ (S80)
which reproduces the logarithmic reactive singularity discussed in the main text.
Summary. Equations (S64)–(S68) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
(VII) Conceptual structure: from hidden modes to infrared memory: The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states (TDOS),
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)e^{-\lambda t}.$ (S81)
This relation is purely kinematic and does not rely on perturbation theory, weak coupling, or specific microscopic assumptions. The infrared dynamical structure of the collective field is therefore completely determined by the low-$\lambda$ scaling of $\rho(\lambda)$.
If the TDOS vanishes at the origin, $\rho(\lambda)\to 0$ as $\lambda\to 0$, the kernel decays exponentially and the dynamics reduces to effectively Markovian overdamping. By contrast, when the TDOS remains finite at zero relaxation rate,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (S82)
the Laplace transform produces the universal long-time tail
$K(t)\simeq\frac{\rho_{0}}{t},\qquad t\gg\Lambda^{-1}.$ (S83)
Thus long-memory infrared dynamics is not an additional assumption but an unavoidable consequence of a flat slow-mode spectrum. The TDOS provides the fundamental organizing principle linking microscopic stability structure to macroscopic temporal response.
In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
Appendix C Appendix C: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel. We contrast the conventional Eliashberg regime with the memory-dominated regime generated by a flat relaxation-rate density of states. The derivation follows from shell integration of the linearized gap equation and clarifies the origin of marginal versus relevant superconducting instabilities.
Fermionic Matsubara frequencies are denoted by $\omega$ and $\omega^{\prime}$, and $\Omega\equiv\omega-\omega^{\prime}$ is the bosonic transfer frequency entering the pairing kernel $g(\Omega)$.
(I) Linearized gap equation and RG shell construction: At low energies the linearized gap equation takes the schematic form
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (S84)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff. The factor $1/|\omega^{\prime}|$ originates from the fermionic Cooper bubble. The effective pairing kernel $g(\Omega)$ is generated by collective fluctuations.
The RG procedure consists of integrating out a thin frequency shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$ with $b>1$ and absorbing its contribution into a renormalized coupling $g\rightarrow g+dg$. Defining the RG scale $l=\ln b$, the flow equation follows from the shell contribution.
(II) Conventional Eliashberg regime: Near metallic quantum criticality with Ohmic Landau damping, the bosonic propagator obeys
$D(\Omega)\sim\frac{1}{|\Omega|},$ (S85)
so that the pairing kernel scales as
$g(\Omega)\sim g\,D(\Omega)\sim\frac{g}{|\Omega|},$ (S86)
where $g$ denotes the dimensionless pairing interaction.
The leading shell correction arises from the ladder diagram and is proportional to
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}.$ (S87)
The shell integral is elementary:
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln\Lambda-\ln(\Lambda/b)=\ln b\equiv dl.$ (S88)
Hence the RG equation takes the marginal form
$\frac{dg}{dl}=C\,g^{2},$ (S89)
with $C$ a positive numerical constant.
The solution is
$g(l)=\frac{g_{0}}{1-Cg_{0}l}\simeq g_{0}+Cg_{0}^{2}l,$ (S90)
demonstrating logarithmic enhancement characteristic of a marginal superconducting instability.
(III) Memory-dominated regime: In the memory-dominated regime derived in Sec. II, the collective susceptibility behaves as
$\chi(\Omega)\sim\frac{1}{\rho_{0}\ln(\Lambda/|\Omega|)},$ (S91)
reflecting the flat TDOS of slow relaxation modes.
The pairing kernel therefore scales as
$g(\Omega)\sim\frac{g}{\ln(\Lambda/|\Omega|)}.$ (S92)
The shell correction becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\frac{1}{\ln(\Lambda/\omega^{\prime})}.$ (S93)
Introducing the variable
$u=\ln\frac{\Lambda}{\omega^{\prime}},\qquad du=-\frac{d\omega^{\prime}}{\omega^{\prime}},$ (S94)
the integration limits transform as
$\omega^{\prime}=\Lambda\Rightarrow u=0,\qquad\omega^{\prime}=\Lambda/b\Rightarrow u=\ln b\equiv dl.$ (S95)
Equation (S93) becomes
$dg\propto g^{2}\int_{0}^{dl}\frac{du}{u}.$ (S96)
The apparent divergence at $u\to 0$ is cut off at a microscopic scale $u_{\min}=O(1)$ marking the onset of the logarithmic TDOS regime. The shell contribution reduces to
$dg=\alpha\,g\,dl,$ (S97)
where $\alpha>0$ depends on the spectral weight $\rho_{0}$ and the onset scale of the memory-dominated regime.
The RG equation is therefore
$\frac{dg}{dl}=\alpha\,g,$ (S98)
with solution
$g(l)=g_{0}e^{\alpha l}=g_{0}b^{\alpha}.$ (S99)
The pairing interaction thus grows as a power law under infrared coarse graining, signaling a relevant superconducting instability.
(IV) Physical origin of marginal versus relevant behavior: The qualitative difference between Eqs. (S89) and (S98) originates from the infrared structure of the pairing kernel.
In the Eliashberg regime, each logarithmic shell contributes independently, leading to cumulative logarithmic growth. In contrast, the memory-dominated kernel generated by a flat TDOS correlates fluctuations across all time scales, producing linear amplification of the effective interaction under RG.
Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into a runaway superconducting instability once a slow-mode reservoir is formed.
Summary. Ohmic Landau damping yields a marginal Cooper-channel flow $dg/dl\propto g^{2}$, producing only logarithmic enhancement. A flat relaxation-rate TDOS instead generates a memory-dominated kernel that renders the pairing interaction relevant, with $dg/dl\propto g$ and power-law growth. This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$. This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S100)
[…]
$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S101)
The elementary integral yields
$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S102)
In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S103)
which expresses the nonanalytic infrared enhancement at zero temperature.
(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies. As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S101) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S104)
where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
Evaluating the integral gives
$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S105)
Thus the infrared divergence is universally cut off by the thermal scale,
$\Pi(T)\propto\frac{1}{T}.$ (S106)
(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S107)
Using the standard identity
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S108)
and expanding for $\lambda\ll T$,
$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S109)
one obtains
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S110)
Substitution into Eq. (S107) yields
$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S111)
confirming that finite temperature universally regulates the infrared divergence.
(IV) Implication for the superconducting transition: The Thouless criterion,
$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S112)
therefore gives
$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S113)
or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale. The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime. We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S114)
where
$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S115)
and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S116)
(Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.)
(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S117)
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S118)
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S119)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S120)
(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S121)
For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S122)
Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S120) is maximized when
$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S123)
which yields
$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S124)
For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
(V) Universal scaling collapse: Define the scaling variable
$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S125)
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S120) reduces to the universal form
$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S126)
so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S127)
yielding
$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S128)
For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
(VII) Physical interpretation: Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
Equation (S120) (or equivalently the collapse form (S126)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.
Appendix F Appendix F: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In conventional quantum many-body theory, collective behavior is typically organized in terms of energy eigenmodes of an underlying Hamiltonian,
$H|n\rangle=E_{n}|n\rangle,$ (S129)
with dynamical response expressed through Fourier modes oscillating at frequencies set by energy differences. This structure naturally emphasizes stationary states and conservative dynamics, with dissipation treated as a secondary perturbative effect.
However, this description represents a special limit corresponding to closed, Hamiltonian systems. In general interacting and open systems — particularly near criticality — collective dynamics is governed not by energy eigenmodes but by the stability spectrum of the full dynamical flow.
(I) Dynamical linearization and relaxation modes: For a general high-dimensional dynamical system
$\dot{x}(t)=F(x(t)),$ (S130)
linearization about a metastable configuration $x^{*}$ yields
$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S131)
The eigenvalue problem
$Av_{i}=\lambda_{i}v_{i}$ (S132)
defines a spectrum of collective relaxation modes with temporal evolution
$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S133)
The real parts of $\lambda_{i}$ set intrinsic time scales of decay, while imaginary parts correspond to oscillatory components. Energy eigenmodes arise only in the special Hamiltonian limit where all eigenvalues lie purely on the imaginary axis.
Thus, the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural and generally applicable dynamical basis for collective phenomena.
(II) Liouvillian formulation and relation to energy spectra: In quantum dynamical field theories formulated in Keldysh or MSRJD language, evolution of the density matrix is generated by a Liouvillian superoperator $\mathcal{L}$,
$\frac{d\rho}{dt}=\mathcal{L}\rho.$ (S134)
For closed Hamiltonian dynamics,
$\mathcal{L}\rho=-i[H,\rho],$ (S135)
yielding eigenvalues
$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S136)
which lie purely on the imaginary axis and reproduce energy-difference oscillations.
In interacting open systems and critical regimes, additional dissipative and collective processes shift Liouvillian eigenvalues into the complex plane, producing a dense spectrum of decay rates. The fundamental objects governing long-time dynamics are therefore relaxation rates rather than energy levels.
(III) Memory kernels and continuous relaxation spectra: General linear causal dynamics with memory takes the form
$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S137)
Taking the Laplace transform gives
$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S138)
The poles and branch cuts of $s+\tilde{K}(s)$ define the intrinsic relaxation spectrum. When $\tilde{K}(s)$ possesses branch-cut structure, the response is governed by a continuous distribution of decay rates,
$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S139)
which uniquely defines the time-scale density of states.
Thus continuous relaxation spectra emerge generically from causality and memory, independent of microscopic details.
(IV) Energy modes as a special case: The familiar Fourier representation of conservative dynamics corresponds to the singular limit in which the TDOS collapses onto the imaginary axis,
$\rho(\lambda)\propto\delta(\mathrm{Re}\,\lambda),$ (S140)
recovering purely oscillatory energy eigenmodes. Away from this limit, generic interacting systems naturally develop broad relaxation spectra controlling infrared behavior.
(V) Physical implication: The time-scale density of states therefore represents the fundamental organizing variable of collective dynamics in correlated systems. Energy eigenmodes describe a special integrable limit, while relaxation-rate spectra capture the universal structure of open, critical, and strongly interacting matter.
Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes. The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more general foundation than energy-based descriptions.

### v2

@@
-Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity
+arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026
 Abstract
-Understanding the dynamical origin of high–temperature superconductivity remains a central challenge of strongly correlated quantum matter.
-Conventional approaches to quantum criticality assume overdamped Markovian dissipation governed by Ohmic Landau damping.
-Here we show that infrared collective dynamics is instead generically controlled by the time–scale density of states (TDOS) of relaxation modes.
-Within the Martin–Siggia–Rose–Janssen–De Dominicis formalism, we derive an exact spectral representation of the collective susceptibility in terms of the TDOS.
-A finite infrared TDOS defines a new universality class of memory–dominated critical dynamics characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response.
-This spectral reorganization produces a strong infrared amplification of the particle–particle channel, converting the marginal instabilities of BCS and Eliashberg theories into algebraic superconducting transitions.
-As a result, the transition temperature scales linearly with the infrared spectral weight of slow collective modes, naturally yielding superconducting domes and Uemura scaling without invoking bosonic glue or fine tuning.
-The same slow–mode reservoir governs anomalous normal–state dynamics, including long–time correlations and strange–metal behavior, providing a unified description of thermodynamic and dynamical phenomena in correlated superconductors.
-Our results establish dynamical spectral organization as a fundamental principle of quantum critical matter and identify memory–dominated criticality as a generic route to high–temperature superconductivity.
+[…]
 I I. Introduction
@@
 A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11].
-Such low-dimensional reductions necessarily generate a finite set of soft modes and therefore preclude the emergence of a dense continuum of relaxation time scales.
-A genuine slow-mode spectrum can only arise when the full high-dimensional collective dynamical state space is retained.
-In this work we show that infrared collective dynamics is generically controlled not by a single overdamped mode but by the time–scale density of states (TDOS) of correlated internal relaxation modes.
-By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [12, 13, 14], we derive an exact spectral representation of the collective susceptibility in terms of the TDOS.
+Such descriptions implicitly assume that the infrared behavior can be captured by a finite set of soft bosonic modes with local-in-time damping.
+However, strongly correlated quantum matter is fundamentally governed by Liouvillian evolution of the density matrix, where the full spectrum of relaxation eigenvalues determines long-time dynamics [12, 13, 14].
+In more general, criticality need not correspond to softening of a single excitation mode but may instead involve reorganization of the entire relaxation-rate spectrum.
+Rather than reducing the dynamics to a few collective coordinates, one must consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes.
+Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality.
+The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
+In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels.
+By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
 We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics.
 While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$.
-This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a new universality class governed by self–organized slow modes rather than fast external baths.
-Crucially, we show that such flat relaxation mode spectra arise naturally in strongly correlated collective systems, including self–organized dynamical manifolds explored in our previous work [15, 16, 17].
-The resulting long–memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a natural route to strong superconductivity without invoking finely tuned bosonic mediators.
+This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a distinct universality class governed by the structure of the relaxation spectrum.
+Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20].
+The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
 Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter.
-The infrared structure of the relaxation mode spectra determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 18, 19].
-Charge, spin, and pairing channels reorganize simultaneously within the same slow–mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a common infrared spectral continuum.
-Our results establish the relaxation mode spectra as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity.
-This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified perspective on collective behavior across strongly correlated quantum matter [5, 7, 10].
+The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22].
+Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
+Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity.
+This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 10].
 The remainder of this paper is organized as follows.
@@
 II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
-Strongly correlated electronic matter hosts many intertwined collective fluctuation channels, including charge, spin, current, pairing, and phase–coherence sectors.
-These degrees of freedom continuously exchange energy and information and cannot, in general, be reduced to a small set of independent order parameters.
-To capture this collective complexity in a unified manner, we represent the correlated system as a high–dimensional interacting dynamical system whose instantaneous collective configuration is encoded in a state vector
-$x(t)\in\mathbb{R}^{N},$ (1)
-which evolves under nonlinear many–body feedback.
-The coarse-grained dynamics may therefore be written as a stochastic field theory
-$\dot{x}(t)=F(x(t))+\xi(t),$ (2)
-where $F$ encodes correlated interactions and $\xi$ is Gaussian noise with
-$\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=2D\,\delta_{ij}\delta(t-t^{\prime}).$ (3)
+Strongly correlated electronic matter is intrinsically a quantum many-body system.
+At long times and coarse-grained scales, its dynamics is governed not by isolated quasiparticles, but by collective relaxation channels generated by the full quantum dynamical evolution of the density matrix [12, 13, 14],
+$\frac{d\hat{\varrho}}{dt}=\mathcal{L}\hat{\varrho}$, (1)
+where $\mathcal{L}$ denotes the Liouvillian generator of the effective (open or coarse-grained) dynamics.
+The eigenmodes of $\mathcal{L}$,
+$\mathcal{L}R_{n}=\lambda_{n}R_{n},$ (2)
+define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
+$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
+Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
+Observable collective responses—charge, spin, pairing, current, and phase-coherence fluctuations—arise from projections of the full relaxation spectrum onto experimentally accessible operators.
+The resulting low-energy theory is thus controlled by the statistical structure of relaxation rates.
+We define the TDOS as the infrared distribution of Liouvillian relaxation eigenvalues,
+$\rho(\lambda)=\sum_{n}\delta\!\left(\lambda+\mathrm{Re}\,\lambda_{n}\right),$ (4)
+or its continuum limit.
+The spectral structure introduced here should therefore be understood at the level of the full dynamical generator rather than solely in terms of Hamiltonian energy eigenmodes.
+Energy spectra characterize the conservative limit in which all Liouvillian eigenvalues lie on the imaginary axis, whereas generic interacting or coarse-grained systems develop complex spectra with nonzero decay rates.
+A detailed discussion of this relaxation-rate (Liouvillian) formulation and its connection to memory kernels is provided in Appendix A.
 II.1 A. Dynamical stability spectrum and time-scale density of states
-Near correlated critical regimes the system does not soften a single collective mode but undergoes a reorganization of its full dynamical stability structure.
-Linearizing around a fluctuating metastable configuration $x^{*}$,
-$\delta\dot{x}(t)=-A\,\delta x(t)+\xi(t),$ (4)
-with
-$A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}},$ (5)
-yields a spectrum of collective relaxation modes
-$Av_{i}=\lambda_{i}v_{i},$ (6)
-whose retarded propagators are
-$G_{i}(\omega)=\frac{1}{-i\omega+\lambda_{i}}.$ (7)
-The infrared structure of correlated dynamics is therefore controlled by the distribution of relaxation rates $\lambda_{i}$. We characterize this spectrum by the time-scale density of states,
-$\rho(\lambda)=\Big\langle\sum_{i}\delta(\lambda-\lambda_{i})\Big\rangle.$ (8)
-In weakly interacting systems the TDOS vanishes as $\lambda\to 0$, reflecting a finite number of quasiparticle-like modes.
-By contrast, strongly correlated critical regimes generically generate an extensive accumulation of near-marginal directions in phase space,
+The infrared structure of correlated matter is encoded in the spectrum of collective relaxation rates.
+Rather than softening a single order parameter, correlated critical regimes reorganize the entire decay spectrum, producing a dense accumulation of near-marginal modes.
+We characterize this structure by introducing the TDOS, which measures the density of collective decay channels at relaxation rate $\lambda$.
+This object plays a role analogous to the energy density of states, but in relaxation-rate space.
+It determines how many collective modes decay on a given time scale.
+In weakly interacting systems, the number of slow modes remains finite, and the TDOS vanishes as $\lambda\to 0$, reflecting the presence of well-defined quasiparticles and isolated collective excitations.
+By contrast, in strongly correlated critical regimes, the relaxation spectrum generically develops an extensive accumulation of near-marginal decay channels,
 $\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (5)
-forming a slow-mode reservoir that governs infrared collective dynamics.
-This accumulation is the dynamical signature of correlated criticality: the system approaches marginal stability along many collective directions rather than along a single order-parameter coordinate.
-II.2 B. Exact spectral representation from MSRJD field theory
-To compute collective response functions while retaining the full dynamical spectrum, we employ the Martin–Siggia–Rose–Janssen–De Dominicis formalism associated with Eq. (2),
-$S[x,\tilde{x}]=\int dt\,\tilde{x}(t)\big[\dot{x}(t)-F(x(t))\big]-D\tilde{x}^{2}(t).$ (10)
-We focus on a slow collective field $\phi(\mathbf{r},t)$ obtained by coarse graining the full state vector and regard the remaining fast correlated degrees of freedom as internal relaxation modes.
-The reorganized stability spectrum may be represented exactly as a continuum of Ornstein–Uhlenbeck modes $X_{\lambda}$ labeled by their decay rates,
+forming a slow-mode reservoir that dominates infrared collective dynamics.
+This accumulation signifies a dynamical approach to marginal stability along a macroscopic manifold of collective directions, rather than along a single order-parameter coordinate.
+II.2 B. Functional representation of the relaxation spectrum
+To compute collective response functions, we employ a functional representation of the dissipative dynamics.
+The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24].
+Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
+Let $\phi(\mathbf{r},t)$ denote a coarse–grained collective bosonic field representing charge, spin, pairing, or phase fluctuations.
+Projecting the full quantum dynamics onto this operator sector yields an exact decomposition in terms of relaxation eigenmodes.
+Because the dissipative spectrum consists of decay rates $\lambda\geq 0$, the projected dynamics may be represented as a continuum of relaxational modes labeled by their decay scales,
 $\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (6)
-$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda}.$ (12)
-Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action becomes
+$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
+where $X_{\lambda}$ represents a decay mode with relaxation rate $\lambda$, and $g(\lambda)$ encodes its overlap with $\phi$.
+Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action reads
 $S=$ $\int dt\,d^{d}r\,\tilde{\phi}\Big[\dot{\phi}+(r+c\nabla^{2})\phi-\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}\Big]-D_{\phi}\tilde{\phi}^{2}$
-$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi]$
+$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}\big[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi\big]$
 $-\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,D_{\lambda}\tilde{X}_{\lambda}^{2}.$ (8)
-Since the reservoir variables enter quadratically, they can be integrated out exactly. In Fourier space this yields the effective collective action
-$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots,$ (14)
-with retarded self-energy
-$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (15)
+Because the reservoir fields $X_{\lambda}$ enter quadratically, they can be integrated out exactly.
+Performing the Gaussian functional integration generates a nonlocal-in-time memory kernel in the effective equation of motion,
+$\partial_{t}\phi(t)=-(r+c\nabla^{2})\phi(t)-\int_{0}^{t}dt^{\prime}\,K(t-t^{\prime})\,\phi(t^{\prime})+\eta_{\rm eff}(t),$ (9)
+where the memory kernel is given exactly by
+$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (10)
+with the time-scale density of states
+$\rho(\lambda)\equiv g(\lambda)^{2}.$ (11)
+Equation (10) makes explicit that the slow-mode reservoir is entirely controlled by the distribution of relaxation rates.
+The long-time structure of $K(t)$ therefore encodes the infrared organization of decay modes in the interacting system.
+Fourier transforming the kernel yields the retarded self-energy,
+$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
+so that the effective action becomes
+$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
 Separating real and imaginary parts,
-$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\frac{\omega}{\lambda^{2}+\omega^{2}},$ (16)
-gives
-$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (17)
-$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{1}{\lambda^{2}+\omega^{2}}.$ (18)
-Identifying the TDOS as
-$\rho(\lambda)\equiv g(\lambda)^{2},$ (19)
-the collective inverse susceptibility becomes
+$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
+$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
+The collective inverse susceptibility therefore assumes the exact spectral representation
 $\chi_{R}^{-1}(\omega,\mathbf{q})$ $=r+cq^{2}+\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}$ (16)
 $\quad+i\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$
-Equation (20) constitutes an exact spectral representation of collective infrared dynamics in terms of the dynamical stability spectrum.
-The complete field-theoretic derivation of the collective susceptibility kernel from the MSRJD formalism, including the explicit Gaussian integration over the relaxation-mode reservoir and the resulting self-energy structure, is presented in Appendix A.
+Equation (16) provides an exact infrared response fully controlled by the relaxation-rate density.
+The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
+The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
 II.3 C. Infrared universality classes
@@
 $\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (18)
-Here $\omega$ denotes the frequency conjugate to the collective relaxation dynamics obtained after integrating out the slow–mode reservoir in the MSRJD formulation, rather than a microscopic quasiparticle energy scale.
+Here $\omega$ denotes the collective relaxation frequency arising from the projected Liouvillian dynamics, rather than a microscopic quasiparticle excitation energy.
 It is therefore natural to write the long–wavelength dynamical response in the scaling form
 $\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (19)
-in direct analogy with the conventional Hertz–Millis representation. The TDOS scaling immediately implies the universal identification $s=\alpha$.
+in direct analogy with the conventional Hertz–Millis representation [3, 4]. The TDOS scaling immediately implies the universal identification $s=\alpha$.
 Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$.
@@
 $z=\frac{2}{s}.$ (20)
-In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the presence of an extensive reservoir of slow collective relaxation modes.
+In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the organization of the underlying relaxation spectrum.
 Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$.
-Infrared universality classes are thus classified directly by the low–energy organization of relaxation modes encoded in the TDOS.
+Infrared universality classes are thus classified directly by the low–energy structure of the relaxation-rate density.
 In particular, when the TDOS remains finite at the origin,
@@
 $\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (24)
-corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, which is derived explicitly in Appendix B.
+corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
 This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory.
-Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow internal modes.
-The TDOS thus provides a unified organizing principle for infrared critical dynamics, with different universality classes emerging from the scaling of the dynamical stability spectrum itself.
+Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow decay modes.
 III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
-The emergence of a slow–mode reservoir qualitatively reorganizes the infrared structure of the Cooper pairing channel.
+The emergence of a slow–mode reservoir reorganizes the infrared structure of the Cooper pairing channel.
 In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal.
@@
 The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining.
-Physically, overdamped collective modes rapidly dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
+Physically, overdamped collective modes dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
 III.2 B. Infrared reorganization by slow collective modes
@@
 II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$.
-After integrating out the relaxation reservoir within the MSRJD framework, the retarded collective susceptibility acquires the exact spectral form
+After integrating out the relaxation reservoir, the retarded collective susceptibility takes the exact spectral form
 $\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (29)
@@
 Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct.
-In Eliashberg theory the singularity reflects instantaneous overdamped dissipation arising from coupling to a fast electronic bath [20, 21].
-Here, by contrast, the same infrared scaling is a consequence of self–organized spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
+In Eliashberg theory the singularity reflects overdamped dissipation arising from coupling to a fast electronic bath [25, 26].
+Here, by contrast, the same infrared scaling results from spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
 This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales.
@@
 2.
-A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix C.
-Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and dynamically amplify intrinsic electronic pairing tendencies.
-Superconductivity in strongly correlated systems thus emerges from infrared spectral self–organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
+A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix D.
+Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and amplify intrinsic electronic pairing tendencies.
+Superconductivity in strongly correlated systems thus emerges from infrared spectral organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
 Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction.
@@
 Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
-In this section we show explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, yielding power–law growth of the effective pairing interaction and algebraic scaling of the transition temperature.
+In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
 IV.1 A. Intrinsic electronic pairing and dynamical formulation
@@
 favoring local singlet formation.
-More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated by electronic interactions, independent of long–wavelength collective mediators.
+More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated purely by electronic interactions, independent of long–wavelength collective mediators.
 This intrinsic scale provides the microscopic seed for superconductivity.
@@
 $\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (36)
-with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies this preexisting pairing tendency.
+with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
 IV.2 B. Memory–dominated Cooper kernel from flat TDOS
-Because single–particle propagators admit a relaxational spectral representation controlled by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
+Because single–particle propagators admit a relaxational spectral representation governed by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
 $\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (37)
@@
 $\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (39)
-This nonanalytic enhancement reflects the accumulation of collective memory across all time scales.
-Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel is dominated by long–time memory effects generated by the slow–mode TDOS.
-IV.3 C. Algebraic superconducting instability and universality
+This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales.
+Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel becomes relevant due to the flat TDOS.
+The slow–mode reservoir therefore reorganizes the pairing problem into a memory–dominated critical regime.
+IV.3 C. Algebraic Instability and Transition Scale
 The superconducting transition follows from the Thouless criterion
@@
 $T_{c}\sim\rho_{0}\,E_{\rm pair},$ (41)
-up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
+up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix E.
 The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir.
 High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
-Because the TDOS exponent controls infrared dynamics independently of microscopic details, the pairing amplification mechanism is universal across correlated systems exhibiting flat relaxation spectra.
-The slow–mode reservoir therefore acts as a universal dynamical amplifier, converting electronic correlation energy into enhanced superconducting transition temperatures without invoking specific bosonic glue.
+It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence.
+The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
+$k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right),$ (42)
+with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
+The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$.
+The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
+In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios.
+The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
 Figure 3: Experimental consequences of memory–dominated critical dynamics.
@@
 The linearized transition scale derived in Eq.
-(45) shows that the superconducting temperature is controlled by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir.
-Near correlated quantum criticality, $\rho_{0}$ is not constant but is governed by universal scaling.
-Let $p$ denote a tuning parameter (e.g., doping or pressure) and $p_{c}$ the critical point. The mass term of the collective kernel obeys
-$r(p)\sim|p-p_{c}|,$ (46)
-so that the corresponding infrared cutoff scale takes the universal form
-$\Lambda(p)\sim r^{z\nu}\sim|p-p_{c}|^{z\nu},$ (47)
-with $z$ the dynamical exponent and $\nu$ the correlation-length exponent.
-Away from criticality the slow–mode reservoir is therefore truncated in the infrared, yielding an effective low-energy TDOS weight
-$\rho_{0}(p)\sim\int_{0}^{\Lambda(p)}d\lambda\,\rho(\lambda).$ (48)
-For a memory–dominated regime in which the TDOS remains finite at the origin, $\rho(\lambda\to 0)=\rho_{*}$, one obtains parametrically
-$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{p_{0}}\right),$ (49)
-where $f$ is a universal decreasing scaling function and $p_{0}$ denotes a nonuniversal crossover scale.
-Substituting Eq. (49) into Eq. (45) gives
-$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{x_{0}}\right),$ (50)
-which is maximal near criticality and suppressed on both sides, producing a superconducting dome without fine tuning.
-The dome structure is thus a direct consequence of the competition between intrinsic short–range electronic pairing and the finite infrared extent of the slow–mode reservoir away from criticality.
-Importantly, no competing order parameters or additional microscopic suppression mechanisms are required: suppression of $T_{c}$ arises purely from the truncation of long–time memory encoded in the TDOS.
-In this interpretation, optimal doping corresponds to maximal infrared spectral reorganization rather than to the peak of any specific bosonic fluctuation.
+(41) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir.
+We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
+In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
+$\rho(\lambda\to 0)=\rho_{*}>0,$ (43)
+without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
+External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir.
+Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal.
+The collective mass term behaves as
+$r(p)\sim|p-p_{\star}|,$ (44)
+so that the associated cutoff obeys
+$\Lambda(p)\sim r^{z\nu}\sim|p-p_{\star}|^{z\nu}.$ (45)
+Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
+$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (46)
+with $f$ a decreasing crossover function.
+Substituting into Eq. (41) gives
+$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (47)
+which naturally produces a superconducting dome.
+Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
+In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point.
+Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
 V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
-While the microscopic pairing seed is set by the intrinsic electronic scale $E_{\rm pair}$, the establishment of superconductivity requires global phase coherence of the Cooper collective field.
-In systems with low superfluid density, the transition may be controlled by phase stiffness rather than by gap formation.
-The effective phase–only action for the Cooper field reads
+While Sec.
+IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence.
+In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
+The long-wavelength phase dynamics of the Cooper field is described by the effective action
 $S_{\theta}=\frac{\rho_{s}}{2}\int d\tau d^{d}r\left[(\partial_{\tau}\theta)^{2}+v_{s}^{2}(\nabla\theta)^{2}\right],$ (48)
-where $\rho_{s}$ is the superfluid stiffness. The phase–ordering temperature therefore scales as
+where $\rho_{s}$ denotes the superfluid stiffness. The corresponding phase-ordering temperature scales as
 $T_{\theta}\sim\rho_{s}.$ (49)
-Within the slow–mode reservoir framework, the stiffness is controlled by the same infrared spectral weight that enhances pairing. Integrating out amplitude fluctuations yields the scaling
+Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport.
+The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
+In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$.
+Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$ [Eq.
+(39)], the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight.
+Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
 $\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (50)
-reflecting the collective participation of slow modes in coherent transport. Combining Eqs. (52) and (53) gives
+up to nonuniversal numerical factors.
+Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
+Combining Eqs. (49) and (50) gives
 $T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (51)
 establishing a direct proportionality between the superconducting transition temperature and superfluid density.
-Equation (54) reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors.
-In contrast to BCS superconductors, where phase stiffness is large and irrelevant to $T_{c}$, correlated systems operate in a regime where infrared collective dynamics simultaneously controls pairing and coherence.
-In the present framework, this is not an additional assumption but follows directly from the single infrared control parameter $\rho_{0}$.
+This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density.
+In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
+Within the present framework, this proportionality is not imposed phenomenologically.
+It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness.
+The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
 V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
@@
 corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
-Such long–time tails naturally generate 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors.
+Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors.
 Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure.
@@
 We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
-The TDOS controls: (i) the extent of the slow–mode reservoir away from criticality [Eq.
-(47)], (ii) the dynamical amplification of intrinsic pairing tendencies [Eq.
-(50)], (iii) the superfluid stiffness and phase coherence [Eq.
-(54)], and (iv) the emergence of non-Markovian long–time response.
-Quantum criticality thus acts not merely as a source of fluctuations but as a self–organized dynamical amplifier that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
-This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) should simultaneously display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling.
-The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
+The TDOS controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response.
+Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
+This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling.
+The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
 VI VI. Discussion
-The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is governed not by local Markovian dissipation but by the time–scale density of states of collective relaxation modes.
-Criticality corresponds to a reorganization of the entire dynamical spectrum, and universality classes are naturally classified by the low–energy scaling of the TDOS.
-In particular, a finite TDOS at vanishing relaxation rate defines a memory–dominated infrared fixed point characterized by long–time kernels $K(t)\sim 1/t$.
-This perspective clarifies the limited domain of validity of conventional coarse–grained field theories.
-Away from criticality, relaxation spectra are gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective order parameters.
-Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending to zero.
-Internal modes cannot be integrated out: the full relaxation spectrum remains dynamically active at all time scales.
-The breakdown of Markovian coarse–graining is therefore a universal consequence of dynamical criticality rather than a model-dependent artifact.
-This observation explains why Hertz–Millis–type approaches, which assume overdamped local dynamics of a small set of collective fields, cannot capture the infrared structure of strongly correlated critical matter.
-Criticality does not correspond to the softening of a single bosonic mode but to a spectral condensation of near–marginal collective degrees of freedom.
-The TDOS thus emerges as the natural organizing variable of infrared dynamics, replacing phenomenological damping parameters and providing a microscopic foundation for anomalous relaxation and transport.
-A central physical consequence of this spectral reorganization is the qualitative enhancement of superconducting instabilities.
-Rather than requiring finely tuned bosonic mediators, superconductivity emerges generically from dynamical amplification of intrinsic electronic pairing tendencies by the slow–mode reservoir.
+The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by softening of a single order parameter.
+The time–scale density of states provides a quantitative measure of this organization.
+When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response.
+This viewpoint clarifies the limits of conventional coarse-grained field theories.
+Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields.
+Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero.
+In this regime, internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales.
+The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
+This observation highlights the structural difference from Hertz–Millis–type approaches.
+Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping.
+In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels.
+Criticality is reinterpreted not as single-mode softening but as reorganization of the entire decay spectrum.
+The TDOS thus replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
+A direct physical consequence of this spectral reorganization is qualitative enhancement of superconducting instabilities.
+Superconductivity need not rely on finely tuned bosonic mediators.
+Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir.
 The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
-A single infrared parameter, $\rho_{0}$, governs thermodynamic and dynamical phenomena alike.
-The superconducting dome reflects truncation of long–time memory away from criticality, Uemura scaling follows from the shared spectral origin of pairing and phase coherence, and anomalous normal-state dynamics arises from the same non-Markovian kernel.
-Quantum criticality therefore acts as a self-organized dynamical amplifier that simultaneously controls fluctuations, transport, and superconducting order.
-Although electrons propagate as Bloch quasiparticles on a crystalline lattice, collective particle–hole and particle–particle excitations form extended many-body coordinates not constrained to lattice energy eigenstates.
-In strongly correlated regimes these collective modes are generically overdamped by the electronic continuum, rendering a relaxation-rate description more fundamental than an energy-eigenmode picture.
-While conventional lattice theories organize collective behavior in terms of Hamiltonian eigenmodes, the present framework shows that near criticality the relevant organizing structure is instead the dynamically self-organized spectrum of relaxation rates.
-This relaxation–rate organization is not a phenomenological assumption but a direct consequence of the fundamental structure of nonequilibrium quantum and classical dynamics.
-As shown in the Appendix F, both Liouvillian evolution in open quantum systems and generalized memory–kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues rather than by Hamiltonian energy eigenstates.
-Even in the absence of spatial degrees of freedom, the Laplace–domain representation of non-Markovian dynamics generically yields continuous spectra of relaxation rates whenever long–time memory kernels are present.
-The emergence of a dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, not a model-dependent artifact of coarse–graining.
-From this perspective, energy eigenmodes describe conservative reversible dynamics, while relaxation spectra encode the physically relevant irreversible organization governing collective behavior near criticality.
-Strongly correlated quantum matter near dynamical critical points is therefore more naturally classified by its relaxation-rate spectrum than by its Hamiltonian excitation spectrum.
-The TDOS framework introduced here thus elevates relaxation modes to the status of fundamental collective coordinates, providing a principled dynamical basis for infrared universality, memory effects, and emergent macroscopic coherence.
-Away from criticality the dynamical spectrum collapses onto a trivial fast mode.
-Near memory-dominated criticality an extensive continuum of slow collective modes emerges and governs macroscopic coherence.
-Superconducting pairing in this regime corresponds to condensation within a slow-mode reservoir rather than binding mediated by a single sharp bosonic excitation.
-This unified picture resolves long-standing puzzles of unconventional superconductivity, including the coexistence of strange-metal behavior with robust pairing and the apparent universality of dome-shaped phase diagrams across disparate materials.
-Rather than invoking competing orders or material-specific bosonic glue, these phenomena arise naturally from universal infrared spectral structure.
-More broadly, the TDOS framework establishes dynamical spectral organization as a fundamental principle of correlated quantum matter.
-Memory-dominated criticality represents a new universality class beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and self-organized critical systems.
-Future directions include quantitative determination of TDOS spectra in microscopic models and experiments, extension to nonequilibrium driven systems, and exploration of engineered platforms in which slow-mode reservoirs can be dynamically controlled.
-In this light, high-temperature superconductivity appears not as an anomaly requiring special microscopic ingredients, but as a generic consequence of infrared spectral self-organization in strongly correlated matter.
+In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
+A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena.
+The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; anomalous normal-state dynamics arises from the same non-Markovian kernel.
+Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
+The relaxation-rate organization emphasized here is not merely a phenomenological assumption.
+Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues.
+Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra.
+A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, rather than a model-specific construction.
+From this perspective, Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality.
+Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
+More broadly, the TDOS framework elevates dynamical spectral organization to a fundamental organizing principle of correlated quantum matter.
+Memory-dominated criticality represents a universality class distinct from traditional dissipative field theories, with potential implications for non-Fermi liquids, glassy systems, and self-organized critical dynamics.
 Memory–dominated dynamics as a guiding principle for room-temperature superconductivity.
-Recent experimental efforts toward room-temperature superconductivity have largely focused on engineering stronger microscopic pairing interactions through pressure, lattice design, or exotic chemical compositions.
+Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition.
 Within the present framework, however, enhancing the bare pairing scale alone is insufficient.
-Our results show that the decisive control parameter for high transition temperatures is not the magnitude of an attractive interaction but the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
-A dense reservoir of near-marginal slow modes fundamentally reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$.
-From this perspective, many existing high-pressure hydride systems and correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
-The present theory therefore suggests a new design principle for superconductivity beyond material-specific glue engineering: the deliberate stabilization of dynamical phases supporting extensive slow-mode continua.
-Controlling electronic structure, frustration, dimensionality, and feedback mechanisms to promote near-marginal collective dynamics may be as crucial as increasing microscopic attraction itself.
-Room-temperature superconductivity, in this view, corresponds to a dynamical phase characterized by maximal infrared memory rather than by extreme coupling strength.
+The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
+A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$.
+From this viewpoint, certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
+The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27].
+Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself.
+Room-temperature superconductivity, in this perspective, corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
 VII VII. Conclusion
 We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter.
-Formulating collective dynamics in terms of the time–scale density of states reveals that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
+By formulating collective dynamics in terms of the time–scale density of states, we show that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
 A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response.
-Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting the marginal instabilities of conventional theories into algebraic superconducting transitions governed by infrared spectral weight.
-This mechanism provides a unified explanation of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state.
-These signatures arise generically from infrared spectral reorganization, without invoking specific bosonic glue, competing orders, or fine tuning.
-More broadly, our results recast quantum criticality as a constructive dynamical process that enhances collective quantum phenomena through self–organized memory effects.
-Memory–dominated criticality thus represents a new paradigm beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
-In this perspective, high–temperature superconductivity emerges not as an exception requiring special microscopic ingredients, but as a natural consequence of infrared spectral self–organization in strongly correlated matter.
-Future exploration of TDOS-controlled dynamics may enable systematic design of materials and devices that exploit memory–dominated criticality to stabilize and enhance collective quantum states.
+Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
+This mechanism provides a unified account of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state.
+These signatures arise naturally from infrared spectral reorganization, without requiring specific bosonic mediators or material-dependent fine tuning.
+More broadly, the results recast quantum criticality as a dynamical process that enhances collective quantum phenomena through extended temporal correlations.
+Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
+In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter.
+Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
 Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
@@
 Supplementary Materials
-Appendix A Appendix A: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
-In this appendix we derive, without omitting algebraic steps, how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism.
+Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
+In conventional quantum many-body theory, collective behavior is often organized in terms of energy eigenmodes of an underlying Hamiltonian,
+$H|n\rangle=E_{n}|n\rangle,$ (S1)
+with dynamical response expressed through oscillatory Fourier components at frequencies set by energy differences.
+This viewpoint is natural for closed systems whose dynamics is unitary and time-reversal invariant, and it emphasizes stationary states and conservative motion.
+However, energy eigenmodes provide only a special organizing basis.
+In generic interacting settings—especially in open, driven, or internally coarse-grained systems—the physically relevant long-time organization is not set by Hamiltonian energies but by the stability spectrum of the effective dynamical generator.
+In this work, the central spectral object is therefore the distribution of relaxation rates governing irreversible decay and memory.
+(I) Dynamical linearization and relaxation modes: For a general high-dimensional (possibly coarse-grained) dynamical system,
+$\dot{x}(t)=F(x(t)),$ (S2)
+linearization about a metastable configuration $x^{*}$ yields
+$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S3)
+The eigenvalue problem
+$Av_{i}=\lambda_{i}v_{i}$ (S4)
+defines a spectrum of collective relaxation modes with temporal evolution
+$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S5)
+The real parts of $\lambda_{i}$ set decay time scales, while imaginary parts encode oscillatory components.
+Purely oscillatory energy-mode dynamics corresponds to the singular limit $\mathrm{Re}\,\lambda_{i}=0$ for all modes, which is realized only in strictly Hamiltonian, non-dissipative evolution.
+Thus the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural dynamical basis for generic collective phenomena, including metastability, critical slowing down, and non-Markovian memory.
+(II) Liouvillian formulation (closed, open, and non-Markovian dynamics): A conceptually complete formulation of nonequilibrium quantum dynamics is given in terms of the density matrix $\hat{\varrho}(t)$ and its dynamical generator.
+For a closed system, $\hat{\varrho}(t)$ evolves by the von Neumann equation,
+$\frac{d\hat{\varrho}}{dt}=-i[H,\hat{\varrho}]\equiv\mathcal{L}_{H}\hat{\varrho}$, (S6)
+where $\mathcal{L}_{H}$ is the Hamiltonian Liouvillian superoperator. If $H|n\rangle=E_{n}|n\rangle$, then the operator basis $|n\rangle\langle m|$ diagonalizes $\mathcal{L}_{H}$:
+$\mathcal{L}_{H}\big(|n\rangle\langle m|\big)=-i(E_{n}-E_{m})\,|n\rangle\langle m|,$ (S7)
+so the Liouvillian eigenvalues are
+$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S8)
+lying purely on the imaginary axis and reproducing undamped oscillations.
+For open quantum systems, the reduced density matrix typically obeys a master equation with a dissipative contribution. In the Markovian limit, the evolution takes Lindblad form,
+[…]
+where the jump operators $L_{\mu}$ encode decoherence and dissipation.
+The resulting Liouvillian $\mathcal{L}$ is generically non-Hermitian as a superoperator and therefore possesses complex eigenvalues
+$\mathcal{L}R_{\alpha}=\lambda_{\alpha}R_{\alpha},\qquad\mathrm{Re}\,\lambda_{\alpha}\leq 0,$ (S10)
+with decay rates set by $\mathrm{Re}\,\lambda_{\alpha}$. The density matrix admits the modal expansion
+$\hat{\varrho}(t)=\hat{\varrho}_{\rm ss}+\sum_{\alpha\neq 0}c_{\alpha}\,e^{\lambda_{\alpha}t}\,R_{\alpha},$ (S11)
+where $\hat{\varrho}_{\rm ss}$ is the stationary state ($\lambda_{0}=0$) and the long-time approach is controlled by the Liouvillian spectral gap and any near-marginal continuum.
+More generally, when the environment or eliminated internal coordinates retain memory, the reduced dynamics becomes non-Markovian and takes a memory-kernel form (Nakajima–Zwanzig structure),
+$\frac{d\hat{\varrho}}{dt}=\int_{0}^{t}ds\;\mathcal{K}(t-s)\,\hat{\varrho}(s),$ (S12)
+where $\mathcal{K}(t)$ is a causal superoperator kernel. In Laplace space,
+$\tilde{\hat{\varrho}}(s)=\big[s-\tilde{\mathcal{K}}(s)\big]^{-1}\hat{\varrho}(0),$ (S13)
+so the poles and branch cuts of $s-\tilde{\mathcal{K}}(s)$ define the intrinsic relaxation spectrum.
+In this sense, even for quantum systems the relevant organizing structure for long-time dynamics is generically a relaxation-rate spectrum (possibly continuous), not an energy spectrum.
+(III) Memory kernels and continuous relaxation spectra: The same spectral logic appears already at the level of linear causal dynamics for a collective coordinate $\phi(t)$:
+$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S14)
+Taking the Laplace transform gives
+$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S15)
+The analytic structure of $s+\tilde{K}(s)$ determines the intrinsic relaxation content: isolated poles correspond to discrete decay modes, while branch cuts encode continua of decay rates.
+Whenever long-time memory is present, the causal kernel admits a spectral representation of the form
+$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S16)
+which defines the time-scale spectrum. A flat infrared TDOS produces the universal tail $K(t)\sim 1/t$.
+Thus continuous relaxation spectra arise generically from causality and memory, independent of microscopic Hamiltonian structure.
+(IV) Energy modes as a special limit: The familiar Fourier-mode picture of conservative dynamics corresponds to the singular limit in which the relevant spectrum collapses onto the imaginary axis.
+In the closed-system Liouvillian case this appears as $\lambda_{nm}=-i(E_{n}-E_{m})$, while in the memory-kernel representation it corresponds to eliminating all decay rates (no weight at $\mathrm{Re}\,\lambda>0$), recovering purely oscillatory behavior.
+Away from this limit, generic interacting and/or open systems naturally develop broad spectra of relaxation rates controlling infrared dynamics.
+(V) Physical implication: The TDOS—the infrared distribution of collective relaxation rates— is therefore the fundamental organizing variable of collective dynamics in correlated systems.
+Energy eigenmodes describe an integrable, conservative limit, whereas relaxation-rate spectra capture the universal structure of irreversible, critical, and strongly interacting matter.
+Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes.
+The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more generally applicable foundation than energy-based descriptions.
+Appendix B Appendix B: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
+In this appendix we derive how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism.
 The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
-(I) Coupled stochastic dynamics: collective variable plus OU reservoir: We consider a collective degree of freedom $x(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$.
+(I) Coupled stochastic dynamics: We consider a collective degree of freedom $\phi(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$.
 The coupled stochastic dynamics is
-$\dot{x}(t)$ $=-r\,x(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S1)
-$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,x(t)+\xi_{\lambda}(t).$ (S2)
+$\dot{\phi}(t)$ $=-r\,\phi(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S17)
+$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\xi_{\lambda}(t).$ (S18)
 The noises are assumed Gaussian, stationary, and white:
-$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{x}\,\delta(t-t^{\prime}),$ (S3)
+$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{\phi}\,\delta(t-t^{\prime}),$ (S19)
 $\langle\xi_{\lambda}(t)\,\xi_{\lambda^{\prime}}(t^{\prime})\rangle$ $=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\,\delta(t-t^{\prime}).$ (S20)
-No equilibrium fluctuation–dissipation relation is assumed; $D_{\lambda}$ may be arbitrary.
-Our goal is to integrate out $\{X_{\lambda}\}$ and obtain an exact effective MSRJD action for (x,x~)($\tilde{x}$) containing a memory kernel.
+[…]
 (II) MSRJD action for the coupled system: For a generic additive-noise Langevin equation $\dot{y}=f(y)+\zeta$ with $\langle\zeta(t)\zeta(t^{\prime})\rangle=2D\,\delta(t-t^{\prime})$, the MSRJD functional integral may be written (Ito convention) as
@@
 up to an overall normalization independent of $y$. Applying (S21) to (S17)–(S18) gives
-$Z=\int\mathcal{D}x\,\mathcal{D}\tilde{x}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[x,\tilde{x},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S6)
+$Z=\int\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[\phi,\tilde{\phi},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S22)
 with total action
-$S=S_{x}+\int d\lambda\;S_{\lambda},$ (S7)
+$S=S_{\phi}+\int d\lambda\;S_{\lambda},$ (S23)
 where
-$S_{x}$ $=\int dt\;\tilde{x}(t)\!\left[\dot{x}(t)+r\,x(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
-$\quad-\int dt\;D_{x}\,\tilde{x}(t)^{2},$ (S8)
-$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,x(t)\right]$
+$S_{\phi}$ $=\int dt\;\tilde{\phi}(t)\!\left[\dot{\phi}(t)+r\,\phi(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
+$\quad-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2},$ (S24)
+$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,\phi(t)\right]$
 $\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S25)
 (III) Isolating a single $\lambda$-mode contribution: Fix $\lambda$ and collect all terms in $S$ that contain $X_{\lambda}$ or $\tilde{X}_{\lambda}$. From (S24) we extract
-$S_{x}\supset-\int dt\;\tilde{x}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S10)
+$S_{\phi}\supset-\int dt\;\tilde{\phi}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S26)
 Combining (S26) with (S25) yields the $\lambda$-sector action
-$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};x,\tilde{x}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S11)
-$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;g(\lambda)\,\tilde{x}(t)\,X_{\lambda}(t)$
+$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};\phi,\tilde{\phi}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S27)
+$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;g(\lambda)\,\tilde{\phi}(t)\,X_{\lambda}(t)$
 $\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
-In what follows, $x,\tilde{x}$ are treated as external sources.
-(IV) Functional integration over $X_{\lambda}$: emergence of a delta constraint: Define the linear operator
+In what follows, $\phi,\tilde{\phi}$ are treated as external sources.
+(IV) Functional integration over $X_{\lambda}$: Define the linear operator
 $\mathcal{L}_{\lambda}\equiv\partial_{t}+\lambda.$ (S28)
@@
 Then (S27) becomes
-$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t)\Big)$ (S18)
-$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
+$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t)\Big)$ (S34)
+$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
 Now $X_{\lambda}$ enters only linearly. Hence its functional integral produces a functional delta constraint:
@@
 where here
-$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t).$ (S20)
+$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t).$ (S36)
 Therefore,
 […]
-$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
-The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $x,\tilde{x}$ and can therefore be absorbed into the overall normalization of $Z$.
+$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
+The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $\phi,\tilde{\phi}$ and can therefore be absorbed into the overall normalization of $Z$.
 (V) Solving the delta constraint for $\tilde{X}_{\lambda}$ via Green functions: The constraint
-$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{x}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S22)
+$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{\phi}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S38)
 is a linear inhomogeneous ODE for $\tilde{X}_{\lambda}(t)$. Introduce the Green function $G_{\lambda}^{A}(t-t^{\prime})$ solving
 $(-\partial_{t}+\lambda)\,G_{\lambda}^{A}(t-t^{\prime})=\delta(t-t^{\prime}).$ (S39)
-Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $x$-sector response).
+Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $\phi$-sector response).
 The solution with advanced support is
@@
 (Indeed, for $t\neq 0$, $(-\partial_{t}+\lambda)G_{\lambda}^{A}=0$ and the jump at $t=0$ produces the delta distribution.) Multiplying (S38) by $G_{\lambda}^{A}$ and integrating, we obtain
-$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime}).$ (S25)
-Substituting (S25) into (S21) eliminates $\tilde{X}_{\lambda}$ entirely and yields an effective contribution to the (x,x~)($\tilde{x}$) action.
+$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime}).$ (S41)
+[…]
 (VI) Effective action: memory kernel (self-energy) and induced colored noise: We now evaluate the two remaining terms in the exponent of (S37) using (S41).
-1. The $\tilde{X}_{\lambda}x$ term (deterministic memory kernel).
+1. The $\tilde{X}_{\lambda}\phi$ term (deterministic memory kernel).
 Start from
-$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t).$ (S26)
+$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t).$ (S42)
 Insert (S41):
-$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime})\,x(t).$ (S27)
+$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime})\,\phi(t).$ (S43)
 Relabel integration variables $(t,t^{\prime})\to(t^{\prime},t)$:
-$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{x}(t)\,x(t^{\prime}).$ (S28)
+$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{\phi}(t)\,\phi(t^{\prime}).$ (S44)
 Using the identity $G_{\lambda}^{A}(t^{\prime}-t)=G_{\lambda}^{R}(t-t^{\prime})$ with the retarded Green function
@@
 we obtain
-$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)$ $=\int dt\,dt^{\prime}\;\tilde{x}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,x(t^{\prime}).$ (S30)
+$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)$ $=\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,\phi(t^{\prime}).$ (S46)
 Thus a single OU mode contributes the causal memory kernel
@@
 Insert (S41):
-$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$ (S33)
-$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{x}(t^{\prime})\,\tilde{x}(t^{\prime\prime}).$ (S34)
+$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$
+$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{\phi}(t^{\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S49)
 Define the induced (generally nonlocal) noise kernel
@@
 so that
-$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{x}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{x}(t^{\prime\prime}).$ (S36)
+$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{\phi}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S51)
 In equilibrium one may impose a fluctuation–dissipation relation linking $\mathcal{N}_{\lambda}$ and $K_{\lambda}$, but this is not required for the deterministic self-energy structure derived below.
@@
 Combining (S46)–(S51) for all $\lambda$ yields the exact effective action
-$S_{\mathrm{eff}}[x,\tilde{x}]=\int dt\;\tilde{x}(t)\big(\dot{x}(t)+rx(t)\big)-\int dt\;D_{x}\,\tilde{x}(t)^{2}$
-$-\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{x}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{x}(t^{\prime}),$ (S37)
+$S_{\mathrm{eff}}[\phi,\tilde{\phi}]=\int dt\;\tilde{\phi}(t)\big(\dot{\phi}(t)+r\phi(t)\big)-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2}$
+$-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$ (S52)
 where the total memory kernel is the reservoir sum
@@
 The bilinear term involving $K$ becomes
-$\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{x}(-\omega)\,K(\omega)\,x(\omega),$ (S40)
+$\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{\phi}(-\omega)\,K(\omega)\,\phi(\omega),$ (S55)
 where $K(\omega)$ is the Fourier transform of the causal kernel (S53):
@@
 $K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S58)
-In the MSRJD response sector, the retarded inverse propagator for $x$ is read off from the $\tilde{x}(-\omega)x(\omega)$ coefficient in $S_{\mathrm{eff}}$:
+In the MSRJD response sector, the retarded inverse propagator for $\phi$ is read off from the $\tilde{\phi}(-\omega)\phi(\omega)$ coefficient in $S_{\mathrm{eff}}$:
 $G_{R}^{-1}(\omega)=-i\omega+r+K(\omega).$ (S59)
@@
 where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
-In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates (e.g.
-from Jacobian eigenvalues).
+In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates.
 In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
 $\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S65)
-This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective variable.
+This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
 (VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
@@
 This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
-Appendix B Appendix B: Exact emergence of a memory kernel from integrating out an OU reservoir
+(X) Relation to the Keldysh functional integral: We briefly clarify the formal relation between the MSRJD formulation employed in the present work and the real–time Keldysh functional integral for driven–dissipative quantum systems.
+The MSRJD action for a stochastic dynamical variable $\phi(t)$ with memory kernel and Gaussian noise may be written as
+$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S69)
+$\quad\phi(t^{\prime})-\int dt\,dt^{\prime}\,\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$
+where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
+The real–time Keldysh functional integral for an open quantum system is naturally formulated in terms of classical and quantum fields, $\phi_{\mathrm{cl}}$ and $\phi_{q}$, defined by the Keldysh rotation
+$\phi_{\mathrm{cl}}=\frac{\phi_{+}+\phi_{-}}{2},\qquad\phi_{q}=\phi_{+}-\phi_{-},$ (S70)
+where $\phi_{\pm}$ denote fields on the forward and backward branches of the time contour.
+To quadratic order in the quantum field, the Keldysh action assumes the general structure
+$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S71)
+where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component.
+The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
+In the regime of strong decoherence or large occupation, where quantum interference effects are suppressed, the quantum field $\phi_{q}$ acts as a response field. Identifying
+$\phi_{q}\;\leftrightarrow\;i\tilde{\phi}$, (S72)
+and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$.
+The corresponding MSRJD action then takes the form (S69), with the correspondence
+$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S73)
+Thus the MSRJD formulation used in the present work can be viewed as the classical (or strongly dissipative) limit of the Keldysh functional integral, in which the relevant dynamical information is encoded in the retarded kernel and its associated spectral properties.
+Importantly, our theory does not rely on coherent quantum interference or operator-valued field structure.
+Instead, the central object is the spectrum of the dynamical generator (equivalently, the retarded kernel), whose low-frequency density of states controls the infrared behavior and pairing amplification mechanism discussed in the main text.
+Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
 In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate.
 The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$.
-(I) Coupled Markovian dynamics: collective field plus relaxational reservoir: For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space.
+This time-domain elimination is equivalent in content to the Gaussian MSRJD integration, but is presented here directly at the level of the coupled equations of motion.
+(I) Coupled Markovian dynamics (collective field plus relaxational reservoir): For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space.
 We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
@@
 Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources.
+For $\lambda>0$ each reservoir coordinate is exponentially stable; the sector $\lambda\to 0^{+}$ should be understood as the infrared limit of a continuous spectrum of increasingly slow relaxational modes rather than an exactly conserved degree of freedom.
 Equations (S74)–(S75) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
@@
 Solving for $X_{\lambda}(t)$ yields the exact causal representation
-$X_{\lambda}(t)$ (S59)
-$\quad=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
+$X_{\lambda}(t)$ $=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$ (S79)
+$\quad+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
 This expression already shows that the reservoir coordinate at time $t$ depends on the entire history $\{\phi(s)\}_{s<t}$.
@@
 $+\eta(t)$
-$=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$
-$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
-$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s)+\eta(t).$ (S60)
-We now isolate the term proportional to the history $\phi(s)$:
-$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
-$\quad=\int_{t_{0}}^{t}ds\,\Bigg[\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}e^{-\lambda(t-s)}\Bigg]\phi(s).$ (S61)
-This motivates the definition of the (causal) memory kernel
-$K(t-s)\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},\qquad t\geq s,$ (S62)
-so that (S61) becomes simply
+This motivates the definition of the causal memory kernel
+$K(t-s)\equiv\Theta(t-s)\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},$ (S81)
+so that (LABEL:eq:kernel_identification_step_app) becomes
 $\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S82)
-With this definition, Eq. (S60) can be written as
+With this definition, Eq. (LABEL:eq:phi_substituted_app) can be written as
 $\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S83)
@@
 For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays.
-More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0$ sector does not produce a non-decaying contribution.
-(IV) TDOS form and Laplace-transform structure: Defining the time-scale density of states as
-$\rho(\lambda)\equiv g(\lambda)^{2},$ (S67)
-(Here $\rho(\lambda)$ should be understood as an effective spectral weight for the coupling of relaxation rates to $\phi$; in a microscopic discrete-mode representation one naturally has $\rho(\lambda)\overline{g^{2}}(\lambda)$, and we absorb any slowly varying $\overline{g^{2}}(\lambda)$ into $\rho(\lambda)$ for notational simplicity.)
-the memory kernel (S62) becomes
+More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0^{+}$ sector does not produce a non-decaying contribution.
+(IV) TDOS form and Laplace-transform structure: The memory kernel identified above depends only on the combination $g(\lambda)^{2}$ that weights the contribution of each relaxational mode to the collective dynamics.
+It is therefore natural to introduce the TDOS as the effective coupling-weighted spectral density of relaxation rates:
+$\rho(\lambda)\equiv g(\lambda)^{2}.$ (S86)
+More generally, in a microscopic description with discrete relaxational modes of density $\nu(\lambda)$ and coupling amplitudes $g_{\lambda}$, the kernel involves the combination $\nu(\lambda)|g_{\lambda}|^{2}$.
+For notational simplicity, we absorb any slowly varying factors into the definition of $\rho(\lambda)$.
+Thus $\rho(\lambda)$ should be understood as the effective spectral weight governing how strongly relaxation modes at scale $\lambda$ influence the collective coordinate $\phi$.
+With this definition, the memory kernel becomes
 $K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S87)
-Thus $K(t)$ is precisely the Laplace transform of the TDOS (with cutoff $\Lambda$).
-In this sense the TDOS is the fundamental spectral object, and the time-domain memory kernel is its direct transform.
+Equation (S87) shows that the memory kernel is precisely the Laplace transform of the TDOS (with ultraviolet cutoff $\Lambda$).
+In this sense the TDOS is the fundamental spectral object of the theory, while the time-domain memory kernel is its direct transform.
+The infrared dynamical behavior of the collective coordinate is therefore completely determined by the low-$\lambda$ structure of $\rho(\lambda)$.
 (V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
@@
 then for times $t\gg\Lambda^{-1}$ the infrared part of (S87) dominates and we may approximate $\rho(\lambda)\simeq\rho_{0}$ over the relevant range. In that case the kernel can be evaluated exactly:
-$K(t)$ $=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}=\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}.$ (S70)
+$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}.$ (S89)
 Perform the change of variables $u=\lambda t$ (so $d\lambda=du/t$):
-$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\int_{0}^{\Lambda t}du\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S71)
+$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S90)
 Therefore, at long times $t\gg\Lambda^{-1}$,
@@
 $\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S94)
-(We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
+We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
 Separating real and imaginary parts using
@@
 $=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
-$=\frac{\rho_{0}}{2}\Big(\ln(\Lambda^{2}+\omega^{2})-\ln(\omega^{2})\Big)$
 $=\frac{\rho_{0}}{2}\ln\!\left(\frac{\Lambda^{2}+\omega^{2}}{\omega^{2}}\right)\simeq\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$
@@
 This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
-(VII) Conceptual structure: from hidden modes to infrared memory: The derivation above establishes a general structural result.
+(VII) Conceptual structure (from hidden modes to infrared memory): The derivation above establishes a general structural result.
 A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated.
-The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states (TDOS),
+The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states,
 $K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)e^{-\lambda t}.$ (S100)
@@
 the Laplace transform produces the universal long-time tail
-$K(t)\simeq\frac{\rho_{0}}{t},\qquad t\gg\Lambda^{-1}.$ (S83)
+$K(t)\simeq\frac{\rho_{0}}{t}.$ (S102)
 Thus long-memory infrared dynamics is not an additional assumption but an unavoidable consequence of a flat slow-mode spectrum.
@@
 In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
-Appendix C Appendix C: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
+Appendix D Appendix D: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
 In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel.
@@
 This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
-Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
+Appendix E Appendix E: Finite-temperature regulation of the infrared divergence
 In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$.
@@
 The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
-Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
+Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
 In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime.
@@
 Equation (S139) (or equivalently the collapse form (S145)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.
-Appendix F Appendix F: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
-In conventional quantum many-body theory, collective behavior is typically organized in terms of energy eigenmodes of an underlying Hamiltonian,
-$H|n\rangle=E_{n}|n\rangle,$ (S129)
-with dynamical response expressed through Fourier modes oscillating at frequencies set by energy differences.
-This structure naturally emphasizes stationary states and conservative dynamics, with dissipation treated as a secondary perturbative effect.
-However, this description represents a special limit corresponding to closed, Hamiltonian systems.
-In general interacting and open systems — particularly near criticality — collective dynamics is governed not by energy eigenmodes but by the stability spectrum of the full dynamical flow.
-(I) Dynamical linearization and relaxation modes: For a general high-dimensional dynamical system
-$\dot{x}(t)=F(x(t)),$ (S130)
-linearization about a metastable configuration $x^{*}$ yields
-$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S131)
-The eigenvalue problem
-$Av_{i}=\lambda_{i}v_{i}$ (S132)
-defines a spectrum of collective relaxation modes with temporal evolution
-$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S133)
-The real parts of $\lambda_{i}$ set intrinsic time scales of decay, while imaginary parts correspond to oscillatory components.
-Energy eigenmodes arise only in the special Hamiltonian limit where all eigenvalues lie purely on the imaginary axis.
-Thus, the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural and generally applicable dynamical basis for collective phenomena.
-(II) Liouvillian formulation and relation to energy spectra: In quantum dynamical field theories formulated in Keldysh or MSRJD language, evolution of the density matrix is generated by a Liouvillian superoperator $\mathcal{L}$,
-$\frac{d\rho}{dt}=\mathcal{L}\rho.$ (S134)
-For closed Hamiltonian dynamics,
-$\mathcal{L}\rho=-i[H,\rho],$ (S135)
-yielding eigenvalues
-$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S136)
-which lie purely on the imaginary axis and reproduce energy-difference oscillations.
-In interacting open systems and critical regimes, additional dissipative and collective processes shift Liouvillian eigenvalues into the complex plane, producing a dense spectrum of decay rates.
-The fundamental objects governing long-time dynamics are therefore relaxation rates rather than energy levels.
-(III) Memory kernels and continuous relaxation spectra: General linear causal dynamics with memory takes the form
-$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S137)
-Taking the Laplace transform gives
-$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S138)
-The poles and branch cuts of $s+\tilde{K}(s)$ define the intrinsic relaxation spectrum.
-When $\tilde{K}(s)$ possesses branch-cut structure, the response is governed by a continuous distribution of decay rates,
-$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S139)
-which uniquely defines the time-scale density of states.
-Thus continuous relaxation spectra emerge generically from causality and memory, independent of microscopic details.
-(IV) Energy modes as a special case: The familiar Fourier representation of conservative dynamics corresponds to the singular limit in which the TDOS collapses onto the imaginary axis,
-$\rho(\lambda)\propto\delta(\mathrm{Re}\,\lambda),$ (S140)
-recovering purely oscillatory energy eigenmodes. Away from this limit, generic interacting systems naturally develop broad relaxation spectra controlling infrared behavior.
-(V) Physical implication: The time-scale density of states therefore represents the fundamental organizing variable of collective dynamics in correlated systems.
-Energy eigenmodes describe a special integrable limit, while relaxation-rate spectra capture the universal structure of open, critical, and strongly interacting matter.
-Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes.
-The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more general foundation than energy-based descriptions.

### v3

@@
-arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026
+Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity
 Abstract
@@
 These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
-A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11].
-Such descriptions implicitly assume that the infrared behavior can be captured by a finite set of soft bosonic modes with local-in-time damping.
-However, strongly correlated quantum matter is fundamentally governed by Liouvillian evolution of the density matrix, where the full spectrum of relaxation eigenvalues determines long-time dynamics [12, 13, 14].
-In more general, criticality need not correspond to softening of a single excitation mode but may instead involve reorganization of the entire relaxation-rate spectrum.
-Rather than reducing the dynamics to a few collective coordinates, one must consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes.
+A key limitation of many conventional approaches lies not in the use of collective fields themselves, but in the assumption that the infrared dynamics can be adequately described by a small number of soft bosonic modes with local-in-time damping [10, 11].
+Such descriptions effectively close the theory within a finite set of order-parameter degrees of freedom.
+While this approximation is often adequate in weakly interacting systems, its validity becomes less clear in strongly correlated electronic matter, where experimentally relevant observables correspond to coarse-grained collective sectors that continuously exchange coherence and energy with a large number of internal degrees of freedom.
+In such situations the effective dynamics need not be governed by a finite set of collective modes, but may instead involve a broad spectrum of relaxation processes spanning many time scales.
+Although the full many-body system evolves unitarily under a Hamiltonian, the effective dynamics of these collective sectors is more naturally described as that of an open subsystem.
+In this situation the appropriate dynamical generator is not simply the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes of the reduced density matrix [12, 13, 14].
+The long-time behavior of the system is therefore controlled by the spectrum of relaxation eigenvalues of this Liouvillian operator.
+From this perspective, dynamical criticality need not correspond to the softening of a single excitation mode but may instead arise from a reorganization of the entire relaxation-rate spectrum.
+Rather than reducing the dynamics to a few collective coordinates, one must therefore consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes.
 Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality.
@@
 By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
-We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics.
-While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$.
-This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a distinct universality class governed by the structure of the relaxation spectrum.
+We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior.
+Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility.
+While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to the case $\alpha=1$, arising microscopically from particle–hole polarization of a Fermi liquid, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic kernel proportional to $\mathrm{sgn}(\omega)$.
+In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$.
+The latter regime reflects a fundamentally different physical mechanism: rather than overdamped order-parameter fluctuations generated by a fermionic bath, the infrared dynamics is governed by the collective organization of the relaxation spectrum itself.
+This memory-dominated regime therefore defines a distinct universality class controlled by the structure of the TDOS.
 Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20].
@@
 $\rho(\lambda)\sim\lambda^{\alpha},\qquad\lambda\to 0,$ (17)
-the dissipative part of the collective kernel follows directly from the spectral representation in Eq. (16),
+the collective susceptibility follows directly from the spectral representation in Eq. (16).
+The dissipative part of the kernel is
 $\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (18)
-Here $\omega$ denotes the collective relaxation frequency arising from the projected Liouvillian dynamics, rather than a microscopic quasiparticle excitation energy.
-It is therefore natural to write the long–wavelength dynamical response in the scaling form
+Because the response function is causal, the reactive component is not independent but is related to the dissipative part through the Kramers–Kronig relation.
+The real part therefore inherits the same infrared scaling structure and, depending on the TDOS exponent, produces logarithmic or subleading corrections to the static stiffness.
+It is convenient to express the long–wavelength response in the scaling form
 $\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (19)
-in direct analogy with the conventional Hertz–Millis representation [3, 4]. The TDOS scaling immediately implies the universal identification $s=\alpha$.
-Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$.
-Requiring both contributions to remain of comparable order at criticality ($r=0$) fixes the dynamical exponent to
-$z=\frac{2}{s}.$ (20)
-In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the organization of the underlying relaxation spectrum.
-Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$.
-Infrared universality classes are thus classified directly by the low–energy structure of the relaxation-rate density.
-In particular, when the TDOS remains finite at the origin,
-$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (21)
-the dissipative kernel simplifies to
+where the TDOS scaling directly implies the identification $s=\alpha$.
+While this form resembles the conventional Hertz–Millis representation, its physical interpretation is different.
+In the present framework the exponent $s$ does not originate from fermionic damping of a single order–parameter mode but instead reflects the infrared organization of the entire relaxation spectrum.
+The appropriate renormalization viewpoint is therefore naturally formulated in terms of temporal coarse–graining.
+Integrating out fast relaxation processes with decay rates $\lambda>\Lambda/b$ corresponds to eliminating short time scales $t\lesssim b^{-1}\Lambda^{-1}$.
+Under this temporal renormalization step the remaining effective theory is characterized by a renormalized TDOS
+$\rho_{b}(\lambda)=b^{\alpha}\rho(b\lambda),$ (20)
+so that the infrared scaling exponent $\alpha$ defines the universality class of the relaxation spectrum.
+Infrared universality classes are therefore classified directly by the low–energy structure of the relaxation–rate density.
+A particularly important case occurs when the TDOS remains finite at the origin,
+$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0}.$ (21)
+In this flat–TDOS regime the dissipative kernel simplifies to
 $\mathrm{Im}\,\chi_{R}^{-1}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{0}}{\lambda^{2}+\omega^{2}}$
 $\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (22)
-while the reactive part acquires a logarithmic infrared singularity,
+while the reactive component develops a logarithmic infrared singularity,
 $\mathrm{Re}\,\chi_{R}^{-1}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{\lambda}{\lambda^{2}+\omega^{2}}\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (23)
-Collective infrared dynamics is therefore governed by the marginal non-Markovian kernel
+Collective infrared dynamics is therefore governed by the marginal non–Markovian kernel
 $\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (24)
-corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
-This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory.
-Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow decay modes.
+which corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
+This flat–TDOS fixed point represents a dynamical regime in which the infrared behavior is governed by the accumulation of a continuum of slow decay modes.
+Rather than arising from overdamped fluctuations of a single soft order parameter, collective relaxation is controlled by the spectral organization of the full relaxation spectrum.
 III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
-The emergence of a slow–mode reservoir reorganizes the infrared structure of the Cooper pairing channel.
-In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal.
-Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
-III.1 A. Marginal pairing from Ohmic collective dynamics
-Near metallic quantum critical points, fermions couple to overdamped collective modes with retarded propagator
+The dynamical framework developed in the previous section implies that strongly correlated quantum matter can self–organize into a regime dominated by a continuum of slow collective modes.
+Instead of a single overdamped bosonic mediator, the infrared dynamics is governed by the full relaxation spectrum characterized by the time–scale density of states $\rho(\lambda)$.
+This reorganization of the collective spectrum qualitatively alters the structure of the Cooper channel.
+In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”.
+In contrast, in the present framework superconductivity emerges from the dynamical amplification of pairing correlations generated by the slow–mode reservoir itself.
+In this section we derive the infrared structure of the Cooper channel in this regime and contrast it with the conventional Eliashberg mechanism.
+III.1 A. Marginal pairing from overdamped collective dynamics
+Near metallic quantum critical points, electrons often couple to overdamped collective modes whose retarded propagator takes the form
 $D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (25)
-The linear frequency dependence reflects Landau damping by the electronic continuum and produces a pairing kernel scaling as
-$D(\omega)\sim\frac{1}{|\omega|}.$ (26)
-The infrared shell correction to the dimensionless Cooper–channel coupling $g$ therefore reads
-$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=g^{2}\ln b.$ (27)
-Introducing the RG scale $l=\ln b$ yields
-$\frac{dg}{dl}=Cg^{2},$ (28)
-with $C>0$.
-The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining.
-Physically, overdamped collective modes dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
-III.2 B. Infrared reorganization by slow collective modes
-In the memory–dominated regime derived in Sec.
-II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$.
-After integrating out the relaxation reservoir, the retarded collective susceptibility takes the exact spectral form
-$\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (29)
-For a flat TDOS in the infrared, $\rho(\lambda\to 0)=\rho_{0}$, the effective polarization kernel in the Cooper channel inherits the infrared scaling
-$\Pi(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{1}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},\qquad|\omega|\to 0.$ (30)
-Unlike the conventional Ohmic case, this singular behavior does not originate from Landau damping by fast electronic continua.
-Instead, it emerges from the accumulation of an extensive continuum of near–marginal relaxation modes encoded in a flat TDOS.
-The resulting kernel corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$.
-Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct.
-In Eliashberg theory the singularity reflects overdamped dissipation arising from coupling to a fast electronic bath [25, 26].
-Here, by contrast, the same infrared scaling results from spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
-This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales.
-The slow–mode reservoir therefore converts the Cooper channel into a relevant infrared interaction.
-III.3 C. Relevant infrared amplification of the Cooper channel
-The infrared shell correction in the memory–dominated regime becomes
-$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}\ln(\Lambda/\omega^{\prime})}.$ (31)
-Introducing $u=\ln(\Lambda/\omega^{\prime})$ yields
-$dg\simeq\alpha g\,dl,$ (32)
-with $\alpha>0$ proportional to the infrared TDOS weight.
-The resulting RG flow reads
-$\frac{dg}{dl}=\alpha g,$ (33)
-with solution $g(l)=g_{0}e^{\alpha l}$.
-The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one.
-This qualitative contrast with Eliashberg theory is illustrated in Fig.
+Here $D_{R}$ denotes the effective propagator of a coarse-grained order-parameter (particle–hole) collective mode in the Hertz–Millis sense: the $|\omega|$ term is generated by the electron–hole polarization of the metallic continuum (Landau damping) [3, 4].
+This object should be distinguished from the Cooper-channel polarization $\Pi(\omega)$ introduced below, which diagnoses the pairing instability itself.
+The linear frequency dependence reflects Landau damping due to decay of collective excitations into the electronic continuum.
+After momentum integration, the pairing kernel entering the Cooper channel behaves approximately as
+$D_{R}(\omega)\sim\frac{1}{|\omega|}.$ (26)
+The linearized gap equation may be written schematically as
+$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
+where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff.
+To analyze the infrared flow we eliminate fast frequency modes with $\omega^{\prime}\in[\Lambda/b,\Lambda]$, corresponding to coarse–graining of short time scales.
+The ladder diagram gives the leading correction
+$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}.$ (28)
+The shell integral yields
+$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln b.$ (29)
+Introducing the RG scale $l=\ln b$ gives
+$\frac{dg}{dl}=Cg^{2}.$ (30)
+The Cooper interaction therefore grows only logarithmically under infrared coarse graining.
+In RG language the pairing interaction is marginal.
+Physically, overdamped collective modes dissipate into the electronic continuum and repeatedly lose phase coherence.
+As a result superconducting correlations accumulate only logarithmically as the energy scale is reduced.
+III.2 B. Collective relaxation spectrum and TDOS representation
+Strongly correlated quantum matter, however, can develop a much richer infrared structure.
+Instead of a single overdamped bosonic excitation, the system may host a dense continuum of slow collective modes.
+The collective dynamics can then be represented by the TDOS $\rho(\lambda)$, which measures the density of relaxation rates $\lambda$.
+Projecting the dynamics onto the slow collective sector yields the spectral representation
+$\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
+The polarization kernel entering the Cooper channel becomes
+$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (32)
+Here $\Pi(\omega)$ represents the polarization of the Cooper channel generated by the collective relaxation spectrum.
+It should be distinguished from the conventional bosonic polarization that appears in propagator renormalization.
+In the present framework $\Pi(\omega)$ directly measures the dynamical pair susceptibility produced by the slow collective modes.
+For a flat TDOS in the infrared,
+$\rho(\lambda\to 0)=\rho_{0},$ (33)
+the polarization acquires a universal singular form
+$\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (34)
+In the time domain this corresponds to a long–memory kernel
+$K(t)\sim\frac{1}{t}.$ (35)
+Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves.
+III.3 C. Infrared amplification of pairing correlations
+The Cooper susceptibility takes the form
+$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (36)
+Because $\Pi(\omega)$ diverges as $|\omega|\to 0$, pairing correlations are strongly enhanced as the energy scale is lowered.
+To understand the infrared flow more generally, consider a TDOS scaling
+$\rho(\lambda)\sim\lambda^{\eta}.$ (37)
+Substituting this form into the polarization integral yields
+$\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (38)
+The effective pairing kernel therefore scales as
+$g_{\mathrm{eff}}(\omega)\sim|\omega|^{-(1-\eta)}.$ (39)
+Under RG rescaling $\omega\rightarrow\omega/b$ the kernel transforms as
+$g^{\prime}=b^{1-\eta}g.$ (40)
+Introducing $l=\ln b$ gives the tree–level RG flow
+$\frac{dg}{dl}=(1-\eta)g.$ (41)
+For $\eta<1$ the Cooper interaction becomes relevant in the infrared. Including the usual ladder correction yields
+$\frac{dg}{dl}=(1-\eta)g+Cg^{2}.$ (42)
+The linear term dominates the early RG flow, leading to algebraic growth
+$g(l)=g_{0}e^{(1-\eta)l}.$ (43)
+Physical interpretation.
+The qualitative difference between the Eliashberg regime and the present memory–dominated regime originates from the structure of the collective spectrum, as illustrated schematically in Fig.
 2.
-A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix D.
-Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and amplify intrinsic electronic pairing tendencies.
-Superconductivity in strongly correlated systems thus emerges from infrared spectral organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
+In the Eliashberg case, each logarithmic shell contributes independently, leading to marginal enhancement of pairing correlations.
+In contrast, a flat TDOS corresponds to a dense continuum of near–marginal relaxation modes.
+Fluctuations across widely separated time scales remain dynamically correlated, allowing pairing correlations to accumulate coherently.
+Superconductivity therefore emerges as a dynamical instability of the collective relaxation spectrum itself, rather than from the exchange of a specific bosonic mediator.
 Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction.
@@
 $H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (45)
-favoring local singlet formation.
-More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated purely by electronic interactions, independent of long–wavelength collective mediators.
-This intrinsic scale provides the microscopic seed for superconductivity.
+which favors local singlet formation. This interaction provides a natural microscopic origin of electronic pairing in systems near Mott localization.
+More generally, interacting electrons are described by a generic four–fermion interaction
+$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma}.$ (46)
+This interaction contains multiple collective channels.
+In conventional diagrammatic approaches these channels are analyzed through resummations such as particle–hole bubble diagrams or particle–particle ladder diagrams.
+Rather than focusing on a specific mediator, the present framework emphasizes the dynamical organization of the collective spectrum itself.
+While the four–fermion interaction generically allows multiple ordering channels, their ordering tendencies are not equivalent.
+Particle–hole instabilities such as charge or spin density waves typically require specific momentum structure, most notably Fermi surface nesting or commensurate wave vectors determined by the lattice.
+Such kinematic constraints restrict the available phase space for density–wave ordering.
+Superconductivity, by contrast, corresponds to condensation in the particle–particle channel at zero total momentum,
+$\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (47)
+and therefore does not rely on special nesting conditions of the Fermi surface.
+Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
 Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
 $\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (48)
-with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
+where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations.
+Diagrammatically $\Pi(\omega)$ corresponds to the Cooper bubble, which resums particle–particle ladder diagrams.
+In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory.
+Near correlated criticality, however, the situation changes qualitatively.
+Because the relaxation spectrum becomes dense near $\lambda\to 0$, an extensive reservoir of slow collective modes emerges.
+This slow–mode reservoir reorganizes the infrared dynamics of the electronic system and strongly enhances the Cooper channel.
+Within the present framework critical dynamics does not create attraction.
+Instead it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
+The intrinsic pairing scale $E_{\rm pair}$ therefore acts as the microscopic seed, while the slow collective spectrum dynamically amplifies the Cooper instability.
 IV.2 B. Memory–dominated Cooper kernel from flat TDOS
-Because single–particle propagators admit a relaxational spectral representation governed by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
+Because single–particle propagation inherits the relaxational spectral structure generated by the slow–mode reservoir, the Cooper bubble reduces to a continuum integral over relaxation rates,
 $\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (49)
@@
 This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales.
-Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel becomes relevant due to the flat TDOS.
-The slow–mode reservoir therefore reorganizes the pairing problem into a memory–dominated critical regime.
+Physically, the flat TDOS implies that an extensive set of slow collective modes contributes coherently to the Cooper channel.
+Unlike conventional Eliashberg kernels, which remain marginal in the infrared, the present Cooper kernel becomes strongly enhanced by the infrared spectral weight of the slow–mode reservoir.
+The pairing interaction is therefore dynamically amplified by the collective memory of the system.
 IV.3 C. Algebraic Instability and Transition Scale
@@
 $\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (52)
-which yields
+which yields the algebraic transition scale
 $T_{c}\sim\rho_{0}\,E_{\rm pair},$ (53)
-up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix E.
+up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
 The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir.
@@
 In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$.
-Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$ [Eq.
-(39)], the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight.
+Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight.
 Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
@@
 We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
-The TDOS controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response.
+The density of relaxation rates controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response.
 Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
 This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling.
-The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
+The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
+Figure 4: Universality map relating the relaxation-rate spectrum to infrared collective dynamics and superconducting pairing.
+The TDOS exponent $\alpha$ determines the scaling of the polarization kernel $\Pi(\omega)\sim|\omega|^{\alpha-1}$.
+The Ohmic Hertz–Millis theory corresponds to the marginal point $\alpha=1$, while a flat TDOS ($\alpha=0$) produces memory-dominated critical dynamics with $\Pi(\omega)\sim 1/|\omega|$ and enhanced pairing.
 VI VI. Discussion
-The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by softening of a single order parameter.
+The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by the softening of a single order parameter.
 The time–scale density of states provides a quantitative measure of this organization.
-When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response.
+When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
 This viewpoint clarifies the limits of conventional coarse-grained field theories.
@@
 Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero.
-In this regime, internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales.
+In this regime internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales.
 The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
-This observation highlights the structural difference from Hertz–Millis–type approaches.
+This observation highlights a structural difference from Hertz–Millis–type approaches.
 Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping.
 In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels.
-Criticality is reinterpreted not as single-mode softening but as reorganization of the entire decay spectrum.
-The TDOS thus replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
-A direct physical consequence of this spectral reorganization is qualitative enhancement of superconducting instabilities.
+Criticality is reinterpreted not as single-mode softening but as a reorganization of the entire decay spectrum.
+The TDOS therefore replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
+This structure is summarized schematically in Fig.
+4.
+The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states.
+For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form
+$\Pi(\omega)\sim|\omega|^{\alpha-1}.$ (65)
+Different dynamical regimes therefore emerge as special cases of the same spectral organization.
+For $\alpha>1$ the density of slow modes vanishes in the infrared, so that $\Pi(\omega)\rightarrow 0$ and the collective dynamics reduces to weakly damped propagating modes with $\chi^{-1}\simeq r+cq^{2}$.
+The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$.
+Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges.
+In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
+The lower panel of Fig.
+4 illustrates the direct consequence of this spectral organization for superconductivity.
+Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic kernel $\Pi(T)\sim\ln(\Omega/T)$ and the exponentially small BCS transition temperature.
+In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$.
+The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
+A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities.
 Superconductivity need not rely on finely tuned bosonic mediators.
@@
 A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena.
-The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; anomalous normal-state dynamics arises from the same non-Markovian kernel.
+The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel.
 Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
+Liouvillian description of infrared collective dynamics.
 The relaxation-rate organization emphasized here is not merely a phenomenological assumption.
@@
 Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra.
-A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, rather than a model-specific construction.
-From this perspective, Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality.
+A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations rather than a model-specific constructione, is derived in Appendix F.
+This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics.
+While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem continuously exchanging coherence and information with the remaining degrees of freedom.
+From the perspective of such a subsystem, the relevant generator of dynamics is not the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes.
+The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues rather than solely by the spectrum of excitation energies.
+In this sense the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
+From this perspective Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality.
 Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
-More broadly, the TDOS framework elevates dynamical spectral organization to a fundamental organizing principle of correlated quantum matter.
-Memory-dominated criticality represents a universality class distinct from traditional dissipative field theories, with potential implications for non-Fermi liquids, glassy systems, and self-organized critical dynamics.
+Temporal scaling and dynamical criticality.
+An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality.
+In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges indirectly through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$.
+The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
+In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves.
+When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales.
+The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and logarithmic infrared response functions.
+Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality.
+This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena.
 Memory–dominated dynamics as a guiding principle for room-temperature superconductivity.
@@
 A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$.
-From this viewpoint, certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
+From this viewpoint certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat-TDOS regime.
 The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27].
 Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself.
-Room-temperature superconductivity, in this perspective, corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
+In this perspective high-temperature superconductivity corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
 VII VII. Conclusion
@@
 Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
-The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during early development.
+The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during development.
 Supplementary Materials
@@
 Appendix B Appendix B: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
-In this appendix we derive how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism.
-The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
+In this appendix we show how a continuum of relaxational (Ornstein–Uhlenbeck) modes generates a nonlocal memory kernel and the associated frequency-dependent self-energy $\Sigma(\omega)$ through Gaussian functional integration within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism.
+The derivation follows the standard structure of Gaussian path integration but is formulated in a way adapted to the slow-mode reservoir perspective developed in this work.
+The construction should be understood as an effective infrared description of collective dynamics rather than a microscopic derivation from an explicit fermionic Hamiltonian.
+At the microscopic level one may introduce collective Hubbard–Stratonovich fields and integrate out fast fermionic degrees of freedom, in the spirit of Hertz-type constructions, which generally produces a nonlocal dynamical kernel for the collective sector.
+Instead of evaluating this kernel microscopically, we represent its infrared structure through a continuum of relaxational modes.
+Within this coarse-grained description the relaxational reservoir is not an additional assumption but a spectral representation of a causal memory kernel.
+Integrating out the reservoir therefore yields a retarded dynamical correction and a frequency-dependent self-energy for the collective field.
+Such a formulation is particularly natural in strongly correlated systems where the infrared dynamics is governed not by a single soft mode but by a broad spectrum of slow relaxation processes.
+In this regime the collective response is controlled by the spectral weight distribution of these modes, leading naturally to the TDOS-based description employed throughout this work.
 (I) Coupled stochastic dynamics: We consider a collective degree of freedom $\phi(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$.
@@
 and likewise $\mathcal{N}=\int d\lambda\,\mathcal{N}_{\lambda}$.
-(VI) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
+(VII) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
 $f(\omega)=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,f(t),\qquad f(t)=\int\frac{d\omega}{2\pi}\;e^{-i\omega t}\,f(\omega).$ (S54)
@@
 $\Sigma_{R}(\omega)\equiv K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S60)
-(Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
+Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
 Equations (S58)–(S60) demonstrate that integrating out a continuum of OU relaxation modes produces a frequency- dependent self-energy whose analytic structure is determined by the relaxation-rate spectrum and couplings.
-(VII) Connection to TDOS (relaxation-rate density of states): If the couplings are sufficiently structureless over the low-$\lambda$ sector, one may absorb $g(\lambda)^{2}$ into an effective spectral density.
-Defining
-$\mathcal{J}(\lambda)\equiv g(\lambda)^{2},$ (S61)
-we can write
-$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega}.$ (S62)
-Equivalently, starting from a discrete set of modes labeled by $\alpha$,
+(VIII) Connection to TDOS (relaxation-rate spectral weight): The spectral representation derived above shows that each relaxational mode contributes to the retarded self-energy with pole $(\lambda_{\alpha}-i\omega)^{-1}$ weighted by the square of its overlap with the collective field.
+Starting from the discrete mode expansion,
 $\Sigma_{R}(\omega)=\sum_{\alpha}\frac{g_{\alpha}^{2}}{\lambda_{\alpha}-i\omega},$ (S61)
-the continuum limit gives
-$\Sigma_{R}(\omega)$ $=\int d\lambda\;\rho(\lambda)\,\frac{\overline{g^{2}}(\lambda)}{\lambda-i\omega}\equiv\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega},$
-$\qquad\mathcal{J}(\lambda)\equiv\rho(\lambda)\,\overline{g^{2}}(\lambda),$ (S64)
-where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
-In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates.
-In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
-$\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S65)
-This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
-(VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
+it is therefore natural to define the TDOS directly as the coupling-weighted spectral distribution of relaxation rates,
+$\rho(\lambda)\equiv\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S62)
+With this definition, the continuum form of the self-energy becomes
+$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S63)
+In this sense, $\rho(\lambda)$ plays a role directly analogous to the Lehmann spectral function in Hamiltonian many-body theory.
+For an operator $O$ with energy eigenstates $|n\rangle$, the usual spectral representation involves weights $|\langle 0|O|n\rangle|^{2}$ multiplying poles at energies $E_{n}$.
+Here, by contrast, the relevant poles are located at the relaxation rates $\lambda_{\alpha}$, and the corresponding spectral weights are the squared overlap amplitudes $g_{\alpha}^{2}$ between the collective field and the relaxation eigenmodes.
+Thus the TDOS should be understood not merely as a bare counting density of modes, but as the physically observable spectral weight distribution of the relaxation spectrum for the collective field under consideration.
+Equation (S63) makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
+(IX) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
 $S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
@@
 with $K(\omega)=\int d\lambda\,g(\lambda)^{2}/(\lambda-i\omega)\equiv\Sigma_{R}(\omega)$. Hence,
-$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega)$ (S68)
-(up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response).
-(IX) Summary of the exact Gaussian result: Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables.
+$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),$ (S66)
+up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response.
+Summary of the exact Gaussian result.
+Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables.
 Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S52) with a causal memory kernel (S53).
@@
 $\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S80)
-With this definition, Eq. (LABEL:eq:phi_substituted_app) can be written as
+With this definition, Eq. (LABEL:eq:kernel_identification_step_app) can be written as
 $\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S81)
@@
 More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0^{+}$ sector does not produce a non-decaying contribution.
-(IV) TDOS form and Laplace-transform structure: The memory kernel identified above depends only on the combination $g(\lambda)^{2}$ that weights the contribution of each relaxational mode to the collective dynamics.
-It is therefore natural to introduce the TDOS as the effective coupling-weighted spectral density of relaxation rates:
-$\rho(\lambda)\equiv g(\lambda)^{2}.$ (S86)
-More generally, in a microscopic description with discrete relaxational modes of density $\nu(\lambda)$ and coupling amplitudes $g_{\lambda}$, the kernel involves the combination $\nu(\lambda)|g_{\lambda}|^{2}$.
-For notational simplicity, we absorb any slowly varying factors into the definition of $\rho(\lambda)$.
-Thus $\rho(\lambda)$ should be understood as the effective spectral weight governing how strongly relaxation modes at scale $\lambda$ influence the collective coordinate $\phi$.
-With this definition, the memory kernel becomes
-$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S87)
-Equation (S87) shows that the memory kernel is precisely the Laplace transform of the TDOS (with ultraviolet cutoff $\Lambda$).
-In this sense the TDOS is the fundamental spectral object of the theory, while the time-domain memory kernel is its direct transform.
-The infrared dynamical behavior of the collective coordinate is therefore completely determined by the low-$\lambda$ structure of $\rho(\lambda)$.
+(IV) TDOS form and Laplace-transform structure:
+As derived in Appendix B, integrating out the continuum of relaxation modes produces a retarded self-energy of the form (S63), where the time-scale density of states is defined as the spectral weight distribution of relaxation eigenmodes,
+$\rho(\lambda)=\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S84)
+Here $g_{\alpha}$ represents the overlap amplitude between the collective field and the relaxation eigenmode $\alpha$, so that $g_{\alpha}^{2}$ plays the role of the spectral weight of that mode.
+In this sense the TDOS is directly analogous to the Lehmann spectral function of Hamiltonian many-body theory, but defined for the relaxation spectrum of the Liouvillian dynamics.
+With this definition the memory kernel introduced above can be written in the compact form
+$K(t)=\int_{0}^{\Lambda}d\lambda\;\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0,$ (S85)
+showing that the memory kernel is the Laplace transform of the TDOS. Consequently the long-time dynamics of the collective field is fully controlled by the infrared structure of $\rho(\lambda)$.
 (V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
@@
 In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
-Appendix D Appendix D: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
-In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel.
-We contrast the conventional Eliashberg regime with the memory-dominated regime generated by a flat relaxation-rate density of states.
-The derivation follows from shell integration of the linearized gap equation and clarifies the origin of marginal versus relevant superconducting instabilities.
-Fermionic Matsubara frequencies are denoted by $\omega$ and $\omega^{\prime}$, and $\Omega\equiv\omega-\omega^{\prime}$ is the bosonic transfer frequency entering the pairing kernel $g(\Omega)$.
-(I) Linearized gap equation and RG shell construction: At low energies the linearized gap equation takes the schematic form
-$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (S103)
-where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff.
-The factor $1/|\omega^{\prime}|$ originates from the fermionic Cooper bubble.
-The effective pairing kernel $g(\Omega)$ is generated by collective fluctuations.
-The RG procedure consists of integrating out a thin frequency shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$ with $b>1$ and absorbing its contribution into a renormalized coupling $g\rightarrow g+dg$.
-Defining the RG scale $l=\ln b$, the flow equation follows from the shell contribution.
-(II) Conventional Eliashberg regime: Near metallic quantum criticality with Ohmic Landau damping, the bosonic propagator obeys
-$D(\Omega)\sim\frac{1}{|\Omega|},$ (S104)
-so that the pairing kernel scales as
-$g(\Omega)\sim g\,D(\Omega)\sim\frac{g}{|\Omega|},$ (S105)
-where $g$ denotes the dimensionless pairing interaction.
-The leading shell correction arises from the ladder diagram and is proportional to
-$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}.$ (S106)
-The shell integral is elementary:
-$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln\Lambda-\ln(\Lambda/b)=\ln b\equiv dl.$ (S107)
-Hence the RG equation takes the marginal form
-$\frac{dg}{dl}=C\,g^{2},$ (S108)
-with $C$ a positive numerical constant.
-The solution is
-$g(l)=\frac{g_{0}}{1-Cg_{0}l}\simeq g_{0}+Cg_{0}^{2}l,$ (S109)
-demonstrating logarithmic enhancement characteristic of a marginal superconducting instability.
-(III) Memory-dominated regime: In the memory-dominated regime derived in Sec. II, the collective susceptibility behaves as
-$\chi(\Omega)\sim\frac{1}{\rho_{0}\ln(\Lambda/|\Omega|)},$ (S110)
-reflecting the flat TDOS of slow relaxation modes.
-The pairing kernel therefore scales as
-$g(\Omega)\sim\frac{g}{\ln(\Lambda/|\Omega|)}.$ (S111)
-The shell correction becomes
-$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\frac{1}{\ln(\Lambda/\omega^{\prime})}.$ (S112)
-Introducing the variable
-$u=\ln\frac{\Lambda}{\omega^{\prime}},\qquad du=-\frac{d\omega^{\prime}}{\omega^{\prime}},$ (S113)
-the integration limits transform as
-$\omega^{\prime}=\Lambda\Rightarrow u=0,\qquad\omega^{\prime}=\Lambda/b\Rightarrow u=\ln b\equiv dl.$ (S114)
-Equation (S112) becomes
-$dg\propto g^{2}\int_{0}^{dl}\frac{du}{u}.$ (S115)
-The apparent divergence at $u\to 0$ is cut off at a microscopic scale $u_{\min}=O(1)$ marking the onset of the logarithmic TDOS regime. The shell contribution reduces to
-$dg=\alpha\,g\,dl,$ (S116)
-where $\alpha>0$ depends on the spectral weight $\rho_{0}$ and the onset scale of the memory-dominated regime.
-The RG equation is therefore
-$\frac{dg}{dl}=\alpha\,g,$ (S117)
-with solution
-$g(l)=g_{0}e^{\alpha l}=g_{0}b^{\alpha}.$ (S118)
-The pairing interaction thus grows as a power law under infrared coarse graining, signaling a relevant superconducting instability.
-(IV) Physical origin of marginal versus relevant behavior: The qualitative difference between Eqs. (S108) and (S117) originates from the infrared structure of the pairing kernel.
-In the Eliashberg regime, each logarithmic shell contributes independently, leading to cumulative logarithmic growth.
-In contrast, the memory-dominated kernel generated by a flat TDOS correlates fluctuations across all time scales, producing linear amplification of the effective interaction under RG.
-Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into a runaway superconducting instability once a slow-mode reservoir is formed.
-Summary.
-Ohmic Landau damping yields a marginal Cooper-channel flow $dg/dl\propto g^{2}$, producing only logarithmic enhancement.
-A flat relaxation-rate TDOS instead generates a memory-dominated kernel that renders the pairing interaction relevant, with $dg/dl\propto g$ and power-law growth.
-This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
-Appendix E Appendix E: Finite-temperature regulation of the infrared divergence
+Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
 In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$.
@@
 The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
-Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
+Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics.
+(a) Absolute spectral line shape at $q=0$: the TDOS theory yields a broad continuum, whereas HM produces a sharper peak.
+(b) Momentum dependence of the TDOS spectral function: varying $q$ shifts the hump through $\Delta_{q}=r+cq^{2}$.
+(c) TDOS prediction (collapse): when plotted versus the scaling variable $u=\Delta_{q}/\rho_{0}+\ln(\Lambda/|\omega|)$, the TDOS spectra collapse onto a universal curve, $A(\omega,q)=\frac{\pi}{\rho_{0}}\,[u^{2}+(\pi/2)^{2}]^{-1}$.
+(d) Control test: HM spectra do not exhibit the same log-shift collapse.
+(e) Finite-temperature regulation implemented by $|\omega|\rightarrow\max\{|\omega|,cT\}$ yields infrared saturation.
+(f) Corresponding $T$-diagnostic at $q=0$: $A_{T}(\omega\!\to\!0)$ varies systematically with $\ln(\Lambda/T)$, providing a practical fitting protocol.
+All curves are shown in arbitrary units with consistent prefactor conventions.
+Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
 In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime.
@@
 $\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S117)
-(Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.)
+Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.
 (II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
@@
 For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
-(VII) Physical interpretation: Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part.
+Physical interpretation.
+Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part.
 The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
-Equation (S139) (or equivalently the collapse form (S145)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.
+(VII) Experimental diagnostics and scaling tests: Figure S1 summarizes a set of practical diagnostics that distinguish the memory–dominated TDOS regime from conventional overdamped critical dynamics such as the Hertz–Millis (HM) theory.
+Panel (a) compares the absolute spectral line shapes at $q=0$ predicted by the two frameworks.
+The HM theory produces a relatively narrow Lorentzian-like peak, reflecting dynamics controlled by a small number of overdamped bosonic modes.
+In contrast, Eq.
+(S121) yields a much broader continuum with slowly varying logarithmic tails, arising from the collective response of an extensive continuum of slow relaxation modes.
+Panel (b) illustrates the predicted momentum dependence of the TDOS spectral function.
+Because the tuning parameter enters through $\Delta_{\mathbf{q}}=r+cq^{2}$, varying the momentum effectively shifts the spectral hump according to the scale $|\omega_{*}|=\Lambda\exp(\Delta_{\mathbf{q}}/\rho_{0})$ derived in Eq.
+(S125).
+This produces a characteristic logarithmic displacement of the hump position as $q$ changes.
+Panels (c) and (d) demonstrate the most direct experimental test of the TDOS framework. Introducing the scaling variable
+$u(\omega,\mathbf{q})=\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S130)
+all TDOS spectra collapse onto the universal curve (S127).
+This collapse reflects the fact that the full line shape depends only on the combined variable $u$.
+By contrast, spectra generated by the conventional Hertz–Millis form do not collapse under the same transformation, providing a clear diagnostic distinction.
+Panel (e) illustrates the finite–temperature regulation discussed in Eq.
+(S129).
+Replacing $|\omega|$ by $\max\{|\omega|,cT\}$ cuts off the logarithmic infrared growth and produces saturation of the low–frequency spectral weight.
+The resulting spectra evolve smoothly from the zero–temperature form to a temperature–dominated plateau.
+Finally, panel (f) shows a convenient experimental diagnostic at $q\simeq 0$.
+In the regime $|\omega|\ll T$, Eq.
+(S129) predicts that the infrared spectral weight varies systematically with $\ln(\Lambda/T)$.
+Measuring $A_{T}(\omega\!\rightarrow 0)$ as a function of temperature therefore provides a direct way to extract the TDOS parameter $\rho_{0}$ and test the predicted logarithmic scaling.
+Taken together, these diagnostics provide a practical experimental protocol for identifying the memory–dominated TDOS universality class.
+In particular, the logarithmic peak shift and the universal collapse of line shapes constitute strong signatures of collective dynamics governed by an extensive slow–mode reservoir rather than by a single overdamped critical mode.
+Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
+In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text.
+The central quantity of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
+Near dynamical critical regimes, the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates.
+In particular, an important limiting case is a finite infrared density of slow modes,
+$\rho(\lambda\rightarrow 0)=\rho_{0}\neq 0.$ (S131)
+This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales.
+As a result, the system cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
+(I) Long–time memory kernel: The dynamical kernel associated with the relaxation spectrum can be written as
+$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (S132)
+where $\Lambda$ denotes the ultraviolet cutoff of the relaxation spectrum.
+If the TDOS approaches a constant for $\lambda\ll\Lambda$, $\rho(\lambda)\approx\rho_{0}$, the long–time behavior becomes
+$K(t)\approx\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}=\frac{\rho_{0}}{t}\left(1-e^{-\Lambda t}\right).$ (S133)
+For $t\gg\Lambda^{-1}$ one therefore obtains the algebraic decay
+$K(t)\sim\frac{\rho_{0}}{t},$ (S134)
+indicating the emergence of long–time memory.
+(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex systems, including models of self–organized criticality (SOC).
+In many SOC models the lifetime distribution of events takes the approximate form
+$P(\tau)\sim\tau^{-2}.$ (S135)
+Using the relation between lifetime and relaxation rate $\lambda=1/\tau$, one finds
+$\rho(\lambda)=P(\tau)\left|\frac{d\tau}{d\lambda}\right|=P(1/\lambda)\frac{1}{\lambda^{2}}.$ (S136)
+Substituting $P(\tau)\sim\tau^{-2}$ yields
+$\rho(\lambda)\sim{\rm const}.$ (S137)
+Thus a flat infrared TDOS corresponds to a relaxation spectrum that is consistent with scale–free lifetime statistics.
+This observation suggests a possible dynamical connection between the TDOS framework and the emergence of SOC–type temporal scaling.
+(III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra.
+Within the present framework the noise spectrum can be written as
+$S(\omega)\sim\int d\lambda\,\frac{\rho(\lambda)\lambda}{\lambda^{2}+\omega^{2}}.$ (S138)
+When the infrared TDOS approaches a constant, the dominant contribution arises from $\lambda\lesssim\omega$, which yields approximately
+$S(\omega)\sim\frac{1}{|\omega|}.$ (S139)
+This provides a natural dynamical interpretation of the ubiquitous $1/f$ noise observed in many correlated systems.
+(IV) Relation to the infrared spectral function: As shown in Appendix E, a flat TDOS produces a characteristic logarithmic structure in the retarded susceptibility,
+$\chi_{R}^{-1}(\omega,q)=\Delta_{q}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,{\rm sgn}(\omega).$ (S140)
+The resulting spectral function
+$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}$ (S141)
+describes a broad continuum rather than a sharp quasiparticle peak.
+Such continua have been reported in several strongly correlated systems, including strange metals, although the precise microscopic origin remains an open question.
+(V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
+$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (S142)
+with
+$\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S143)
+For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one finds
+$\Pi(\omega)\sim\frac{1}{|\omega|},$ (S144)
+which leads to the algebraic transition scale
+$T_{c}\sim\rho_{0}E_{\rm pair}.$ (S145)
+The slow–mode reservoir therefore acts as a dynamical amplifier of intrinsic electronic pairing correlations.
+Summary.
+These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems.
+Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes.
+Further investigation of this spectral viewpoint may help clarify the role of dynamical criticality in a wide range of nonequilibrium and strongly correlated systems.

### v4

@@
 Fourier transforming the kernel yields the retarded self-energy,
-$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
+$\Sigma_{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
 so that the effective action becomes
-$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
+$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma_{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
 Separating real and imaginary parts,
-$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
-$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
+$\mathrm{Re}\,\Sigma_{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
+$\mathrm{Im}\,\Sigma_{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
 The collective inverse susceptibility therefore assumes the exact spectral representation
@@
 In this section we derive the infrared structure of the Cooper channel in this regime and contrast it with the conventional Eliashberg mechanism.
-III.1 A. Marginal pairing from overdamped collective dynamics
-Near metallic quantum critical points, electrons often couple to overdamped collective modes whose retarded propagator takes the form
-$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (25)
-Here $D_{R}$ denotes the effective propagator of a coarse-grained order-parameter (particle–hole) collective mode in the Hertz–Millis sense: the $|\omega|$ term is generated by the electron–hole polarization of the metallic continuum (Landau damping) [3, 4].
-This object should be distinguished from the Cooper-channel polarization $\Pi(\omega)$ introduced below, which diagnoses the pairing instability itself.
-The linear frequency dependence reflects Landau damping due to decay of collective excitations into the electronic continuum.
-After momentum integration, the pairing kernel entering the Cooper channel behaves approximately as
-$D_{R}(\omega)\sim\frac{1}{|\omega|}.$ (26)
-The linearized gap equation may be written schematically as
-$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
+III.1 A. Marginal pairing in overdamped bosonic-mediator theories
+As a reference point, it is useful to recall the structure of pairing in conventional theories based on overdamped collective mediators, such as spin-fluctuation approaches near metallic quantum criticality [25, 26].
+In such treatments one first introduces an effective bosonic propagator for a particle–hole collective mode,
+$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|,$ (25)
+where the $|\omega|$ term arises from Landau damping by the electronic continuum.
+This object describes the retarded susceptibility of an overdamped order-parameter fluctuation and should not be identified with the Cooper-channel polarization kernel itself.
+In bosonic-mediator theories [27], the overdamped collective mode is then treated as an effective pairing mediator.
+After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form
+$V(\omega)\sim\frac{1}{|\omega|},$ (26)
+up to model-dependent form factors and ultraviolet regularization.
+If one further assumes that pairing can still be analyzed within a quasiparticle-based ladder or gap-equation framework, the linearized pairing equation takes the schematic form
+$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,V(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
 where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff.
-To analyze the infrared flow we eliminate fast frequency modes with $\omega^{\prime}\in[\Lambda/b,\Lambda]$, corresponding to coarse–graining of short time scales.
-The ladder diagram gives the leading correction
-$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}.$ (28)
-The shell integral yields
+The corresponding infrared flow can be understood by integrating out fast frequency modes in the shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$. The leading ladder correction is then
+$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}},$ (28)
+which yields
 $\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln b.$ (29)
-Introducing the RG scale $l=\ln b$ gives
+Introducing the RG scale $l=\ln b$, one obtains the standard marginal flow
 $\frac{dg}{dl}=Cg^{2}.$ (30)
-The Cooper interaction therefore grows only logarithmically under infrared coarse graining.
-In RG language the pairing interaction is marginal.
-Physically, overdamped collective modes dissipate into the electronic continuum and repeatedly lose phase coherence.
-As a result superconducting correlations accumulate only logarithmically as the energy scale is reduced.
+Thus, within overdamped bosonic-mediator theories, pairing grows only logarithmically in the infrared [28, 29].
+The physical reason is that the mediator itself is continuously damped by decay into the electronic continuum, so coherence is not retained across an extended hierarchy of time scales.
+The resulting enhancement is therefore marginal rather than algebraic.
+This benchmark should be distinguished sharply from the framework developed in the present work.
+Here we do not assume a well-defined quasiparticle description nor an effective pairing glue represented by a single bosonic propagator.
+Instead, after integrating out microscopic electronic degrees of freedom, the infrared Cooper sector is organized by the collective relaxation spectrum itself.
+The relevant object is not the retarded propagator of an overdamped order-parameter mode, but the polarization kernel generated by a continuum of relaxational pair modes, as derived below.
 III.2 B. Collective relaxation spectrum and TDOS representation
@@
 $\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
-The polarization kernel entering the Cooper channel becomes
+While the quantity above governs the retarded response of the collective sector, the object relevant for superconducting pairing is the Cooper-channel fluctuation kernel.
+After integrating out microscopic electronic degrees of freedom, pairing correlations are encoded in the two-point correlator of the collective pair field
+$\langle\Delta(\omega)\Delta^{*}(\omega)\rangle.$ (32)
+To understand its structure, consider a single relaxational pair mode with decay rate $\lambda$. At long times its dynamics is well approximated by a linear stochastic equation
+$\dot{\Delta}_{\lambda}=-\lambda\,\Delta_{\lambda}+\xi_{\lambda},$ (33)
+where $\xi_{\lambda}$ denotes Gaussian noise generated by microscopic degrees of freedom. The stationary two-point correlation function of such a mode is
+$\langle\Delta_{\lambda}(t)\Delta_{\lambda}^{*}(0)\rangle\propto e^{-\lambda|t|}.$ (34)
+Fourier transforming this correlation function gives the fluctuation spectrum of a relaxational mode
+$S_{\lambda}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (35)
+The Cooper-channel polarization kernel therefore arises by summing the contributions of all relaxational pair modes weighted by the TDOS,
 $\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (36)
-Here $\Pi(\omega)$ represents the polarization of the Cooper channel generated by the collective relaxation spectrum.
-It should be distinguished from the conventional bosonic polarization that appears in propagator renormalization.
-In the present framework $\Pi(\omega)$ directly measures the dynamical pair susceptibility produced by the slow collective modes.
+Equation (36) represents the dynamical polarization of the Cooper channel generated by the collective relaxation spectrum.
+Importantly, this object differs from the reactive part of the retarded susceptibility discussed above.
+Rather than describing the propagation of a single bosonic mediator, it measures the fluctuation spectrum of the entire slow-mode reservoir.
 For a flat TDOS in the infrared,
@@
 $\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (38)
-In the time domain this corresponds to a long–memory kernel
+In the time domain this corresponds to a long-memory kernel
 $K(t)\sim\frac{1}{t}.$ (39)
-Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves.
+Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath.
+Instead it arises from the spectral accumulation of slow relaxation modes themselves.
+A complete derivation of Eq.
+(36) from the stochastic dynamics of relaxational pair modes is presented in Appendix D.
 III.3 C. Infrared amplification of pairing correlations
-The Cooper susceptibility takes the form
+In conventional superconductivity theories the transition temperature is often obtained from a microscopic gap equation derived from fermionic quasiparticle propagators.
+Such formulations rely on the existence of reasonably well-defined single-particle excitations, so that the Cooper kernel can be written in terms of products of fermionic Green functions.
+In strongly correlated systems, however, the quasiparticle picture may cease to be valid.
+The single-particle spectral function can become broad and incoherent, and the low-energy dynamics may instead be governed by a continuum of collective relaxation processes.
+Within the present framework the infrared dynamics is therefore organized in terms of the relaxation-rate spectrum $\rho(\lambda)$ rather than quasiparticle propagation.
+Because the microscopic quasiparticle propagator is not introduced explicitly, the superconducting transition is identified through the instability of the Cooper-channel susceptibility.
+The corresponding Thouless criterion reads
+$\chi_{\Delta}^{-1}(0,T_{c})=0,\quad\text{or equivalently}\quad 1-g_{\rm pair}\Pi(T_{c})=0.$ (40)
+Here $g_{\rm pair}$ represents the intrinsic electronic interaction that seeds Cooper correlations, while the relaxation spectrum controls their dynamical amplification through the infrared structure of $\Pi(\omega)$.
+The Cooper susceptibility therefore takes the form
 $\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (41)
@@
 $\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (43)
-The effective pairing kernel therefore scales as
-$g_{\mathrm{eff}}(\omega)\sim|\omega|^{-(1-\eta)}.$ (39)
-Under RG rescaling $\omega\rightarrow\omega/b$ the kernel transforms as
+The Cooper-channel kernel therefore acquires the infrared scaling
+$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (44)
+Under RG rescaling $\omega\rightarrow\omega/b$ the effective pairing interaction transforms as
 $g^{\prime}=b^{1-\eta}g.$ (45)
@@
 Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue).
-In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange).
+In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=(1-\eta)g$, leading to algebraic amplification of the pairing interaction (orange).
 The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
@@
 IV.1 A. Intrinsic electronic pairing and dynamical formulation
-In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes.
+In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes [1, 30, 31].
 Starting from the Hubbard Hamiltonian,
@@
 Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
-Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
-$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (48)
-where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations.
-Diagrammatically $\Pi(\omega)$ corresponds to the Cooper bubble, which resums particle–particle ladder diagrams.
+Within the dynamical framework developed in Sec.
+III, the pairing susceptibility assumes the universal tt–matrix form in Eq.
+(41), where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations.
+Diagrammatically this corresponds to the Cooper-channel polarization obtained after integrating out microscopic electronic degrees of freedom.
 In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory.
@@
 IV.2 B. Memory–dominated Cooper kernel from flat TDOS
-Because single–particle propagation inherits the relaxational spectral structure generated by the slow–mode reservoir, the Cooper bubble reduces to a continuum integral over relaxation rates,
+Because the collective pair fluctuations inherit the relaxational spectral structure generated by the slow-mode reservoir, the Cooper-channel polarization kernel reduces to a continuum integral over relaxation rates,
 $\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (53)
@@
 IV.3 C. Algebraic instability and transition scale
+At finite temperature the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$.
+As a consequence the system cannot probe arbitrarily long time scales.
+In the imaginary-time formalism the dynamical fields are expanded in bosonic Matsubara frequencies
+$\Omega_{n}=2\pi nT.$ (56)
+The smallest nonzero dynamical frequency therefore sets a natural infrared scale of order
+$|\omega|\sim T.$ (57)
+Substituting this scale into the infrared expression (55) gives
+$\Pi(T)\sim\frac{\rho_{0}}{T}.$ (58)
+Thus the infrared divergence present at zero temperature is universally regulated by the thermal scale.
 The superconducting transition follows from the Thouless criterion
@@
 $T_{c}\sim\rho_{0}\,E_{\rm pair},$ (60)
-up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
+up to numerical factors of order unity.
 The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir.
@@
 Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight.
-Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
+Integrating out amplitude fluctuations in the presence of the slow-mode reservoir generically yields
 $\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (69)
@@
 corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
-Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors.
+Such long-time tails naturally produce 1/f-type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors.
 Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure.
@@
 The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
-Figure 4: Universality map relating the relaxation-rate spectrum to infrared collective dynamics and superconducting pairing.
-The TDOS exponent $\alpha$ determines the scaling of the polarization kernel $\Pi(\omega)\sim|\omega|^{\alpha-1}$.
-The Ohmic Hertz–Millis theory corresponds to the marginal point $\alpha=1$, while a flat TDOS ($\alpha=0$) produces memory-dominated critical dynamics with $\Pi(\omega)\sim 1/|\omega|$ and enhanced pairing.
+Figure 4: Dynamical classification determined by the TDOS exponent $\alpha$.
+The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi(\omega)\sim|\omega|^{\alpha-1}$ and separates the system into memory–dominated ($\alpha<1$), marginal Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences.
+Here $\omega$ denotes Matsubara frequency.
 VI VI. Discussion
@@
 The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states.
-For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form
-$\Pi(\omega)\sim|\omega|^{\alpha-1}.$ (65)
+For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form $\Pi(\omega)\sim|\omega|^{\alpha-1}$.
 Different dynamical regimes therefore emerge as special cases of the same spectral organization.
-For $\alpha>1$ the density of slow modes vanishes in the infrared, so that $\Pi(\omega)\rightarrow 0$ and the collective dynamics reduces to weakly damped propagating modes with $\chi^{-1}\simeq r+cq^{2}$.
+For $\alpha>1$ the density of slow modes vanishes in the infrared, so that the polarization kernel remains finite and the collective dynamics approaches the weakly damped limit, with the infrared response dominated by the nondissipative part $\chi^{-1}\simeq r+cq^{2}$.
 The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$.
@@
 In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
+Interestingly, the infrared response generated by a flat TDOS is structurally identical to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 32].
+In particular, the resulting spectrum corresponds in frequency space to a logarithmic real part and a nearly constant absorptive component of the susceptibility.
+Similar infrared structures were introduced phenomenologically in the marginal Fermi liquid theory of strange metals and also arise in solvable strongly correlated spin models such as the Sachdev–Ye spin liquid.
+In the present framework, however, this behavior follows directly from the accumulation of slow relaxation modes and therefore emerges naturally as a consequence of a finite infrared relaxation-rate density of states.
 The lower panel of Fig.
 4 illustrates the direct consequence of this spectral organization for superconductivity.
-Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic kernel $\Pi(T)\sim\ln(\Omega/T)$ and the exponentially small BCS transition temperature.
+Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic Cooper kernel $\Pi(T)\sim\ln(\Omega/T)$ and hence to the exponentially small transition scale characteristic of conventional marginal pairing theories.
 In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$.
@@
 The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
-In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
+In this sense, superconducting order is governed by condensation in a Cooper sector embedded within an extended slow-mode manifold rather than by binding through a sharp bosonic excitation.
 A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena.
@@
 Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra.
-A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations rather than a model-specific constructione, is derived in Appendix F.
+A dense TDOS therefore reflects a generic mathematical feature of dynamical systems with extended temporal correlations rather than a model-specific construction.
+Related derivations are presented in Appendix F.
 This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics.
@@
 When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales.
-The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and logarithmic infrared response functions.
+The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions, with logarithmic or stronger singular structures depending on the observable sector.
 Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality.
-This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena.
+This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena [33, 34].
 Memory–dominated dynamics as a guiding principle for room-temperature superconductivity.
@@
 Within the present framework, however, enhancing the bare pairing scale alone is insufficient.
-The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
+A decisive control parameter for high transition temperatures may be the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
 A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$.
@@
 Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself.
-In this perspective high-temperature superconductivity corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
+In this perspective high-temperature superconductivity is most naturally favored in regimes characterized by maximal infrared memory rather than simply by extreme coupling strength.
 VII VII. Conclusion
@@
 Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
-In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter.
+In this perspective, high–temperature superconductivity can be viewed as a generic consequence of favorable infrared spectral organization in strongly correlated matter.
 Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
@@
 (IX) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
-$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
+$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
 Coupling $\phi$ locally to an OU reservoir $X_{\lambda}(\mathbf{r},t)$ and integrating out $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the nonlocal-in-time term
@@
 $\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),$ (S66)
-up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response.
+up to the conventional $i\omega$ term of the retarded response.
 Summary of the exact Gaussian result.
@@
 and substituting into (S72) gives
-$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S91)
+$\Big[-i\omega+r+\Sigma_{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S91)
 with the retarded self-energy
-$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S92)
-We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
+$\Sigma_{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S92)
+We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
 Separating real and imaginary parts using
@@
 one obtains
-$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S94)
-$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S95)
+$\mathrm{Re}\,\Sigma_{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S94)
+$\mathrm{Im}\,\Sigma_{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S95)
 For a flat TDOS $\rho(\lambda)\simeq\rho_{0}$ and $|\omega|\ll\Lambda$,
-$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
+$\mathrm{Im}\,\Sigma_{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
 $\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (S96)
 and
-$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
+$\mathrm{Re}\,\Sigma_{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
 $=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
@@
 The kernel is the Laplace transform of the TDOS.
-A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part.
+A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma_{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part.
 This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
@@
 In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
-Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
-In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$.
-This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
-(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
-$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S101)
-[…]
-$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S102)
-The elementary integral yields
-$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S103)
-In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
-$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S104)
-which expresses the nonanalytic infrared enhancement at zero temperature.
-(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies.
-As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
-Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S102) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
-$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S105)
-where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
-Evaluating the integral gives
-$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S106)
-Thus the infrared divergence is universally cut off by the thermal scale,
-$\Pi(T)\propto\frac{1}{T}.$ (S107)
-(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
-$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S108)
-Using the standard identity
-$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S109)
-and expanding for $\lambda\ll T$,
-$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S110)
+Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
+In this Appendix we present an explicit derivation of the polarization kernel used in the Cooper channel within the relaxation-spectrum framework.
+The Cooper-channel object relevant for pairing enhancement is the fluctuation (or correlator) spectrum of the collective pair field.
+The construction below should be understood, in the same spirit as Appendix B, as an effective infrared description after microscopic electronic degrees of freedom have already been integrated out.
+At that stage the relevant dynamical variables are collective pair modes, and the polarization kernel is obtained from their stochastic two-point correlations rather than from an explicit microscopic fermionic loop.
+(I) Single relaxational pair mode: Consider a single complex collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$ obeying the linear Langevin equation
+$\partial_{t}\Delta_{\lambda}(t)=-\lambda\,\Delta_{\lambda}(t)+\xi_{\lambda}(t),$ (S101)
+where $\xi_{\lambda}(t)$ is Gaussian white noise with
+$\langle\xi_{\lambda}(t)\rangle=0,\qquad\langle\xi_{\lambda}(t)\xi_{\lambda^{\prime}}^{*}(t^{\prime})\rangle=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(t-t^{\prime}).$ (S102)
+Here $D_{\lambda}$ denotes the noise strength associated with the mode $\lambda$.
+In the stationary regime, obtained by sending the initial time to -$\infty$, the exact solution of Eq. (S101) is
+$\Delta_{\lambda}(t)=\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s).$ (S103)
+Differentiating Eq. (S103) directly yields
+$\partial_{t}\Delta_{\lambda}(t)$ $=\xi_{\lambda}(t)-\lambda\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s)$
+$=\xi_{\lambda}(t)-\lambda\Delta_{\lambda}(t),$ (S104)
+which confirms Eq. (S101).
+(II) Time-domain two-point correlation function: We now compute the stationary correlator
+$C_{\lambda}(\tau)\equiv\langle\Delta_{\lambda}(\tau)\Delta_{\lambda}^{*}(0)\rangle.$ (S105)
+For $\tau>0$, inserting Eq. (S103) gives
+$C_{\lambda}(\tau)$ $=\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\,\langle\xi_{\lambda}(s)\xi_{\lambda}^{*}(s^{\prime})\rangle$
+$=2D_{\lambda}\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\delta(s-s^{\prime}).$ (S106)
+Because the delta function enforces $s=s^{\prime}$, while the second integral restricts $s^{\prime}\leq 0$, the effective integration domain reduces to $s\leq 0$. Therefore
+$C_{\lambda}(\tau)$ $=2D_{\lambda}\int_{-\infty}^{0}ds\,e^{-\lambda(\tau-s)}e^{\lambda s}$
+$=2D_{\lambda}e^{-\lambda\tau}\int_{-\infty}^{0}ds\,e^{2\lambda s}$
+$=2D_{\lambda}e^{-\lambda\tau}\left[\frac{e^{2\lambda s}}{2\lambda}\right]_{-\infty}^{0}$
+$=\frac{D_{\lambda}}{\lambda}e^{-\lambda\tau}.$ (S107)
+By stationarity and Hermitian symmetry, the result extends to all times as
+$C_{\lambda}(t)=\frac{D_{\lambda}}{\lambda}\,e^{-\lambda|t|}.$ (S108)
+Equation (S108) is the exact two-point correlation function of a single Ornstein–Uhlenbeck pair mode.
+The factor $1/\lambda$ in front of the exponential is crucial.
+It is precisely this factor that removes the extra numerator $\lambda$ appearing in the Fourier transform of the bare retarded kernel.
+(III) Frequency-domain spectral density: The mode-resolved fluctuation spectrum is defined as
+$S_{\lambda}(\omega)\equiv\int_{-\infty}^{\infty}dt\,e^{i\omega t}C_{\lambda}(t).$ (S109)
+Substituting Eq. (S108) gives
+$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\int_{-\infty}^{\infty}dt\,e^{i\omega t}e^{-\lambda|t|}.$ (S110)
+Splitting the integral into positive and negative times,
+$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}\right].$ (S111)
+Changing variables u=tu=-t in the second term yields
+$\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}=\int_{0}^{\infty}du\,e^{-(\lambda+i\omega)u},$ (S112)
+so that
+$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{0}^{\infty}dt\,e^{-(\lambda+i\omega)t}\right].$ (S113)
+Using
+$\int_{0}^{\infty}dt\,e^{-at}=\frac{1}{a},\qquad\Re a>0,$ (S114)
+we obtain
+$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left(\frac{1}{\lambda-i\omega}+\frac{1}{\lambda+i\omega}\right)$
+$=\frac{D_{\lambda}}{\lambda}\frac{2\lambda}{\lambda^{2}+\omega^{2}}.$ (S115)
+Hence
+$S_{\lambda}(\omega)=\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S116)
+This is the key result.
+The Fourier transform of the exponential memory kernel itself produces a numerator $\lambda$, but the stationary correlator contains a compensating factor $1/\lambda$.
+As a result, the exact fluctuation spectrum carries no additional factor of $\lambda$ in the numerator.
+(IV) Equivalent derivation in frequency space: The same result follows directly in frequency space. Fourier transforming Eq. (S101) gives
+$(-i\omega+\lambda)\Delta_{\lambda}(\omega)=\xi_{\lambda}(\omega),$ (S117)
+hence
+$\Delta_{\lambda}(\omega)=\frac{1}{\lambda-i\omega}\,\xi_{\lambda}(\omega).$ (S118)
+The noise correlator in frequency space is
+$\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle=2\pi\,2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime}).$ (S119)
+Therefore
+$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle$ $=\frac{\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle}{(\lambda-i\omega)(\lambda^{\prime}+i\omega^{\prime})}$
+$=2\pi\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime})\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S120)
+Defining the spectral density through
+$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda}^{*}(\omega^{\prime})\rangle=2\pi\delta(\omega-\omega^{\prime})S_{\lambda}(\omega),$ (S121)
+one again recovers Eq. (S116).
+This derivation makes clear that the mode-resolved pair correlator is proportional to $D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)$, not to the real part of a single retarded propagator. Indeed,
+$D_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad D_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega},$ (S122)
+so that
+$D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)=\frac{1}{\lambda^{2}+\omega^{2}}.$ (S123)
+(V) From a single mode to the Cooper-channel polarization: After microscopic electronic degrees of freedom are integrated out, the collective Cooper sector is represented as a continuum of relaxational pair modes.
+Let $\rho_{\rm pair}(\lambda)$ denote the effective density of pair relaxation rates, including both mode degeneracy and noise weight.
+The full polarization kernel is then obtained by summing the mode-resolved fluctuation spectrum over the continuum,
+$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,S_{\lambda}(\omega).$ (S124)
+Absorbing the factor $2D_{\lambda}$ into the definition of $\rho_{\rm pair}(\lambda)$ gives
+$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S125)
+This is the Cooper-channel polarization kernel quoted in the main text.
+(VI) Infrared scaling of the polarization kernel: For a generic infrared TDOS
+$\rho_{\rm pair}(\lambda)\sim\lambda^{\eta},\qquad\lambda\to 0,$ (S126)
+Eq. (S125) gives
+$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}.$ (S127)
+Setting $\lambda=|\omega|x$,
+$\Pi(\omega)$ $\sim|\omega|^{\eta-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\eta}}{1+x^{2}}.$ (S128)
+Hence the universal infrared scaling is
+$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (S129)
+In particular, for a flat TDOS,
+$\rho_{\rm pair}(\lambda\to 0)=\rho_{0},$ (S130)
 one obtains
-$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S111)
-Substitution into Eq. (S108) yields
-$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S112)
-confirming that finite temperature universally regulates the infrared divergence.
-(IV) Implication for the superconducting transition: The Thouless criterion,
-$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S113)
-therefore gives
-$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S114)
-or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale.
-The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
+$\Pi(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}$
+$=\frac{\rho_{0}}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right)$
+$\xrightarrow[|\omega|\ll\Lambda]{}\frac{\pi\rho_{0}}{2}\frac{1}{|\omega|}.$ (S131)
+Equation (S131) establishes the memory-dominated Cooper enhancement.
+A flat relaxation-rate spectrum therefore converts the marginal logarithmic pairing growth of conventional Ohmic dynamics into an algebraically enhanced infrared instability of the Cooper channel.
 Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics.
@@
 In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text.
-The central quantity of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
-Near dynamical critical regimes, the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates.
-In particular, an important limiting case is a finite infrared density of slow modes,
+The central object of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
+Near dynamical critical regimes the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates.
+In particular an important limiting case is a finite infrared density of slow modes,
 $\rho(\lambda\rightarrow 0)=\rho_{0}\neq 0.$ (S148)
 This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales.
-As a result, the system cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
-(I) Long–time memory kernel: The dynamical kernel associated with the relaxation spectrum can be written as
+The system therefore cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
+(I) Long–time memory kernel: The dynamical memory kernel associated with the relaxation spectrum can be written as
 $K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (S149)
 where $\Lambda$ denotes the ultraviolet cutoff of the relaxation spectrum.
-If the TDOS approaches a constant for $\lambda\ll\Lambda$, $\rho(\lambda)\approx\rho_{0}$, the long–time behavior becomes
+If the TDOS approaches a constant in the infrared, $\rho(\lambda)\approx\rho_{0}$ for $\lambda\ll\Lambda$, one obtains
 $K(t)\approx\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}=\frac{\rho_{0}}{t}\left(1-e^{-\Lambda t}\right).$ (S150)
-For $t\gg\Lambda^{-1}$ one therefore obtains the algebraic decay
+For long times $t\gg\Lambda^{-1}$ this yields the algebraic decay
 $K(t)\sim\frac{\rho_{0}}{t},$ (S151)
 indicating the emergence of long–time memory.
-(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex systems, including models of self–organized criticality (SOC).
+(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex dynamical systems, including models of self–organized criticality (SOC).
 In many SOC models the lifetime distribution of events takes the approximate form
 $P(\tau)\sim\tau^{-2}.$ (S152)
-Using the relation between lifetime and relaxation rate $\lambda=1/\tau$, one finds
-$\rho(\lambda)=P(\tau)\left|\frac{d\tau}{d\lambda}\right|=P(1/\lambda)\frac{1}{\lambda^{2}}.$ (S136)
-Substituting $P(\tau)\sim\tau^{-2}$ yields
-$\rho(\lambda)\sim{\rm const}.$ (S137)
-Thus a flat infrared TDOS corresponds to a relaxation spectrum that is consistent with scale–free lifetime statistics.
-This observation suggests a possible dynamical connection between the TDOS framework and the emergence of SOC–type temporal scaling.
+The relation between lifetime $\tau$ and relaxation rate $\lambda$ is
+$\tau=\lambda^{-1}.$ (S153)
+Since the number of modes must be invariant under a change of variables, the densities satisfy
+$\rho(\lambda)\,d\lambda=P(\tau)\,d\tau.$ (S154)
+Using
+$\lambda=\tau^{-1},\qquad\left|\frac{d\lambda}{d\tau}\right|=\tau^{-2},$ (S155)
+one obtains
+$P(\tau)=\rho(\tau^{-1})\,\tau^{-2}.$ (S156)
+Thus a flat infrared TDOS, $\rho(\lambda\rightarrow 0)=\rho_{0}$, implies in Eq.
+(S152).
+A scale–free lifetime distribution therefore corresponds to a relaxation spectrum that is approximately constant near $\lambda=0$.
 (III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra.
-Within the present framework the noise spectrum can be written as
-$S(\omega)\sim\int d\lambda\,\frac{\rho(\lambda)\lambda}{\lambda^{2}+\omega^{2}}.$ (S138)
-When the infrared TDOS approaches a constant, the dominant contribution arises from $\lambda\lesssim\omega$, which yields approximately
+Within the present framework the power spectrum associated with a superposition of relaxational modes takes the form
+$S(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S157)
+Changing variables to the lifetime $\tau=1/\lambda$ and using Eq. (S156) gives
+$S(\omega)\sim\int_{0}^{\infty}d\tau\,P(\tau)\,\frac{\tau^{2}}{1+(\omega\tau)^{2}}.$ (S158)
+For a scale–free distribution $P(\tau)\sim\tau^{-2}$ one finds
+$S(\omega)\sim\int d\tau\frac{1}{1+(\omega\tau)^{2}}.$ (S159)
+Introducing the dimensionless variable $x=\omega\tau$ yields
+$S(\omega)\sim\frac{1}{\omega}\int dx\,\frac{1}{1+x^{2}}.$ (S160)
+Since the remaining integral is finite, the low–frequency scaling is
 $S(\omega)\sim\frac{1}{|\omega|}.$ (S161)
-This provides a natural dynamical interpretation of the ubiquitous $1/f$ noise observed in many correlated systems.
+Thus a flat TDOS naturally produces a 1/f1/f–type noise spectrum.
 (IV) Relation to the infrared spectral function: As shown in Appendix E, a flat TDOS produces a characteristic logarithmic structure in the retarded susceptibility,
 $\chi_{R}^{-1}(\omega,q)=\Delta_{q}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,{\rm sgn}(\omega).$ (S162)
-The resulting spectral function
-$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}$ (S141)
+The corresponding spectral function,
+$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}},$ (S163)
 describes a broad continuum rather than a sharp quasiparticle peak.
-Such continua have been reported in several strongly correlated systems, including strange metals, although the precise microscopic origin remains an open question.
 (V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
@@
 $\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S165)
-For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one finds
+For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one obtains
 $\Pi(\omega)\sim\frac{1}{|\omega|},$ (S166)
-which leads to the algebraic transition scale
+which leads to an algebraically enhanced transition scale
 $T_{c}\sim\rho_{0}E_{\rm pair}.$ (S167)
@@
 Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes.
-Further investigation of this spectral viewpoint may help clarify the role of dynamical criticality in a wide range of nonequilibrium and strongly correlated systems.
+Flat infrared TDOS simultaneously produces long–time memory, scale–free lifetime statistics, $1/f$ noise, and algebraic enhancement of superconducting pairing.

### v5

@@
 Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4].
-This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics.
-While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
+This framework, formalized in Hertz–Millis theory, predicts an approximately universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics.
+While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are widely observed [5, 6, 7, 8, 9, 10].
 These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
@@
 In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels.
-By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
+Starting from the real–time Keldysh functional integral formulation of dissipative many-body dynamics [16, 17, 18], and passing to its classical stochastic infrared limit described by the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [19, 20, 21], we construct a dynamical coarse–graining theory in which the collective susceptibility admits an exact spectral representation in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
 We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior.
 Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility.
-While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to the case $\alpha=1$, arising microscopically from particle–hole polarization of a Fermi liquid, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic kernel proportional to $\mathrm{sgn}(\omega)$.
+While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to an effective TDOS scaling with $\alpha=1$, a flat TDOS ($\alpha=0$) instead produces a marginal nonanalytic response characterized by logarithmic frequency dependence and long-time memory.
 In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$.
@@
 Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [22, 23, 24].
-The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
+The resulting long-memory kernel enhances intrinsic electronic pairing tendencies, establishing a natural dynamical environment for strengthened pairing without relying on finely tuned bosonic mediators.
+Rather than arising from the formation of tightly bound Cooper pairs, superconductivity in this framework is understood as a collective dynamical phenomenon.
+A weak intrinsic pairing tendency generates fluctuating pairs, whose instability is driven by the accumulation of slow collective modes in the relaxation spectrum.
 Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter.
-The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22].
+The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing tendencies, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 25, 26].
 Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
-Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity.
+Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a natural dynamical environment for enhanced pairing and superconducting instabilities.
 This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 11].
@@
 In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies.
-We conclude in Section VI with a summary and outlook.
+In Section VI we place these results in a broader conceptual framework, discussing their implications for dynamical criticality, Liouvillian descriptions of collective dynamics, and routes toward enhanced superconductivity, before concluding with a summary and outlook.
 Figure 1: Dynamical universality map of correlated quantum matter.
 (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes.
-Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality.
-(b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory.
-(c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
+Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, a flat TDOS defines memory–dominated criticality, and singular infrared spectra ($\alpha<0$) correspond to a dense accumulation of slow modes.
+(b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory, with $K(t)\sim 1/t$ for the marginal (flat TDOS) case and slower decay for IR-singular spectra.
+(c) Physical consequences of the TDOS-controlled dynamics.
+Memory–dominated criticality produces scale-free dynamics, long-time correlations, and logarithmic enhancement of the retarded pairing kernel, while additional infrared accumulation of slow modes ($\alpha<0$) leads to power-law singularities and algebraic enhancement of superconducting pairing.
 II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
@@
 $\mathcal{L}R_{n}=\lambda_{n}R_{n},$ (2)
-define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
-$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
+define the fundamental decay channels of the interacting system.
+For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as $\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0$.
 Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
@@
 To compute collective response functions, we employ a functional representation of the dissipative dynamics.
-The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24].
+The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework.
 Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
@@
 $\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (5)
-$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
+$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\xi_{\lambda},$ (6)
 where $X_{\lambda}$ represents a decay mode with relaxation rate $\lambda$, and $g(\lambda)$ encodes its overlap with $\phi$.
@@
 Equation (15) provides an exact infrared response fully controlled by the relaxation-rate density.
-The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
+The slow-mode reservoir thus appears as a memory-dominated dynamical regime, whose long-time kernel structure governs the infrared dynamical behavior discussed below.
 The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
@@
 In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”.
-In contrast, in the present framework superconductivity emerges from the dynamical amplification of pairing correlations generated by the slow–mode reservoir itself.
+In contrast, in the present framework superconductivity emerges from the infrared structure of pairing correlations reorganized by the slow-mode reservoir.
 In this section we derive the infrared structure of the Cooper channel in this regime and contrast it with the conventional Eliashberg mechanism.
@@
 $D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|,$ (24)
-where the $|\omega|$ term arises from Landau damping by the electronic continuum.
-This object describes the retarded susceptibility of an overdamped order-parameter fluctuation and should not be identified with the Cooper-channel polarization kernel itself.
+where the $|\omega|$ term arises from Landau damping by the electronic continuum. This object describes the retarded susceptibility of an overdamped order-parameter fluctuation.
 In bosonic-mediator theories [29], the overdamped collective mode is then treated as an effective pairing mediator.
-After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form
-$V(\omega)\sim\frac{1}{|\omega|},$ (26)
-up to model-dependent form factors and ultraviolet regularization.
+After momentum integration, the resulting interaction kernel entering the Cooper channel is schematically of the form $V(\omega)\sim 1/|\omega|$, up to model-dependent form factors and ultraviolet regularization.
 If one further assumes that pairing can still be analyzed within a quasiparticle-based ladder or gap-equation framework, the linearized pairing equation takes the schematic form
@@
 Instead, after integrating out microscopic electronic degrees of freedom, the infrared Cooper sector is organized by the collective relaxation spectrum itself.
-The relevant object is not the retarded propagator of an overdamped order-parameter mode, but the polarization kernel generated by a continuum of relaxational pair modes, as derived below.
-III.2 B. Collective relaxation spectrum and TDOS representation
-Strongly correlated quantum matter, however, can develop a much richer infrared structure.
-Instead of a single overdamped bosonic excitation, the system may host a dense continuum of slow collective modes.
-The collective dynamics can then be represented by the TDOS $\rho(\lambda)$, which measures the density of relaxation rates $\lambda$.
-Projecting the dynamics onto the slow collective sector yields the spectral representation
-$\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
-While the quantity above governs the retarded response of the collective sector, the object relevant for superconducting pairing is the Cooper-channel fluctuation kernel.
-After integrating out microscopic electronic degrees of freedom, pairing correlations are encoded in the two-point correlator of the collective pair field
-$\langle\Delta(\omega)\Delta^{*}(\omega)\rangle.$ (32)
-To understand its structure, consider a single relaxational pair mode with decay rate $\lambda$. At long times its dynamics is well approximated by a linear stochastic equation
-$\dot{\Delta}_{\lambda}=-\lambda\,\Delta_{\lambda}+\xi_{\lambda},$ (33)
-where $\xi_{\lambda}$ denotes Gaussian noise generated by microscopic degrees of freedom. The stationary two-point correlation function of such a mode is
-$\langle\Delta_{\lambda}(t)\Delta_{\lambda}^{*}(0)\rangle\propto e^{-\lambda|t|}.$ (34)
-Fourier transforming this correlation function gives the fluctuation spectrum of a relaxational mode
-$S_{\lambda}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (35)
-The Cooper-channel polarization kernel therefore arises by summing the contributions of all relaxational pair modes weighted by the TDOS,
-$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (36)
-Equation (36) represents the dynamical polarization of the Cooper channel generated by the collective relaxation spectrum.
-Importantly, this object differs from the reactive part of the retarded susceptibility discussed above.
-Rather than describing the propagation of a single bosonic mediator, it measures the fluctuation spectrum of the entire slow-mode reservoir.
-For a flat TDOS in the infrared,
-$\rho(\lambda\to 0)=\rho_{0},$ (37)
-the polarization acquires a universal singular form
-$\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (38)
-In the time domain this corresponds to a long-memory kernel
-$K(t)\sim\frac{1}{t}.$ (39)
-Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath.
-Instead it arises from the spectral accumulation of slow relaxation modes themselves.
-A complete derivation of Eq.
-(36) from the stochastic dynamics of relaxational pair modes is presented in Appendix D.
-III.3 C. Infrared amplification of pairing correlations
-In conventional superconductivity theories the transition temperature is often obtained from a microscopic gap equation derived from fermionic quasiparticle propagators.
-Such formulations rely on the existence of reasonably well-defined single-particle excitations, so that the Cooper kernel can be written in terms of products of fermionic Green functions.
-In strongly correlated systems, however, the quasiparticle picture may cease to be valid.
-The single-particle spectral function can become broad and incoherent, and the low-energy dynamics may instead be governed by a continuum of collective relaxation processes.
-Within the present framework the infrared dynamics is therefore organized in terms of the relaxation-rate spectrum $\rho(\lambda)$ rather than quasiparticle propagation.
-Because the microscopic quasiparticle propagator is not introduced explicitly, the superconducting transition is identified through the instability of the Cooper-channel susceptibility.
-The corresponding Thouless criterion reads
-$\chi_{\Delta}^{-1}(0,T_{c})=0,\quad\text{or equivalently}\quad 1-g_{\rm pair}\Pi(T_{c})=0.$ (40)
-Here $g_{\rm pair}$ represents the intrinsic electronic interaction that seeds Cooper correlations, while the relaxation spectrum controls their dynamical amplification through the infrared structure of $\Pi(\omega)$.
-The Cooper susceptibility therefore takes the form
-$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (41)
-Because $\Pi(\omega)$ diverges as $|\omega|\to 0$, pairing correlations are strongly enhanced as the energy scale is lowered.
-To understand the infrared flow more generally, consider a TDOS scaling
-$\rho(\lambda)\sim\lambda^{\eta}.$ (42)
-Substituting this form into the polarization integral yields
-$\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (43)
-The Cooper-channel kernel therefore acquires the infrared scaling
-$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (44)
-Under RG rescaling $\omega\rightarrow\omega/b$ the effective pairing interaction transforms as
-$g^{\prime}=b^{1-\eta}g.$ (45)
-Introducing $l=\ln b$ gives the tree–level RG flow
-$\frac{dg}{dl}=(1-\eta)g.$ (46)
-For $\eta<1$ the Cooper interaction becomes relevant in the infrared. Including the usual ladder correction yields
-$\frac{dg}{dl}=(1-\eta)g+Cg^{2}.$ (47)
-The linear term dominates the early RG flow, leading to algebraic growth
-$g(l)=g_{0}e^{(1-\eta)l}.$ (48)
-Physical interpretation.
-The qualitative difference between the Eliashberg regime and the present memory–dominated regime originates from the structure of the collective spectrum, as illustrated schematically in Fig.
+The relevant object is not the retarded propagator of an overdamped order-parameter mode, but by the retarded polarization kernel generated by a continuum of relaxational pair modes, as derived below.
+III.2 B. Retarded Cooper kernel and ladder resummation in the relaxation-spectrum framework
+In strongly correlated systems, the Cooper channel must be formulated directly in terms of collective response rather than quasiparticle propagators.
+After microscopic fermions are integrated out, the infrared dynamics of pairing is governed by the relaxation spectrum introduced in Sec.
+II.
+The appropriate object controlling superconducting instability is the retarded response function of the collective pair field,
+$L^{R}(\omega)=\frac{1}{\lambda-i\omega}\quad\text{(single relaxational mode)}.$ (29)
+For a continuum of relaxational modes characterized by the TDOS $\rho(\lambda)$, the retarded Cooper-channel kernel is obtained by integrating over the spectrum,
+$\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (30)
+This expression has the same spectral structure as the retarded self-energy derived in Sec.
+II, reflecting that both quantities are governed by the same underlying relaxation spectrum, although they correspond to distinct physical responses.
+This kernel represents the linear response of the collective pair field and plays the role of the elementary rung in the Cooper ladder.
+The full pair susceptibility is obtained by resumming repeated pair scattering processes,
+$\chi_{\Delta}^{R}(\omega)=\Pi^{R}(\omega)+\Pi^{R}(\omega)\,g_{\rm pair}\,\chi_{\Delta}^{R}(\omega),$ (31)
+which yields
+$\chi_{\Delta}^{R}(\omega)=\frac{\Pi^{R}(\omega)}{1-g_{\rm pair}\Pi^{R}(\omega)},$ (32)
+where $g_{\rm pair}$ denotes the intrinsic electronic interaction that seeds pairing.
+Superconductivity corresponds to a pole of the retarded susceptibility, and the Cooper instability is determined by the Thouless criterion,
+$1-g_{\rm pair}\,\mathrm{Re}\,\Pi^{R}(\omega)=0.$ (33)
+Using Eq. (30), the real part of the kernel is
+$\mathrm{Re}\,\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\;\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}.$ (34)
+For a power-law TDOS, $\rho(\lambda)\sim\lambda^{\alpha}$, the infrared behavior of the retarded kernel depends on the sign of $\alpha$. In the infrared-singular regime $\alpha<0$, one finds
+$\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}\int_{0}^{\infty}dx\,\frac{x^{\alpha+1}}{x^{2}+1}=C_{\alpha}|\omega|^{\alpha},$ (35)
+where $C_{\alpha}$ is a positive finite constant. This nonanalytic power-law behavior directly determines the structure of the superconducting instability.
+The fluctuation spectrum of pairing correlations is encoded in the Keldysh component,
+$L^{K}(\omega)=L^{R}(\omega)\,\Sigma^{K}(\omega)\,L^{A}(\omega),$ (36)
+which captures infrared enhancement of fluctuations and long-time correlations. However, the superconducting instability itself is controlled by the retarded response through Eq. (33).
+Thus, within the relaxation-spectrum framework, superconducting instability is governed by the structure of the retarded kernel, while fluctuation correlators describe the accompanying dynamical noise and non-Markovian correlations.
+A detailed derivation of the response and fluctuation kernels is presented in Appendix D.
+III.3 C. Infrared scaling and renormalization of the Cooper channel
+The superconducting instability is determined by the Thouless criterion derived in Sec.
+III B.
+The infrared behavior of pairing is therefore controlled by the low-frequency scaling of the retarded kernel $\Pi^{R}(\omega)$, which is in turn determined by the structure of the relaxation spectrum.
+In the flat-TDOS regime ($\alpha=0$), the kernel takes the form
+$\Pi^{R}(\omega)\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|},$ (37)
+so that coarse graining over a logarithmic frequency shell $\omega\in[\Lambda/b,\Lambda]$ produces
+$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega}{\omega}=g^{2}\ln b,$ (38)
+leading to the marginal flow $dg/dl=Cg^{2}$. Thus, a flat TDOS yields a logarithmic infrared enhancement of pairing, analogous to marginal behavior in conventional Eliashberg theories.
+When the TDOS develops an infrared divergence ($\alpha<0$), the retarded kernel acquires a power-law form,
+$\Pi^{R}(\omega)\sim|\omega|^{\alpha}.$ (39)
+Under the scale transformation $\omega\to\omega/b$, the kernel transforms as
+$\Pi^{R}\to b^{-\alpha}\Pi^{R},$ (40)
+implying a scale dependence of the effective coupling
+$g^{\prime}=b^{-\alpha}g.$ (41)
+The corresponding RG flow is therefore
+$\frac{dg}{dl}=(-\alpha)\,g+Cg^{2},$ (42)
+so that for $\alpha<0$ the interaction becomes relevant and grows algebraically under coarse graining.
+The qualitative difference from the Eliashberg regime originates from the organization of the relaxation spectrum.
+In the marginal case, each logarithmic shell contributes independently, leading to only logarithmic enhancement.
+By contrast, IR-singular spectra correspond to an accumulation of slow modes toward vanishing relaxation rates, which reorganizes the Cooper channel into a relevant infrared interaction.
+The corresponding RG flows are illustrated schematically in Fig.
 2.
-In the Eliashberg case, each logarithmic shell contributes independently, leading to marginal enhancement of pairing correlations.
-In contrast, a flat TDOS corresponds to a dense continuum of near–marginal relaxation modes.
-Fluctuations across widely separated time scales remain dynamically correlated, allowing pairing correlations to accumulate coherently.
-Superconductivity therefore emerges as a dynamical instability of the collective relaxation spectrum itself, rather than from the exchange of a specific bosonic mediator.
+The schematic curves correspond to the solutions of the RG equations $dg/dl=Cg^{2}$ (marginal case) and $dg/dl=(-\alpha)g$ (IR-singular case), yielding logarithmic and exponential growth of the effective pairing interaction, respectively.
+In the marginal regime associated with a flat TDOS ($\alpha=0$), the pairing interaction increases only logarithmically under coarse graining, reflecting the weak accumulation of infrared spectral weight.
+By contrast, for infrared-singular TDOS ($\alpha<0$), the coupling acquires a finite scaling dimension and grows exponentially with the RG scale, indicating that the Cooper channel becomes a relevant infrared instability.
+The figure thus provides a direct visualization of how the infrared structure of the relaxation spectrum controls the qualitative nature of pairing enhancement.
+Superconductivity thus emerges as an instability of the retarded collective response controlled by the TDOS, with logarithmic enhancement for flat spectra and algebraic growth only in the presence of an IR-singular slow-mode density.
 Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction.
 Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining.
-Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue).
-In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=(1-\eta)g$, leading to algebraic amplification of the pairing interaction (orange).
-The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
+A flat relaxation-rate density of states (marginal TDOS, $\alpha=0$) produces logarithmic growth, $dg/dl\sim g^{2}$, corresponding to a marginal infrared enhancement of pairing (blue).
+This behavior is analogous to the logarithmic scaling familiar from conventional BCS/Eliashberg-type pairing theories.
+In contrast, IR-singular relaxation spectra with enhanced infrared weight ($\alpha<0$) generate a relevant contribution, $dg/dl=(-\alpha)g$, leading to algebraic growth of the pairing interaction (orange).
+.
 IV IV. Pairing from Memory–Dominated Critical Dynamics
-The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes.
-Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
-Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum.
-The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes.
-Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
-In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
+The previous section established the universal infrared structure of the Cooper channel generated by the relaxation spectrum.
+In particular, the slow–mode reservoir reorganizes the retarded Cooper kernel into a singular form, leading to marginal enhancement of pairing when the TDOS remains finite in the infrared, and to a relevant instability only when the TDOS develops an infrared divergence.
+The remaining question is how this infrared amplification is tied to microscopic electronic pairing tendencies in correlated materials and how it determines the observable superconducting transition scale.
+In this section we connect the intrinsic electronic pairing seed to the memory–dominated Cooper kernel and show how their interplay produces an enhanced superconducting instability, which becomes algebraic only when the TDOS develops an infrared divergence.
+In the present framework, superconductivity does not originate from the formation of tightly bound Cooper pairs at the microscopic level.
+Instead, a weak intrinsic pairing tendency is already present in the electronic system, giving rise to fluctuating pair fields.
+The key mechanism is the accumulation of slow collective modes in the relaxation spectrum, which enhances the lifetime and overlap of these fluctuations.
+Superconductivity thus emerges as a collective dynamical instability driven by the reorganization of relaxation modes, rather than by pair formation itself.
+Equivalently, the superconducting transition should be understood not as a bound–state formation problem, but as a dynamical instability of the retarded Cooper kernel controlled by the infrared structure of the relaxation spectrum.
 IV.1 A. Intrinsic electronic pairing and dynamical formulation
-In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes [1, 30, 31].
-Starting from the Hubbard Hamiltonian,
-$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (49)
-projection to the strong–coupling sector yields the superexchange interaction
-$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (50)
-which favors local singlet formation. This interaction provides a natural microscopic origin of electronic pairing in systems near Mott localization.
-More generally, interacting electrons are described by a generic four–fermion interaction
-$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma}.$ (51)
-This interaction contains multiple collective channels.
-In conventional diagrammatic approaches these channels are analyzed through resummations such as particle–hole bubble diagrams or particle–particle ladder diagrams.
-Rather than focusing on a specific mediator, the present framework emphasizes the dynamical organization of the collective spectrum itself.
-While the four–fermion interaction generically allows multiple ordering channels, their ordering tendencies are not equivalent.
-Particle–hole instabilities such as charge or spin density waves typically require specific momentum structure, most notably Fermi surface nesting or commensurate wave vectors determined by the lattice.
-Such kinematic constraints restrict the available phase space for density–wave ordering.
-Superconductivity, by contrast, corresponds to condensation in the particle–particle channel at zero total momentum,
+In strongly correlated materials, short-range electronic interactions naturally generate local singlet correlations, providing a microscopic seed for pairing.
+For example, in systems proximate to Mott insulating phases, superexchange processes arising from strong Coulomb repulsion favor spin-singlet formation [1, 32, 33].
+More generally, however, the specific microscopic origin of pairing is not essential for the present analysis.
+At low energies, interacting electrons may be described by a generic four–fermion interaction,
+$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma},$ (43)
+which contains multiple collective channels.
+Projecting this interaction onto the particle–particle sector defines an effective pairing vertex in the Cooper channel, which serves as the intrinsic pairing interaction $g_{\rm pair}$.
+In conventional approaches these channels are analyzed through diagrammatic resummations, often emphasizing a specific bosonic mediator.
+In contrast, the present framework focuses on the dynamical organization of the collective spectrum itself.
+While several ordering channels are in principle allowed, their infrared efficiency differs.
+Particle–hole instabilities such as density waves require specific momentum structure, restricting their available phase space.
+By contrast, superconductivity corresponds to condensation in the particle–particle channel at zero total momentum,
 $\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (44)
-and therefore does not rely on special nesting conditions of the Fermi surface.
-Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
+and can therefore efficiently accumulate infrared spectral weight.
 Within the dynamical framework developed in Sec.
-III, the pairing susceptibility assumes the universal tt–matrix form in Eq.
-(41), where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations.
-Diagrammatically this corresponds to the Cooper-channel polarization obtained after integrating out microscopic electronic degrees of freedom.
-In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory.
-Near correlated criticality, however, the situation changes qualitatively.
-Because the relaxation spectrum becomes dense near $\lambda\to 0$, an extensive reservoir of slow collective modes emerges.
-This slow–mode reservoir reorganizes the infrared dynamics of the electronic system and strongly enhances the Cooper channel.
-Within the present framework critical dynamics does not create attraction.
-Instead it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
-The intrinsic pairing scale $E_{\rm pair}$ therefore acts as the microscopic seed, while the slow collective spectrum dynamically amplifies the Cooper instability.
-IV.2 B. Memory–dominated Cooper kernel from flat TDOS
-Because the collective pair fluctuations inherit the relaxational spectral structure generated by the slow-mode reservoir, the Cooper-channel polarization kernel reduces to a continuum integral over relaxation rates,
-$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (53)
-When the slow–mode reservoir develops,
-$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (54)
-the infrared behavior becomes
-$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (55)
-This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales.
-Physically, the flat TDOS implies that an extensive set of slow collective modes contributes coherently to the Cooper channel.
-Unlike conventional Eliashberg kernels, which remain marginal in the infrared, the present Cooper kernel becomes strongly enhanced by the infrared spectral weight of the slow–mode reservoir.
-The pairing interaction is therefore dynamically amplified by the collective memory of the system.
-IV.3 C. Algebraic instability and transition scale
-At finite temperature the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$.
-As a consequence the system cannot probe arbitrarily long time scales.
-In the imaginary-time formalism the dynamical fields are expanded in bosonic Matsubara frequencies
-$\Omega_{n}=2\pi nT.$ (56)
-The smallest nonzero dynamical frequency therefore sets a natural infrared scale of order
+III, the pairing susceptibility assumes a universal tt–matrix structure, as given in Eq.
+(32).
+In this expression, $g_{\rm pair}$ represents the intrinsic electronic pairing seed, while $\Pi^{R}(\omega)$ denotes the retarded Cooper kernel generated by the relaxation spectrum after integrating out microscopic degrees of freedom.
+A key point is that the present framework does not rely on a specific microscopic pairing mechanism.
+Instead, short-range electronic correlations provide a generic pairing seed, while the infrared dynamical spectrum controls how this tendency is amplified.
+In weakly interacting systems the pairing instability remains marginal, leading to the exponentially small transition scale of conventional BCS theory.
+Near correlated criticality, however, the relaxation spectrum becomes dense near $\lambda\to 0$, generating an extensive reservoir of slow collective modes.
+This slow–mode reservoir reorganizes the infrared dynamics and strongly enhances the Cooper channel.
+Thus, critical dynamics does not create attraction, but amplifies a preexisting pairing tendency through infrared spectral reorganization.
+IV.2 B. Fluctuation enhancement versus pairing instability
+The analysis above reveals a crucial distinction between two closely related but physically different quantities: the fluctuation correlator of the pair field and the retarded response kernel that governs superconducting instability.
+Within the Keldysh formulation, the fluctuation spectrum of pairing is encoded in the correlator
+$\Pi^{K}(\omega)\equiv\langle\Delta(\omega)\Delta^{\dagger}(\omega)\rangle_{K},$ (45)
+which, for a continuum of relaxational modes, takes the form
+$\Pi^{K}(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm pair}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (46)
+For a flat TDOS, $\rho_{\rm pair}(\lambda\to 0)=\rho_{0}$, this yields
+$\Pi^{K}(\omega)\sim\frac{\rho_{0}}{|\omega|},$ (47)
+indicating a strong infrared enhancement of pairing fluctuations. In the time domain, this corresponds to long-time correlations $K(t)\sim 1/t$ and scale-free noise.
+However, the superconducting transition is not determined by the fluctuation correlator, but by the retarded response function through the Thouless criterion in Eq.
+(33).
+The retarded kernel exhibits a different infrared structure.
+For a flat TDOS, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\frac{\Lambda}{|\omega|}$, leading to a logarithmic (marginal) enhancement and a BCS-like transition scale.
+Thus, although pairing fluctuations are strongly enhanced, this enhancement alone does not produce an algebraic increase of the transition temperature.
+The transition is controlled by the infrared structure of the retarded kernel, not by the fluctuation amplitude itself.
+A qualitatively different regime arises only when the TDOS develops an infrared divergence, $\rho_{\rm pair}(\lambda)\sim\lambda^{\alpha}(\alpha<0)$, for which $\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}$.
+In this case the retarded kernel itself becomes singular, and the pairing interaction acquires a relevant scaling dimension, leading to algebraic growth of the superconducting instability.
+Physical implication.
+Strong infrared fluctuations, long-time memory, and $1/f$-type noise are therefore natural consequences of a flat TDOS, but they do not by themselves imply high transition temperatures.
+Enhanced superconductivity requires not only strong fluctuations, but an infrared-divergent accumulation of slow modes that modifies the retarded response itself.
+This distinction clarifies that superconductivity in the present framework is controlled by the structure of the retarded kernel, while fluctuation correlators characterize the accompanying non-equilibrium and dynamical signatures.
+IV.3 C. Transition scale from infrared spectral structure
+The superconducting transition is determined by the Thouless criterion in Eq.
+(33), which shows that the transition temperature is controlled by the infrared behavior of the retarded Cooper-channel kernel.
+At finite temperature, the imaginary-time direction becomes compact, $\tau\in[0,\beta]$ with $\beta=1/T$, and the system cannot probe arbitrarily long time scales.
+Equivalently, the lowest Matsubara frequency sets the infrared scale,
 $|\omega|\sim T.$ (48)
-Substituting this scale into the infrared expression (55) gives
-$\Pi(T)\sim\frac{\rho_{0}}{T}.$ (58)
-Thus the infrared divergence present at zero temperature is universally regulated by the thermal scale.
-The superconducting transition follows from the Thouless criterion
-$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (59)
-which yields the algebraic transition scale
-$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (60)
-up to numerical factors of order unity.
-The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir.
-High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
-It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence.
-The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
+Thus, thermal fluctuations provide an infrared cutoff, and the transition temperature is determined by evaluating $\Pi^{R}(\omega)$ at an effective scale.
+Flat TDOS ($\alpha=0$): marginal regime. For a finite infrared TDOS,
+$\rho(\lambda\to 0)=\rho_{0},$ (49)
+the retarded kernel exhibits logarithmic scaling,
+$\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (50)
+Substituting $\omega\sim T_{c}$ into Eq. (33) yields
+$1-g_{\rm pair}\,\rho_{0}\ln\frac{\Lambda}{T_{c}}=0,$ (51)
+from which the transition temperature follows as
+$T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (52)
+Thus, a flat TDOS leads to a marginal enhancement of pairing, reproducing a BCS-like exponential transition scale despite the presence of strong infrared fluctuations.
+Infrared-divergent TDOS ($\alpha<0$): IR-singular regime. When the relaxation spectrum develops an infrared divergence,
+$\rho(\lambda)\sim\lambda^{\alpha},\qquad\alpha<0,$ (53)
+the real part of the retarded kernel acquires a power-law infrared singularity,
+$\mathrm{Re}\,\Pi^{R}(\omega)\sim C_{\alpha}|\omega|^{\alpha}.$ (54)
+At finite temperature, thermal fluctuations cut off the infrared divergence at the scale $\lambda_{\rm IR}\sim T$, so that the static retarded kernel should be evaluated with the effective lower limit set by $T$:
+$\Pi^{R}(0;T)\sim\int_{T}^{\Lambda}d\lambda\;\frac{\rho(\lambda)}{\lambda}.$ (55)
+For an infrared-singular TDOS, this gives
+$\Pi^{R}(0;T)\sim\int_{T}^{\Lambda}d\lambda\;\lambda^{\alpha-1}=\frac{T^{\alpha}-\Lambda^{\alpha}}{|\alpha|}\sim\frac{1}{|\alpha|}T^{-|\alpha|}.$ (56)
+The Thouless criterion therefore becomes
+$1-g_{\rm pair}C_{\alpha}T_{c}^{-|\alpha|}=0,$ (57)
+yielding
+$T_{c}\sim(g_{\rm pair}C_{\alpha})^{1/|\alpha|}\sim g_{\rm pair}^{1/|\alpha|}.$ (58)
+The enhancement becomes stronger for larger $|\alpha|$, since the retarded kernel diverges more strongly in the infrared as $T^{-|\alpha|}$.
+In this regime, the accumulation of slow modes toward vanishing relaxation rates reorganizes the Cooper channel into a relevant infrared interaction.
+As a result, the pairing instability is governed by a power-law infrared enhancement of the retarded kernel, and the exponential suppression characteristic of conventional pairing is replaced by an algebraic dependence on the interaction strength.
+Physical interpretation.
+The transition temperature derived above should be understood as the pairing instability scale determined by the retarded Cooper kernel.
+It is therefore controlled by the infrared spectral structure of the relaxation modes.
+A finite TDOS produces strong fluctuations but only logarithmic enhancement of the retarded kernel, leading to an exponentially moderate pairing scale.
+By contrast, an infrared-divergent TDOS modifies the retarded response itself, generating algebraic growth of the pairing interaction and a parametrically enhanced instability scale.
+In real materials, however, the observable superconducting transition requires not only the formation of pairs but also global phase coherence.
+The actual transition temperature is therefore bounded by the smaller of the pairing scale and the superfluid stiffness,
 $k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right),$ (59)
-with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
-The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$.
-The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
-In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios.
-The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
+with $\rho_{s}\sim n_{s}/m^{*}$. The scaling of phase stiffness within the present framework is discussed in Sec. V.
+The central result is that the pairing instability itself is governed by a single infrared characteristic of the system, namely the low-energy structure of the TDOS.
+Superconductivity thus emerges as a consequence of dynamical spectral organization, providing an alternative to descriptions based solely on the exchange of specific bosonic mediators.
 Figure 3: Experimental consequences of memory–dominated critical dynamics.
-(a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality.
-The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest.
-Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning.
-(b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$.
-Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
+(a) Schematic superconducting dome emerging from the infrared spectral structure of the slow–mode reservoir.
+The transition temperature follows $T_{c}(p)\sim\Lambda(p)\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}(p)}\right)$ and is maximal near $p_{\star}$, where both the infrared spectral weight and the extent of the slow–mode reservoir are largest.
+Suppression of $T_{c}$ away from this regime arises from the combined effect of reduced infrared spectral weight and finite infrared cutoff, without invoking competing orders.
+(b) Uemura scaling showing the proportionality between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$, in regimes where phase coherence limits the transition.
+Both pairing and phase stiffness are influenced by the same infrared spectral weight $\rho_{0}$, providing a common dynamical origin for the observed scaling across material families.
 V V. Experimental consequences of memory–dominated critical dynamics
@@
 V.1 A. Infrared spectral control and automatic emergence of the dome
-The linearized transition scale derived in Eq.
-(60) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir.
-We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
-In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
-$\rho(\lambda\to 0)=\rho_{*}>0,$ (62)
-without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
-External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir.
-Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal.
-The collective mass term behaves as
-$r(p)\sim|p-p_{\star}|,$ (63)
-so that the associated cutoff obeys
+[…]
+We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical state characterized by an extended distribution of slow relaxation modes.
+In this state, feedback between electronic correlations and collective dynamics maintains a broad near-marginal sector of the TDOS,
+$\rho(\lambda\to 0)=\rho_{*}(p),$ (60)
+whose magnitude and infrared extent depend on external control parameters such as doping or pressure.
+Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared extent of the slow–mode reservoir is maximal.
+Deviations from $p_{\star}$ introduce a finite mass scale $r(p)$ that truncates the infrared spectrum,
 $\Lambda(p)\sim r^{z\nu}\sim|p-p_{\star}|^{z\nu}.$ (61)
-Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
+At the same time, the low-energy spectral weight evolves smoothly,
 $\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (62)
-with $f$ a decreasing crossover function.
-Substituting into Eq. (60) gives
-$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (66)
+where $f$ is a decreasing crossover function.
+Substituting these dependencies into Eq. (52) yields
+$T_{c}(p)\sim\Lambda(p)\,\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}(p)}\right),$ (63)
 which naturally produces a superconducting dome.
-Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
-In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point.
-Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
+The physical origin of the dome can be understood as a competition between two infrared effects.
+Near $p_{\star}$, the slow–mode reservoir is maximally extended, leading to a large $\rho_{0}$ and a large cutoff $\Lambda$, and hence enhanced pairing.
+Away from this regime, superconductivity is suppressed either by the reduction of the infrared spectral weight (overdoped side) or by the progressive truncation of the slow–mode reservoir (underdoped side).
+In this interpretation, optimal doping does not correspond to a fine-tuned symmetry-breaking quantum critical point.
+Rather, it marks the regime in which the infrared spectral organization of relaxation modes is most fully developed.
+The superconducting dome therefore emerges as a robust consequence of infrared spectral reorganization, without requiring a specific bosonic mediator or fine tuning of microscopic interactions.
 V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
 While Sec.
-IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence.
-In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
+IV establishes that the pairing instability is governed by the infrared structure of the retarded Cooper kernel, the observable transition temperature is ultimately constrained by global phase coherence.
+In systems with reduced carrier density or strong correlations, the superconducting transition may therefore be limited by phase stiffness rather than by pair formation.
 The long-wavelength phase dynamics of the Cooper field is described by the effective action
@@
 $T_{\theta}\sim\rho_{s}.$ (65)
-Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport.
-The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
-In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$.
-Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight.
-Integrating out amplitude fluctuations in the presence of the slow-mode reservoir generically yields
-$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (69)
-up to nonuniversal numerical factors.
-Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
-Combining Eqs. (68) and (69) gives
-$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (70)
-establishing a direct proportionality between the superconducting transition temperature and superfluid density.
-This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density.
-In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
-Within the present framework, this proportionality is not imposed phenomenologically.
-It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness.
-The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
+Within the present framework, both pairing and phase coherence are influenced by the same infrared spectral structure of the relaxation modes.
+The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore controlled by the density of carriers that remain phase coherent over long time scales.
+In a memory–dominated regime, such long-time coherence is governed by the low-energy spectral weight $\rho_{0}$ of the slow–mode reservoir.
+As a result, the superfluid stiffness is expected to scale with the infrared spectral weight,
+$\rho_{s}\sim\rho_{0},$ (66)
+up to material-dependent prefactors that depend on microscopic details such as band structure and effective mass.
+Combining Eqs. (65) and (66), one expects that the superconducting transition temperature tracks the superfluid density,
+$T_{c}\propto\rho_{s},$ (67)
+consistent with the empirical Uemura scaling observed in cuprates and other unconventional superconductors.
+In contrast to weak-coupling BCS superconductors, where phase stiffness is large and essentially decoupled from gap formation, strongly correlated systems operate in a regime where both pairing and coherence are governed by infrared collective dynamics.
+Importantly, however, the transition temperature is determined by the smaller of the pairing scale and the phase stiffness, $k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right)$, so that Uemura scaling emerges most clearly in regimes where phase stiffness provides the dominant limitation.
+Within this picture, the observed proportionality between $T_{c}$ and superfluid density does not arise from a direct identity between pairing strength and phase rigidity, but from their common dependence on the infrared spectral weight of the slow–mode reservoir.
+The TDOS therefore acts as a unifying control parameter that links pairing susceptibility and phase coherence without requiring a specific microscopic pairing mechanism.
 V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
@@
 Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
-V.4 D. Unified infrared scaling picture
+Related derivations are presented in Appendix E.
+Unified infrared scaling picture.
 We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
-The density of relaxation rates controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response.
+The density of relaxation rates plays a central organizing role: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response.
 Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
 This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling.
-The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
+The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
 Figure 4: Dynamical classification determined by the TDOS exponent $\alpha$.
-The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi(\omega)\sim|\omega|^{\alpha-1}$ and separates the system into memory–dominated ($\alpha<1$), marginal Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences.
-Here $\omega$ denotes Matsubara frequency.
+The scaling $\rho(\lambda)\sim\lambda^{\alpha}$ implies $\Pi^{R}(\omega)\sim|\omega|^{\alpha}$ (with logarithmic behavior at $\alpha=0$), and separates the system into infrared-singular ($\alpha<0$), marginal ($\alpha=0$), Ohmic ($\alpha=1$), and analytic ($\alpha>1$) regimes, with corresponding pairing consequences.
+Here $\omega$ denotes frequency (Matsubara or real, depending on context).
 VI VI. Discussion
-The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by the softening of a single order parameter.
+The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum, rather than being described solely by the softening of a single order parameter.
 The time–scale density of states provides a quantitative measure of this organization.
@@
 Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields.
-Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero.
-In this regime internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales.
-The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
+Near correlated critical regimes, however, relaxation rates form a dense continuum extending toward zero.
+In this regime, internal modes cannot be trivially integrated out: the decay spectrum remains dynamically relevant across a wide range of time scales.
+The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a generic consequence of dynamical criticality.
 This observation highlights a structural difference from Hertz–Millis–type approaches.
@@
 In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels.
-Criticality is reinterpreted not as single-mode softening but as a reorganization of the entire decay spectrum.
-The TDOS therefore replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
+Criticality is reinterpreted not as single-mode softening but as a reorganization of the decay spectrum.
+The TDOS therefore provides a microscopic spectral quantity governing infrared universality, complementing phenomenological damping descriptions.
 This structure is summarized schematically in Fig.
@@
 The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states.
-For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form $\Pi(\omega)\sim|\omega|^{\alpha-1}$.
-Different dynamical regimes therefore emerge as special cases of the same spectral organization.
-For $\alpha>1$ the density of slow modes vanishes in the infrared, so that the polarization kernel remains finite and the collective dynamics approaches the weakly damped limit, with the infrared response dominated by the nondissipative part $\chi^{-1}\simeq r+cq^{2}$.
-The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$.
+For a power-law TDOS, $\rho(\lambda)\sim\lambda^{\alpha}$, the retarded polarization kernel exhibits the infrared scaling $\mathrm{Re}\,\Pi^{R}(\omega)\sim|\omega|^{\alpha}$ for $\alpha<1$, with logarithmic behavior in the marginal case $\alpha=0$.
+The inverse susceptibility is correspondingly given by $\chi^{-1}(q,\omega)\simeq r+cq^{2}+\Pi^{R}(\omega)$, so that different dynamical regimes emerge as special cases of a unified spectral organization.
+For $\alpha>1$, the density of slow modes vanishes in the infrared, and the dynamical correction becomes subleading compared to the static terms at low frequencies.
+In this regime the collective dynamics approaches a weakly damped, regular (analytic) limit, with the infrared response dominated by the nondissipative contribution.
+The case $\alpha=1$ corresponds to the Ohmic boundary.
+In this case the kernel acquires a linear frequency dependence up to logarithmic corrections, $\Pi^{R}(\omega)\sim|\omega|$, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$.
 Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges.
-In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
-Interestingly, the infrared response generated by a flat TDOS is structurally identical to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 32].
-In particular, the resulting spectrum corresponds in frequency space to a logarithmic real part and a nearly constant absorptive component of the susceptibility.
+In this regime the polarization exhibits logarithmic infrared behavior, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln(\Lambda/|\omega|)$, while the absorptive part remains finite at low frequencies.
+This leads to long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
+Interestingly, the infrared response generated by a flat TDOS is closely related to anomalous dynamical susceptibilities that have appeared in several apparently unrelated contexts [5, 34].
+In particular, the resulting spectrum corresponds to a logarithmic real part and a weakly frequency-dependent absorptive component of the susceptibility.
 Similar infrared structures were introduced phenomenologically in the marginal Fermi liquid theory of strange metals and also arise in solvable strongly correlated spin models such as the Sachdev–Ye spin liquid.
@@
 The lower panel of Fig.
-4 illustrates the direct consequence of this spectral organization for superconductivity.
-Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic Cooper kernel $\Pi(T)\sim\ln(\Omega/T)$ and hence to the exponentially small transition scale characteristic of conventional marginal pairing theories.
-In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$.
-The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
+4 illustrates the consequence of this spectral organization for superconductivity.
+Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, the transition scale is controlled by the infrared structure of the pairing kernel.
+In conventional Ohmic or marginal pairing scenarios, the Cooper kernel acquires a logarithmic enhancement $\Pi(T)\sim\ln(\Omega/T)$, leading to the exponentially small transition scale characteristic of BCS-like theories.
+In the memory-dominated regime with a flat TDOS, the retarded pairing kernel similarly exhibits logarithmic infrared enhancement, $\Pi(T)\sim\rho_{0}\ln(\Lambda/T)$.
+While the transition scale retains a BCS-like exponential form, the effective parameters are now controlled by the infrared spectral weight of slow collective modes.
+Only when the TDOS becomes infrared-singular ($\alpha<0$) does the pairing kernel acquire a power-law enhancement, $\Pi(T)\sim T^{\alpha}$, leading to algebraic scaling of the transition temperature.
 A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities.
@@
 Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir.
-The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
+In the marginal regime, the exponential suppression characteristic of weak-coupling BCS theory is not removed but can be substantially reduced due to enhanced infrared spectral weight.
+Only when the TDOS becomes infrared-singular does the pairing scale exhibit algebraic enhancement.
 In this sense, superconducting order is governed by condensation in a Cooper sector embedded within an extended slow-mode manifold rather than by binding through a sharp bosonic excitation.
-A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena.
-The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel.
+A single infrared parameter, $\rho_{0}$, plays a central role in both thermodynamic and dynamical phenomena.
+The superconducting dome can be interpreted as reflecting truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel.
 Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
@@
 The relaxation-rate organization emphasized here is not merely a phenomenological assumption.
-Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues.
-Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra.
-A dense TDOS therefore reflects a generic mathematical feature of dynamical systems with extended temporal correlations rather than a model-specific construction.
-Related derivations are presented in Appendix F.
+Both Liouvillian evolution in open quantum systems and generalized memory-kernel formulations admit representations in terms of decay modes characterized by relaxation eigenvalues.
+Whenever long-time correlations are present, Laplace-domain representations naturally lead to continuous relaxation spectra.
+A dense TDOS can therefore be understood as a generic feature of dynamical systems with extended temporal correlations rather than a model-specific construction.
 This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics.
-While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem continuously exchanging coherence and information with the remaining degrees of freedom.
-From the perspective of such a subsystem, the relevant generator of dynamics is not the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes.
-The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues rather than solely by the spectrum of excitation energies.
-In this sense the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
-From this perspective Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality.
-Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
+While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem exchanging coherence and information with the remaining degrees of freedom.
+From this perspective, the relevant generator of dynamics is not the Hamiltonian itself but an effective Liouvillian governing relaxation and memory processes.
+The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues in the effective description.
+In this sense, the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
 Temporal scaling and dynamical criticality.
 An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality.
-In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges indirectly through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$.
+In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$.
 The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
@@
 When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales.
-The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions, with logarithmic or stronger singular structures depending on the observable sector.
-Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality.
-This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena [33, 34].
-Memory–dominated dynamics as a guiding principle for room-temperature superconductivity.
-Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition.
-Within the present framework, however, enhancing the bare pairing scale alone is insufficient.
-A decisive control parameter for high transition temperatures may be the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
-A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$.
-From this viewpoint certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat-TDOS regime.
-The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [35].
-Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself.
-In this perspective high-temperature superconductivity is most naturally favored in regimes characterized by maximal infrared memory rather than simply by extreme coupling strength.
+The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
+Scale-free temporal dynamics can thus be viewed as emerging directly from the structure of the relaxation spectrum, rather than being solely inherited from spatial criticality.
+This perspective is closely related to scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural framework for describing dynamical critical phenomena [35, 36].
+Memory–dominated dynamics and prospects for high-temperature superconductivity.
+Within the present framework, superconducting transition temperatures are controlled not only by the microscopic pairing interaction but also by the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
+In the marginal regime corresponding to a flat TDOS ($\alpha=0$), the retarded pairing kernel exhibits logarithmic infrared enhancement, leading to a BCS-like transition scale, $T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right)$, where $\Lambda$ is an electronic cutoff scale and $\rho_{0}$ is the infrared spectral weight of slow collective modes.
+Because $\Lambda$ is set by electronic energy scales, this form naturally allows transition temperatures in the range of $10^{2}$ K for moderate coupling, providing a consistent framework for understanding high-$T_{c}$ superconductivity without invoking finely tuned bosonic mediators.
+Further enhancement of the transition scale may occur if the relaxation spectrum develops an infrared-singular structure ($\alpha<0$), corresponding to an accumulation of slow modes toward vanishing relaxation rates.
+In this regime, the pairing kernel acquires a power-law infrared enhancement, and the transition temperature follows an algebraic scaling with the effective interaction strength.
+These considerations suggest a complementary perspective on superconductivity design: beyond increasing microscopic pairing interactions, engineering the infrared organization of collective dynamics may provide an effective route toward higher transition temperatures [37].
+In this sense, dynamical phases supporting enhanced slow-mode spectral weight represent promising candidates for realizing superconductivity at elevated temperatures.
 VII VII. Conclusion
@@
 A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response.
-Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
+Within this regime, intrinsic electronic pairing tendencies are dynamically enhanced, leading to logarithmic infrared amplification of the retarded pairing kernel.
+While the resulting transition scale retains a BCS-like exponential form, it is controlled by the infrared spectral weight of slow collective modes.
+Algebraic enhancement of the transition temperature arises only when the TDOS becomes infrared-singular ($\alpha<0$), corresponding to an accumulation of slow modes toward vanishing relaxation rates.
 This mechanism provides a unified account of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state.
@@
 Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
-In this perspective, high–temperature superconductivity can be viewed as a generic consequence of favorable infrared spectral organization in strongly correlated matter.
+In this perspective, high–temperature superconductivity can be understood as arising from favorable infrared spectral organization in strongly correlated matter.
 Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
@@
 The MSRJD action for a stochastic dynamical variable $\phi(t)$ with memory kernel and Gaussian noise may be written as
-$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S67)
+$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-L^{R}(t-t^{\prime})\right]$ (S67)
 $\quad\phi(t^{\prime})-\int dt\,dt^{\prime}\,\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$
-where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
+where $\tilde{\phi}$ denotes the response field, $L^{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
 The real–time Keldysh functional integral for an open quantum system is naturally formulated in terms of classical and quantum fields, $\phi_{\mathrm{cl}}$ and $\phi_{q}$, defined by the Keldysh rotation
@@
 To quadratic order in the quantum field, the Keldysh action assumes the general structure
-$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S69)
-where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component.
-The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
+$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&L^{A}(\omega)\\ L^{R}(\omega)&iL^{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S69)
+where $L^{R,A}$ are retarded/advanced propagators and $L^{K}$ is the Keldysh (fluctuation) component.
+The retarded component encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
 In the regime of strong decoherence or large occupation, where quantum interference effects are suppressed, the quantum field $\phi_{q}$ acts as a response field. Identifying
 $\phi_{q}\;\leftrightarrow\;i\tilde{\phi}$, (S70)
-and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$.
+and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $L^{K}$.
 The corresponding MSRJD action then takes the form (S67), with the correspondence
-$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S71)
+$L^{R}(\omega)\;\leftrightarrow\;L^{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}L^{K}(\omega).$ (S71)
 Thus the MSRJD formulation used in the present work can be viewed as the classical (or strongly dissipative) limit of the Keldysh functional integral, in which the relevant dynamical information is encoded in the retarded kernel and its associated spectral properties.
@@
 In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
-Appendix D Appendix D: Derivation of the Cooper-channel polarization kernel from a relaxational mode continuum
-In this Appendix we present an explicit derivation of the polarization kernel used in the Cooper channel within the relaxation-spectrum framework.
-The Cooper-channel object relevant for pairing enhancement is the fluctuation (or correlator) spectrum of the collective pair field.
-The construction below should be understood, in the same spirit as Appendix B, as an effective infrared description after microscopic electronic degrees of freedom have already been integrated out.
-At that stage the relevant dynamical variables are collective pair modes, and the polarization kernel is obtained from their stochastic two-point correlations rather than from an explicit microscopic fermionic loop.
-(I) Single relaxational pair mode: Consider a single complex collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$ obeying the linear Langevin equation
+Appendix D Appendix D: Retarded Keldysh derivation of the superconducting transition scale from the collective Cooper response
+In this Appendix we reformulate the Cooper-channel instability directly in the real-time Keldysh language.
+The central point is that the superconducting transition temperature must be determined from the retarded linear-response kernel of the collective pair field, rather than from the Keldysh fluctuation correlator itself.
+The fluctuation propagator will be analyzed subsequently, but the transition criterion is fixed first by the retarded ladder pole.
+We therefore proceed in the following order.
+First, we introduce the Keldysh quadratic action for the collective pair field.
+Second, we derive the retarded propagator of a single relaxational pair mode.
+Third, we integrate over a continuum of relaxational modes with a TDOS characterized by a general infrared exponent $\alpha$.
+Fourth, we resum the retarded Cooper ladder and derive the Thouless criterion for $T_{c}$.
+Finally, we analyze explicitly the cases $\alpha=0$ and $\alpha=-1/2$.
+(I) Collective pair field and quadratic Keldysh action: We begin from the collective Cooper pair field
+$\Delta(x,t)=c_{\uparrow}(x,t)c_{\downarrow}(x,t),$ (S101)
+and its Keldysh classical/quantum components
+$\Delta_{cl}(x,t)=\frac{1}{2}\left(\Delta_{+}(x,t)+\Delta_{-}(x,t)\right),$ (S102)
+$\Delta_{q}(x,t)=\Delta_{+}(x,t)-\Delta_{-}(x,t),$
+where ++ and - denote the forward and backward Keldysh contours.
+At quadratic level, the real-time action of the collective pair sector has the generic form
+$S_{2}=$ (S103)
+$\int_{\omega}\begin{pmatrix}\Delta_{cl}^{*}(\omega)&\Delta_{q}^{*}(\omega)\end{pmatrix}\begin{pmatrix}0&\left[L_{0}^{A}(\omega)\right]^{-1}\\ \left[L_{0}^{R}(\omega)\right]^{-1}&\Sigma^{K}(\omega)\end{pmatrix}\begin{pmatrix}\Delta_{cl}(\omega)\\ \Delta_{q}(\omega)\end{pmatrix}$,
+where $\int_{\omega}\equiv\int d\omega/(2\pi)$. The corresponding propagator matrix is
+$\hat{L}(\omega)=\begin{pmatrix}L^{K}(\omega)&L^{R}(\omega)\\ L^{A}(\omega)&0\end{pmatrix}$, (S104)
+with
+$L^{R}(\omega)=\frac{1}{\left[L_{0}^{R}(\omega)\right]^{-1}},\quad L^{A}(\omega)=\frac{1}{\left[L_{0}^{A}(\omega)\right]^{-1}}.$ (S105)
+The superconducting instability is determined by the pole of the retarded pair susceptibility.
+Therefore, throughout the first part of this Appendix, the relevant object is the retarded pair response kernel.
+(II) Retarded propagator of a single relaxational pair mode: Consider first a single collective pair mode $\Delta_{\lambda}(t)$ with relaxation rate $\lambda>0$. Its linear relaxational equation is
 $\partial_{t}\Delta_{\lambda}(t)=-\lambda\,\Delta_{\lambda}(t)+\xi_{\lambda}(t),$ (S106)
-where $\xi_{\lambda}(t)$ is Gaussian white noise with
-$\langle\xi_{\lambda}(t)\rangle=0,\qquad\langle\xi_{\lambda}(t)\xi_{\lambda^{\prime}}^{*}(t^{\prime})\rangle=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(t-t^{\prime}).$ (S102)
-Here $D_{\lambda}$ denotes the noise strength associated with the mode $\lambda$.
-In the stationary regime, obtained by sending the initial time to -$\infty$, the exact solution of Eq. (S101) is
-$\Delta_{\lambda}(t)=\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s).$ (S103)
-Differentiating Eq. (S103) directly yields
-$\partial_{t}\Delta_{\lambda}(t)$ $=\xi_{\lambda}(t)-\lambda\int_{-\infty}^{t}ds\,e^{-\lambda(t-s)}\xi_{\lambda}(s)$
-$=\xi_{\lambda}(t)-\lambda\Delta_{\lambda}(t),$ (S104)
-which confirms Eq. (S101).
-(II) Time-domain two-point correlation function: We now compute the stationary correlator
-$C_{\lambda}(\tau)\equiv\langle\Delta_{\lambda}(\tau)\Delta_{\lambda}^{*}(0)\rangle.$ (S105)
-For $\tau>0$, inserting Eq. (S103) gives
-$C_{\lambda}(\tau)$ $=\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\,\langle\xi_{\lambda}(s)\xi_{\lambda}^{*}(s^{\prime})\rangle$
-$=2D_{\lambda}\int_{-\infty}^{\tau}ds\int_{-\infty}^{0}ds^{\prime}\,e^{-\lambda(\tau-s)}e^{\lambda s^{\prime}}\delta(s-s^{\prime}).$ (S106)
-Because the delta function enforces $s=s^{\prime}$, while the second integral restricts $s^{\prime}\leq 0$, the effective integration domain reduces to $s\leq 0$. Therefore
-$C_{\lambda}(\tau)$ $=2D_{\lambda}\int_{-\infty}^{0}ds\,e^{-\lambda(\tau-s)}e^{\lambda s}$
-$=2D_{\lambda}e^{-\lambda\tau}\int_{-\infty}^{0}ds\,e^{2\lambda s}$
-$=2D_{\lambda}e^{-\lambda\tau}\left[\frac{e^{2\lambda s}}{2\lambda}\right]_{-\infty}^{0}$
-$=\frac{D_{\lambda}}{\lambda}e^{-\lambda\tau}.$ (S107)
-By stationarity and Hermitian symmetry, the result extends to all times as
-$C_{\lambda}(t)=\frac{D_{\lambda}}{\lambda}\,e^{-\lambda|t|}.$ (S108)
-Equation (S108) is the exact two-point correlation function of a single Ornstein–Uhlenbeck pair mode.
-The factor $1/\lambda$ in front of the exponential is crucial.
-It is precisely this factor that removes the extra numerator $\lambda$ appearing in the Fourier transform of the bare retarded kernel.
-(III) Frequency-domain spectral density: The mode-resolved fluctuation spectrum is defined as
-$S_{\lambda}(\omega)\equiv\int_{-\infty}^{\infty}dt\,e^{i\omega t}C_{\lambda}(t).$ (S109)
-Substituting Eq. (S108) gives
-$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\int_{-\infty}^{\infty}dt\,e^{i\omega t}e^{-\lambda|t|}.$ (S110)
-Splitting the integral into positive and negative times,
-$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}\right].$ (S111)
-Changing variables u=tu=-t in the second term yields
-$\int_{-\infty}^{0}dt\,e^{(\lambda+i\omega)t}=\int_{0}^{\infty}du\,e^{-(\lambda+i\omega)u},$ (S112)
-so that
-$S_{\lambda}(\omega)=\frac{D_{\lambda}}{\lambda}\left[\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}+\int_{0}^{\infty}dt\,e^{-(\lambda+i\omega)t}\right].$ (S113)
+where $\xi_{\lambda}(t)$ is a stochastic source. The deterministic linear-response part is
+$(\partial_{t}+\lambda)\Delta_{\lambda}(t)=\xi_{\lambda}(t).$ (S107)
+The retarded Green’s function $L_{\lambda}^{R}(t-t^{\prime})$ is defined as the inverse of the differential operator $\partial_{t}+\lambda$, namely
+[…]
+For $t\neq t^{\prime}$, the delta function vanishes and Eq. (S108) reduces to
+$(\partial_{t}+\lambda)L_{\lambda}^{R}(t-t^{\prime})=0.$ (S109)
+Hence
+$L_{\lambda}^{R}(t-t^{\prime})=A\,e^{-\lambda(t-t^{\prime})}\quad\text{for }t>t^{\prime}.$ (S110)
+Integrating Eq. (S108) across $t=t^{\prime}$ gives
+$\int_{t^{\prime}-\epsilon}^{t^{\prime}+\epsilon}dt\,(\partial_{t}+\lambda)L_{\lambda}^{R}(t-t^{\prime})=\int_{t^{\prime}-\epsilon}^{t^{\prime}+\epsilon}dt\,\delta(t-t^{\prime})=1.$ (S111)
+Since the integral of the regular term $\lambda L_{\lambda}^{R}$ vanishes as $\epsilon\to 0$, one obtains
+$L_{\lambda}^{R}(t^{\prime}+0)-L_{\lambda}^{R}(t^{\prime}-0)=1.$ (S112)
+With the retarded boundary condition $L_{\lambda}^{R}(t^{\prime}-0)=0$, this yields $A=1$, so that
+$L_{\lambda}^{R}(t-t^{\prime})=\theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S113)
+Fourier transforming,
+$L_{\lambda}^{R}(\omega)=\int_{-\infty}^{\infty}dt\,e^{i\omega t}\theta(t)e^{-\lambda t}=\int_{0}^{\infty}dt\,e^{-(\lambda-i\omega)t}=\frac{1}{\lambda-i\omega}.$ (S114)
+Similarly,
+$L_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega}.$ (S115)
+Thus a single relaxational pair mode contributes the retarded response kernel
+$L_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega}.$ (S116)
+(III) Continuum of relaxational pair modes and TDOS: We now assume that, after integrating out microscopic fermionic degrees of freedom, the low-energy Cooper sector is represented by a continuum of relaxational pair modes labeled by $\lambda$.
+Let $\rho_{\rm pair}(\lambda)$ denote the TDOS of pair relaxation rates.
+The irreducible retarded Cooper kernel is then obtained by summing the mode-resolved retarded response over the continuum:
+$\Pi^{R}(\omega;T)=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda-i\omega}.$ (S117)
+Here $\Lambda$ is the ultraviolet cutoff of the collective theory, and the lower limit is taken to be $T$, which serves as the infrared thermal cutoff in the static instability problem.
+Equation (S117) is the retarded analogue of the usual Cooper bubble after projection onto the collective pair sector.
+It is this $\Pi^{R}$, rather than the Keldysh fluctuation correlator, that enters the linear-response instability criterion.
+It is useful to separate real and imaginary parts:
+$\Pi^{R}(\omega;T)$ $=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda+i\omega}{\lambda^{2}+\omega^{2}}$
+$=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda}{\lambda^{2}+\omega^{2}}$
+$\quad+i\omega\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S118)
+Therefore
+$\Re\Pi^{R}(\omega;T)=\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S119)
+and
+$\Im\Pi^{R}(\omega;T)=\omega\int_{T}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S120)
+In the static limit $\omega\to 0$, one obtains
+$\Pi^{R}(0;T)=\int_{T}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda}.$ (S121)
+This is the quantity that controls the transition temperature.
+(IV) Retarded collective ladder and Thouless criterion: Let $g_{\rm pair}>0$ denote the short-distance irreducible pairing vertex in the projected Cooper sector.
+The retarded ladder equation for the full pair susceptibility is
+$\chi_{\Delta}^{R}(\omega;T)=\Pi^{R}(\omega;T)+\Pi^{R}(\omega;T)\,g_{\rm pair}\,\chi_{\Delta}^{R}(\omega;T).$ (S122)
+Solving Eq. (S122) for $\chi_{\Delta}^{R}$, we obtain
+$\chi_{\Delta}^{R}(\omega;T)=\frac{\Pi^{R}(\omega;T)}{1-g_{\rm pair}\Pi^{R}(\omega;T)}.$ (S123)
+Equivalently, the series form is
+$\chi_{\Delta}^{R}(\omega;T)=$ $\Pi^{R}(\omega;T)+\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)$
+$+\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)g_{\rm pair}\Pi^{R}(\omega;T)+\cdots.$ (S124)
+The superconducting instability occurs when the retarded susceptibility develops a pole at zero external frequency and momentum. Therefore the Thouless criterion is
+$1-g_{\rm pair}\Pi^{R}(0;T_{c})=0.$ (S125)
+Using Eq. (S121), the transition scale is determined by
+$1-g_{\rm pair}\int_{T_{c}}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda}=0.$ (S126)
+Equation (S126) is the central formula for the transition temperature in the retarded linear-response formulation.
+(V) General infrared TDOS exponent $\alpha$: We now assume the infrared TDOS
+$\rho_{\rm pair}(\lambda)=\rho_{\alpha}\,\lambda^{\alpha},\qquad T\leq\lambda\leq\Lambda,$ (S127)
+with $\rho_{\alpha}>0$. Then Eq. (S121) becomes
+$\Pi^{R}(0;T)=\rho_{\alpha}\int_{T}^{\Lambda}d\lambda\,\lambda^{\alpha-1}.$ (S128)
+For $\alpha\neq 0$,
+$\int d\lambda\,\lambda^{\alpha-1}=\frac{\lambda^{\alpha}}{\alpha},$ (S129)
+hence
+$\Pi^{R}(0;T)=\frac{\rho_{\alpha}}{\alpha}\left(\Lambda^{\alpha}-T^{\alpha}\right),\qquad\alpha\neq 0.$ (S130)
+For $\alpha=0$,
+$\Pi^{R}(0;T)=\rho_{0}\int_{T}^{\Lambda}\frac{d\lambda}{\lambda}=\rho_{0}\ln\!\left(\frac{\Lambda}{T}\right).$ (S131)
+Substituting into the Thouless criterion $1-g_{\rm pair}\Pi^{R}(0;T_{c})=0$, one finds
+$1-g_{\rm pair}\frac{\rho_{\alpha}}{\alpha}\left(\Lambda^{\alpha}-T_{c}^{\alpha}\right)=0,\qquad\alpha\neq 0,$ (S132)
+and
+$1-g_{\rm pair}\rho_{0}\ln\!\left(\frac{\Lambda}{T_{c}}\right)=0,\qquad\alpha=0.$ (S133)
+It is useful to separate three regimes.
+(a) $\alpha>0$:
+$\Pi^{R}(0;T\to 0)\to\frac{\rho_{\alpha}}{\alpha}\Lambda^{\alpha},$ (S134)
+which is finite. Therefore there is no automatic infrared divergence in the Cooper kernel. A superconducting instability exists only if
+$g_{\rm pair}\frac{\rho_{\alpha}}{\alpha}\Lambda^{\alpha}>1.$ (S135)
+(b) $\alpha=0$: the kernel diverges logarithmically and yields a BCS-like transition scale.
+(c) $\alpha<0$: the kernel diverges algebraically as $T\to 0$, producing an IR-singular enhancement of pairing.
+(VI) Case $\alpha=0$ (flat TDOS and BCS-like exponential scale): For a flat infrared TDOS,
+$\rho_{\rm pair}(\lambda)=\rho_{0},$ (S136)
+Eq. (S131) gives
+$\Pi^{R}(0;T)=\rho_{0}\ln\!\left(\frac{\Lambda}{T}\right).$ (S137)
+The Thouless criterion results in
+$T_{c}=\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (S138)
+Thus a flat TDOS leads to a BCS-like exponential transition scale.
+The physical interpretation, however, is different from weak-coupling BCS: the logarithm is controlled not by the quasiparticle density of states, but by the infrared density of slow collective pair relaxation modes.
+(VII) Case $\alpha=-1/2$ (IR-singular algebraic enhancement): We now consider the singular infrared TDOS
+$\rho_{\rm pair}(\lambda)=\rho_{-1/2}\lambda^{-1/2}.$ (S139)
+Then
+$\Pi^{R}(0;T)=\rho_{-1/2}\int_{T}^{\Lambda}d\lambda\,\lambda^{-3/2}.$ (S140)
 Using
-$\int_{0}^{\infty}dt\,e^{-at}=\frac{1}{a},\qquad\Re a>0,$ (S114)
-we obtain
-$S_{\lambda}(\omega)$ $=\frac{D_{\lambda}}{\lambda}\left(\frac{1}{\lambda-i\omega}+\frac{1}{\lambda+i\omega}\right)$
-$=\frac{D_{\lambda}}{\lambda}\frac{2\lambda}{\lambda^{2}+\omega^{2}}.$ (S115)
+$\Pi^{R}(0;T)$ $=\rho_{-1/2}\left[-2\lambda^{-1/2}\right]_{T}^{\Lambda}$
+$=2\rho_{-1/2}\left(T^{-1/2}-\Lambda^{-1/2}\right),$ (S141)
+the Thouless criterion becomes
+$1-2g_{\rm pair}\rho_{-1/2}\left(T_{c}^{-1/2}-\Lambda^{-1/2}\right)=0.$ (S142)
+The exact transition scale is
+$T_{c}=\left(\Lambda^{-1/2}+\frac{1}{2g_{\rm pair}\rho_{-1/2}}\right)^{-2}.$ (S143)
+In the weak-coupling limit
+$\frac{1}{2g_{\rm pair}\rho_{-1/2}}\gg\Lambda^{-1/2},$ (S144)
+the ultraviolet term may be neglected, so that
+$T_{c}\simeq\left(2g_{\rm pair}\rho_{-1/2}\right)^{2}.$ (S145)
+Thus the $\alpha=-1/2$ IR-singular TDOS converts the BCS-like exponential scale into an algebraically amplified transition scale.
+(VIII) General algebraic regime for $\alpha<0$: For completeness, let $\alpha=-\beta$ with $\beta>0$. Then Eq. (S130) becomes
+$\Pi^{R}(0;T)=\frac{\rho_{-\beta}}{\beta}\left(T^{-\beta}-\Lambda^{-\beta}\right).$ (S146)
+The Thouless criterion yields
+$1-g_{\rm pair}\frac{\rho_{-\beta}}{\beta}\left(T_{c}^{-\beta}-\Lambda^{-\beta}\right)=0,$ (S147)
 hence
-$S_{\lambda}(\omega)=\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S116)
-This is the key result.
-The Fourier transform of the exponential memory kernel itself produces a numerator $\lambda$, but the stationary correlator contains a compensating factor $1/\lambda$.
-As a result, the exact fluctuation spectrum carries no additional factor of $\lambda$ in the numerator.
-(IV) Equivalent derivation in frequency space: The same result follows directly in frequency space. Fourier transforming Eq. (S101) gives
-$(-i\omega+\lambda)\Delta_{\lambda}(\omega)=\xi_{\lambda}(\omega),$ (S117)
-hence
-$\Delta_{\lambda}(\omega)=\frac{1}{\lambda-i\omega}\,\xi_{\lambda}(\omega).$ (S118)
-The noise correlator in frequency space is
-$\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle=2\pi\,2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime}).$ (S119)
-Therefore
-$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle$ $=\frac{\langle\xi_{\lambda}(\omega)\xi_{\lambda^{\prime}}^{*}(\omega^{\prime})\rangle}{(\lambda-i\omega)(\lambda^{\prime}+i\omega^{\prime})}$
-$=2\pi\,\delta(\lambda-\lambda^{\prime})\delta(\omega-\omega^{\prime})\frac{2D_{\lambda}}{\lambda^{2}+\omega^{2}}.$ (S120)
-Defining the spectral density through
-$\langle\Delta_{\lambda}(\omega)\Delta_{\lambda}^{*}(\omega^{\prime})\rangle=2\pi\delta(\omega-\omega^{\prime})S_{\lambda}(\omega),$ (S121)
-one again recovers Eq. (S116).
-This derivation makes clear that the mode-resolved pair correlator is proportional to $D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)$, not to the real part of a single retarded propagator. Indeed,
-$D_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad D_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega},$ (S122)
-so that
-$D_{\lambda}^{R}(\omega)D_{\lambda}^{A}(\omega)=\frac{1}{\lambda^{2}+\omega^{2}}.$ (S123)
-(V) From a single mode to the Cooper-channel polarization: After microscopic electronic degrees of freedom are integrated out, the collective Cooper sector is represented as a continuum of relaxational pair modes.
-Let $\rho_{\rm pair}(\lambda)$ denote the effective density of pair relaxation rates, including both mode degeneracy and noise weight.
-The full polarization kernel is then obtained by summing the mode-resolved fluctuation spectrum over the continuum,
-$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,S_{\lambda}(\omega).$ (S124)
-Absorbing the factor $2D_{\lambda}$ into the definition of $\rho_{\rm pair}(\lambda)$ gives
-$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,\frac{1}{\lambda^{2}+\omega^{2}}.$ (S125)
-This is the Cooper-channel polarization kernel quoted in the main text.
-(VI) Infrared scaling of the polarization kernel: For a generic infrared TDOS
-$\rho_{\rm pair}(\lambda)\sim\lambda^{\eta},\qquad\lambda\to 0,$ (S126)
-Eq. (S125) gives
-$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}.$ (S127)
-Setting $\lambda=|\omega|x$,
-$\Pi(\omega)$ $\sim|\omega|^{\eta-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\eta}}{1+x^{2}}.$ (S128)
-Hence the universal infrared scaling is
-$\Pi(\omega)\sim|\omega|^{\eta-1}.$ (S129)
-In particular, for a flat TDOS,
-$\rho_{\rm pair}(\lambda\to 0)=\rho_{0},$ (S130)
+$T_{c}=\left(\Lambda^{-\beta}+\frac{\beta}{g_{\rm pair}\rho_{-\beta}}\right)^{-1/\beta}.$ (S148)
+In the weak-coupling infrared-dominated regime,
+$T_{c}\simeq\left(\frac{g_{\rm pair}\rho_{-\beta}}{\beta}\right)^{1/\beta}.$ (S149)
+Thus the retarded Keldysh formulation yields the following general result:
+$\alpha=0\;\Rightarrow\;T_{c}\sim\Lambda e^{-1/(g_{\rm pair}\rho_{0})},\quad\alpha<0\;\Rightarrow\;T_{c}\sim g_{\rm pair}^{\,1/|\alpha|}.$ (S150)
+Final remark. The derivation above concerns only the retarded linear-response criterion for the superconducting instability. The Keldysh fluctuation propagator,
+$L^{K}(\omega)=L^{R}(\omega)\Sigma^{K}(\omega)L^{A}(\omega),$ (S151)
+and the corresponding fluctuation correlator will be analyzed separately in the subsequent part of this Appendix.
+The essential logical point is that the transition scale $T_{c}$ is fixed first by the pole of the retarded ladder, while the fluctuation sector is discussed afterwards.
+(IX) Keldysh fluctuation propagator of the collective pair field: Having determined the transition scale from the retarded linear-response kernel, we now briefly summarize the corresponding fluctuation sector.
+In the Keldysh formalism, the fluctuation propagator is obtained from the standard matrix relation
+$L^{K}(\omega)=L^{R}(\omega)\,\Sigma^{K}(\omega)\,L^{A}(\omega),$ (S152)
+where $\Sigma^{K}(\omega)$ is the Keldysh noise kernel.
+For the present purpose, the precise microscopic form of $\Sigma^{K}$ is not essential; what matters is that $L^{K}$ is built from the same retarded and advanced collective propagators that control the instability criterion.
+For a single relaxational pair mode with rate $\lambda$,
+$L_{\lambda}^{R}(\omega)=\frac{1}{\lambda-i\omega},\qquad L_{\lambda}^{A}(\omega)=\frac{1}{\lambda+i\omega}.$ (S153)
+If the corresponding Keldysh noise strength is denoted by $\Sigma_{\lambda}^{K}(\omega)$, then
+$L_{\lambda}^{K}(\omega)=L_{\lambda}^{R}(\omega)\,\Sigma_{\lambda}^{K}(\omega)\,L_{\lambda}^{A}(\omega)=\frac{\Sigma_{\lambda}^{K}(\omega)}{\lambda^{2}+\omega^{2}}.$ (S154)
+In the simplest low-frequency approximation, one may treat $\Sigma_{\lambda}^{K}(\omega)$ as slowly varying and replace it by an effective constant amplitude.
+Then the essential frequency dependence is Lorentzian,
+$L_{\lambda}^{K}(\omega)\propto\frac{1}{\lambda^{2}+\omega^{2}}.$ (S155)
+(X) Continuum TDOS representation of the fluctuation kernel: Summing over the continuum of collective pair modes, the Keldysh fluctuation kernel takes the form
+$\Pi^{K}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{\rm pair}(\lambda)\,L_{\lambda}^{K}(\omega).$ (S156)
+In the low-frequency regime, using Eq. (S155), this reduces to
+$\Pi^{K}(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{\rm pair}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S157)
+For the infrared TDOS
+$\rho_{\rm pair}(\lambda)=\rho_{\alpha}\lambda^{\alpha},$ (S158)
 one obtains
-$\Pi(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}$
-$=\frac{\rho_{0}}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right)$
-$\xrightarrow[|\omega|\ll\Lambda]{}\frac{\pi\rho_{0}}{2}\frac{1}{|\omega|}.$ (S131)
-Equation (S131) establishes the memory-dominated Cooper enhancement.
-A flat relaxation-rate spectrum therefore converts the marginal logarithmic pairing growth of conventional Ohmic dynamics into an algebraically enhanced infrared instability of the Cooper channel.
-Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics.
-(a) Absolute spectral line shape at $q=0$: the TDOS theory yields a broad continuum, whereas HM produces a sharper peak.
-(b) Momentum dependence of the TDOS spectral function: varying $q$ shifts the hump through $\Delta_{q}=r+cq^{2}$.
-(c) TDOS prediction (collapse): when plotted versus the scaling variable $u=\Delta_{q}/\rho_{0}+\ln(\Lambda/|\omega|)$, the TDOS spectra collapse onto a universal curve, $A(\omega,q)=\frac{\pi}{\rho_{0}}\,[u^{2}+(\pi/2)^{2}]^{-1}$.
-(d) Control test: HM spectra do not exhibit the same log-shift collapse.
-(e) Finite-temperature regulation implemented by $|\omega|\rightarrow\max\{|\omega|,cT\}$ yields infrared saturation.
-(f) Corresponding $T$-diagnostic at $q=0$: $A_{T}(\omega\!\to\!0)$ varies systematically with $\ln(\Lambda/T)$, providing a practical fitting protocol.
-All curves are shown in arbitrary units with consistent prefactor conventions.
-Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
-In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime.
-We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
-(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
-$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S132)
-where
-$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S133)
-and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
-$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S134)
-Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.
-(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
-$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S135)
-Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
-$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S136)
-$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S137)
-one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
-$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S138)
-(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
-$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S139)
-For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
-$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S140)
-Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
-(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S138) is maximized when
-$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S141)
-which yields
-$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S142)
-For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
-(V) Universal scaling collapse: Define the scaling variable
-$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S143)
-Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S138) reduces to the universal form
-$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S144)
-so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
-(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
-$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S145)
-yielding
-$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S146)
-For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
-Physical interpretation.
-Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part.
-The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
-(VII) Experimental diagnostics and scaling tests: Figure S1 summarizes a set of practical diagnostics that distinguish the memory–dominated TDOS regime from conventional overdamped critical dynamics such as the Hertz–Millis (HM) theory.
-Panel (a) compares the absolute spectral line shapes at $q=0$ predicted by the two frameworks.
-The HM theory produces a relatively narrow Lorentzian-like peak, reflecting dynamics controlled by a small number of overdamped bosonic modes.
-In contrast, Eq.
-(S138) yields a much broader continuum with slowly varying logarithmic tails, arising from the collective response of an extensive continuum of slow relaxation modes.
-Panel (b) illustrates the predicted momentum dependence of the TDOS spectral function.
-Because the tuning parameter enters through $\Delta_{\mathbf{q}}=r+cq^{2}$, varying the momentum effectively shifts the spectral hump according to the scale $|\omega_{*}|=\Lambda\exp(\Delta_{\mathbf{q}}/\rho_{0})$ derived in Eq.
-(S142).
-This produces a characteristic logarithmic displacement of the hump position as $q$ changes.
-Panels (c) and (d) demonstrate the most direct experimental test of the TDOS framework. Introducing the scaling variable
-$u(\omega,\mathbf{q})=\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S147)
-all TDOS spectra collapse onto the universal curve (S144).
-This collapse reflects the fact that the full line shape depends only on the combined variable $u$.
-By contrast, spectra generated by the conventional Hertz–Millis form do not collapse under the same transformation, providing a clear diagnostic distinction.
-Panel (e) illustrates the finite–temperature regulation discussed in Eq.
-(S146).
-Replacing $|\omega|$ by $\max\{|\omega|,cT\}$ cuts off the logarithmic infrared growth and produces saturation of the low–frequency spectral weight.
-The resulting spectra evolve smoothly from the zero–temperature form to a temperature–dominated plateau.
-Finally, panel (f) shows a convenient experimental diagnostic at $q\simeq 0$.
-In the regime $|\omega|\ll T$, Eq.
-(S146) predicts that the infrared spectral weight varies systematically with $\ln(\Lambda/T)$.
-Measuring $A_{T}(\omega\!\rightarrow 0)$ as a function of temperature therefore provides a direct way to extract the TDOS parameter $\rho_{0}$ and test the predicted logarithmic scaling.
-Taken together, these diagnostics provide a practical experimental protocol for identifying the memory–dominated TDOS universality class.
-In particular, the logarithmic peak shift and the universal collapse of line shapes constitute strong signatures of collective dynamics governed by an extensive slow–mode reservoir rather than by a single overdamped critical mode.
-Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
+$\Pi^{K}(\omega)\propto\rho_{\alpha}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda^{\alpha}}{\lambda^{2}+\omega^{2}}.$ (S159)
+Rescaling $\lambda=|\omega|x$, one finds
+$\Pi^{K}(\omega)$ $\propto\rho_{\alpha}\int_{0}^{\Lambda/|\omega|}d(|\omega|x)\,\frac{(|\omega|x)^{\alpha}}{|\omega|^{2}(1+x^{2})}$
+$=\rho_{\alpha}\,|\omega|^{\alpha-1}\int_{0}^{\Lambda/|\omega|}dx\,\frac{x^{\alpha}}{1+x^{2}}.$ (S160)
+Hence, up to an $\alpha$-dependent numerical factor,
+$\Pi^{K}(\omega)\propto|\omega|^{\alpha-1}.$ (S161)
+(XI) Representative infrared cases: The fluctuation kernel therefore exhibits the following infrared behavior.
+(a) $\alpha=0$: flat TDOS
+$\Pi^{K}(\omega)\propto\frac{1}{|\omega|}.$ (S162)
+Thus a flat TDOS produces a strongly enhanced algebraic fluctuation spectrum even though the transition scale itself remains BCS-like, $T_{c}=\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right)$.
+(b) $\alpha=-1/2$: supercritical TDOS
+$\Pi^{K}(\omega)\propto|\omega|^{-3/2}.$ (S163)
+The fluctuation enhancement is now even stronger than in the flat case, consistent with the algebraically amplified transition scale derived above.
+(c) general $\alpha<0$:
+$\Pi^{K}(\omega)\propto|\omega|^{-1-|\alpha|}.$ (S164)
+showing that increasingly singular infrared TDOS leads to progressively stronger fluctuation enhancement.
+(XII) Physical interpretation: The distinction between the retarded and Keldysh sectors should now be clear.
+The superconducting transition temperature is fixed by the pole of the retarded ladder, $1-g_{\rm pair}\Pi^{R}(0;T_{c})=0$, whereas the Keldysh kernel in Eq.
+(S157) describes the strength of collective Cooper-pair fluctuations.
+Accordingly, a flat TDOS already produces strong algebraic infrared fluctuations,
+$\Pi^{K}(\omega)\propto|\omega|^{-1},$ (S165)
+while a singular TDOS with $\alpha<0$ enhances both the fluctuation spectrum and, through the retarded kernel, the transition scale itself.
+In this sense, the fluctuation sector and the instability sector are closely related but logically distinct: the retarded component determines the transition point, whereas the Keldysh component quantifies the infrared strength of pair-field fluctuations near that transition.
+Appendix E Appendix E: Possible connections between flat TDOS and scale–free dynamical phenomena
 In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text.
@@
 describes a broad continuum rather than a sharp quasiparticle peak.
-(V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
-$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (S164)
-with
-$\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S165)
-For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one obtains
-$\Pi(\omega)\sim\frac{1}{|\omega|},$ (S166)
-which leads to an algebraically enhanced transition scale
-$T_{c}\sim\rho_{0}E_{\rm pair}.$ (S167)
-The slow–mode reservoir therefore acts as a dynamical amplifier of intrinsic electronic pairing correlations.
+(V) Consequences for superconducting pairing: In the main text we have shown that superconducting pairing is governed by the retarded Cooper-channel kernel entering the Thouless criterion.
+Within the relaxation-spectrum framework the pairing susceptibility is written as
+$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi^{R}(\omega),$ (S182)
+where the retarded polarization kernel takes the form
+$\Pi^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S183)
+For a flat infrared TDOS, $\rho(\lambda\rightarrow 0)=\rho_{0}$, the real part of the retarded kernel exhibits logarithmic scaling, $\mathrm{Re}\,\Pi^{R}(\omega)\sim\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)$.
+Evaluating at the thermal scale $\omega\sim T$ yields the BCS–like transition condition $1-g_{\rm pair}\,\rho_{0}\ln\frac{\Lambda}{T_{c}}=0$, leading to
+$T_{c}\sim\Lambda\exp\!\left(-\frac{1}{g_{\rm pair}\rho_{0}}\right).$ (S184)
+It is important to distinguish this retarded response from the fluctuation spectrum of the relaxational modes. The latter is governed by the Keldysh correlator,
+$\Pi^{K}(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},$ (S185)
+which reflects the strong infrared enhancement of fluctuations. While this divergence signals the presence of an extensive reservoir of slow modes, it does not directly determine the transition scale.
+Thus, in the flat TDOS regime the slow-mode reservoir produces strong infrared fluctuations and long-time memory, but the superconducting transition remains marginal, retaining a BCS-like exponential form.
+Only when the TDOS becomes infrared-singular, $\rho(\lambda)\sim\lambda^{\alpha}$ with $\alpha<0$, does the retarded kernel itself acquire a power-law divergence, leading to an algebraically enhanced transition scale.
 Summary.
-These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems.
-Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes.
-Flat infrared TDOS simultaneously produces long–time memory, scale–free lifetime statistics, $1/f$ noise, and algebraic enhancement of superconducting pairing.
+These results indicate that the infrared organization of the relaxation spectrum provides a unified dynamical framework for a wide range of scale–free phenomena observed in correlated systems.
+A flat infrared TDOS generates long–time memory, $K(t)\sim 1/t$, scale–free lifetime distributions, $P(\tau)\sim\tau^{-2}$, and $1/f$ noise, reflecting the presence of a dense continuum of slow relaxation modes.
+At the same time, the superconducting instability is governed by the retarded response of this spectrum.
+In the flat TDOS regime the Cooper channel exhibits logarithmic infrared enhancement, leading to a BCS–like exponential transition scale controlled by the infrared spectral weight.
+A qualitatively new regime emerges only when the TDOS becomes infrared-singular, where the retarded kernel itself becomes power-law divergent and superconductivity is algebraically enhanced.
+Within this perspective, superconductivity and anomalous dynamics arise from a common origin: the infrared spectral structure of the Liouvillian relaxation spectrum, rather than from the exchange of a specific bosonic mediator.
+Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics.
+(a) Absolute spectral line shape at $q=0$: the TDOS theory yields a broad continuum, whereas HM produces a sharper peak.
+(b) Momentum dependence of the TDOS spectral function: varying $q$ shifts the hump through $\Delta_{q}=r+cq^{2}$.
+(c) TDOS prediction (collapse): when plotted versus the scaling variable $u=\Delta_{q}/\rho_{0}+\ln(\Lambda/|\omega|)$, the TDOS spectra collapse onto a universal curve, $A(\omega,q)=\frac{\pi}{\rho_{0}}\,[u^{2}+(\pi/2)^{2}]^{-1}$.
+(d) Control test: HM spectra do not exhibit the same log-shift collapse.
+(e) Finite-temperature regulation implemented by $|\omega|\rightarrow\max\{|\omega|,cT\}$ yields infrared saturation.
+(f) Corresponding $T$-diagnostic at $q=0$: $A_{T}(\omega\!\to\!0)$ varies systematically with $\ln(\Lambda/T)$, providing a practical fitting protocol.
+All curves are shown in arbitrary units with consistent prefactor conventions.
+Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
+In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime.
+We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
+(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
+$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S186)
+where
+$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S187)
+and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
+$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S188)
+Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.
+(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
+$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S189)
+Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
+$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S190)
+$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S191)
+one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
+$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S192)
+(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
+$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S193)
+For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
+$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S194)
+Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
+(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S192) is maximized when
+$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S195)
+which yields
+$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S196)
+For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
+(V) Universal scaling collapse: Define the scaling variable
+$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S197)
+Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S192) reduces to the universal form
+$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S198)
+so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
+(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale.
+A minimal implementation consistent with the Matsubara-frequency structure discussed in Appendix D is
+$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S199)
+yielding
+$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S200)
+For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
+Physical interpretation.
+Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part.
+The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
+(VII) Experimental diagnostics and scaling tests: Figure S1 summarizes a set of practical diagnostics that distinguish the memory–dominated TDOS regime from conventional overdamped critical dynamics such as the Hertz–Millis (HM) theory.
+Panel (a) compares the absolute spectral line shapes at $q=0$ predicted by the two frameworks.
+The HM theory produces a relatively narrow Lorentzian-like peak, reflecting dynamics controlled by a small number of overdamped bosonic modes.
+In contrast, Eq.
+(S192) yields a much broader continuum with slowly varying logarithmic tails, arising from the collective response of an extensive continuum of slow relaxation modes.
+Panel (b) illustrates the predicted momentum dependence of the TDOS spectral function.
+Because the tuning parameter enters through $\Delta_{\mathbf{q}}=r+cq^{2}$, varying the momentum effectively shifts the spectral hump according to the scale $|\omega_{*}|=\Lambda\exp(\Delta_{\mathbf{q}}/\rho_{0})$ derived in Eq.
+(S196).
+This produces a characteristic logarithmic displacement of the hump position as $q$ changes.
+Panels (c) and (d) demonstrate the most direct experimental test of the TDOS framework. Introducing the scaling variable
+$u(\omega,\mathbf{q})=\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S201)
+all TDOS spectra collapse onto the universal curve (S198).
+This collapse reflects the fact that the full line shape depends only on the combined variable $u$.
+By contrast, spectra generated by the conventional Hertz–Millis form do not collapse under the same transformation, providing a clear diagnostic distinction.
+Panel (e) illustrates the finite–temperature regulation discussed in Eq.
+(S200).
+Replacing $|\omega|$ by $\max\{|\omega|,cT\}$ cuts off the logarithmic infrared growth and produces saturation of the low–frequency spectral weight.
+The resulting spectra evolve smoothly from the zero–temperature form to a temperature–dominated plateau.
+Finally, panel (f) shows a convenient experimental diagnostic at $q\simeq 0$.
+In the regime $|\omega|\ll T$, Eq.
+(S200) predicts that the infrared spectral weight varies systematically with $\ln(\Lambda/T)$.
+Measuring $A_{T}(\omega\!\rightarrow 0)$ as a function of temperature therefore provides a direct way to extract the TDOS parameter $\rho_{0}$ and test the predicted logarithmic scaling.
+Taken together, these diagnostics provide a practical experimental protocol for identifying the memory–dominated TDOS universality class.
+In particular, the logarithmic peak shift and the universal collapse of line shapes constitute strong signatures of collective dynamics governed by an extensive slow–mode reservoir rather than by a single overdamped critical mode.